### v2

# Adaptive Width Neural Networks

Abstract
For almost 70 years, researchers have mostly relied on hyper-parameter tuning to pick the width of neural networks’ layers out of many possible choices. This paper challenges the status quo by introducing an easy-to-use technique to learn an unbounded width of a neural network’s layer during training. The technique does not rely on alternate optimization nor hand-crafted gradient heuristics; rather, it jointly optimizes the width and the parameters of each layer via simple backpropagation. We apply the technique to a broad range of data domains such as tables, images, texts, and graphs, showing how the width adapts to the task’s difficulty. By imposing a soft ordering of importance among neurons, it is possible to truncate the trained network at virtually zero cost, achieving a smooth trade-off between performance and compute resources in a structured way. Alternatively, one can dynamically compress the network with no performance degradation. In light of recent foundation models trained on large datasets, believed to require billions of parameters and where hyper-parameter tuning is unfeasible due to huge training costs, our approach stands as a viable alternative for width learning.
Neural Networks, Learning the Number of Neurons, Adaptive Width Learning, Dynamic Architectures, Information Compression, Variational Inference
1 Introduction
Since the construction of the Mark I Perceptron machine (Rosenblatt, 1958) the effective training of neural networks has remained an open research problem of great academic and practical value. The Mark I solved image recognition tasks by exploiting a layer of 512 fixed “association units” that in modern language correspond to the hidden units of a (possibly) Multi-Layer Perceptron (MLP). MLPs possess universal approximation capabilities when assuming arbitrary width (Cybenko, 1989) and sigmoidal activations, and their convergence to good solutions was studied, for instance, in Rumelhart et al. (1986) where the backpropagation algorithm was described as “simple, easy to implement on parallel hardware”, and improvable by other techniques such as momentum that preserve locality of weight updates.
Yet, after almost 70 years of progress (LeCun et al., 2015), the vast majority of neural networks, be they shallow or deep, still rely on a fixed choice of the number of neurons in their hidden layers. This property is typically treated as an architectural design choice, one of the many hyper-parameters that have to be carefully tuned whenever we approach a new task (Wolpert, 1996). The tuning process has many names, such as model selection, hyper-parameter tuning, and cross-validation, and it is associated with non-negligible costs: Different architectural configurations are to be trained until one that performs best on a validation set is selected (Mitchell, 1997). The configurations’ space grows exponentially in the number of layers, so practitioners often resort to shortcuts such as picking a specific number of hidden units for all layers, which greatly reduces the search space together with the chances of selecting a better architecture for the task. Other techniques to tune hyper-parameters include constructive approaches (Fahlman & Lebiere, 1989; Wu et al., 2020), which alternate parameter optimization and creation of new neurons, natural gradient-based heuristics that dynamically modify the network (Mitchell et al., 2023), bi-level optimization (Franceschi et al., 2018), and neural architecture search (White et al., 2023), which often requires separate training runs for each configuration.
The hyper-parameters’ space exploration problem is exacerbated by the steep increase in size of recent neural architectures for language (Brown et al., 2020) and vision (Zhai et al., 2022), for example, where parameters are in the order of billions to (supposedly) accommodate for a huge dataset. Training these models requires an amount of time, compute power, and energy that currently makes it unfeasible for most institutions to perform a thorough model selection and find good width parameters; the commonly accepted compromise is to stick to previously successful hyper-parameter choices. This may also explain why network pruning (Blalock et al., 2020; Mishra et al., 2021), distillation (Zhang et al., 2019) and quantization (Mishra et al., 2021) techniques have recently been in the spotlight, as they trade-off hardware requirements and performance.
This work introduces a simple and easy to use technique to learn the width of each neural network’s layer without imposing upper bounds (we refer to it as unbounded width). The width of each layer is dynamically adjusted during backpropagation using the latest deep learning libraries (Paszke et al., 2017), and it only requires a slight modification to the neural activations that does not alter the ability to parallelize computation. The technical strategy is to impose a soft ordering of hidden units by exploiting any monotonically decreasing function with unbounded support on natural numbers, which has a number of benefits. For instance, we do not need to fix a maximum number of neurons, which is typical of orthogonal approaches like supernetworks (White et al., 2023). As a by-product, we can achieve a straightforward trade-off between parametrization and performance by removing the last rows/columns of weight matrices, which corresponds to ignoring the “least important” neurons in the computation. Finally, we break symmetries in the parametrization of neural networks: it is not possible anymore to obtain an equivalent neural network behavior by permuting the weight matrices, which should in principle reduce the “jostling” effect where symmetric parametrizations compete at the beginning of training (Barber, 2012).
We test the proposed method on an MLP for tabular data, a Convolutional Neural Networks (CNN) (LeCun et al., 1989) for images, a Transformer architecture (Vaswani et al., 2017) for text, and a Deep Graph Network (DGN) (Micheli, 2009; Scarselli et al., 2009) for graphs, to showcase a broad scope of applicability. Empirical results suggest, as may be expected, that the width adapts to the task’s difficulty with the same performance of a fixed-width baseline. We investigate compressing of the network at training time while preserving accuracy, as well as a post-hoc truncation inducing a controlled trade-off at zero additional cost. Ablations suggest that the learned width is not influenced by the starting width (under bounded activations) nor by the batch size, advocating for a potentially large reduction of the hyper-parameter configuration space at model selection time.
2 Related Work
Constructive methods dynamically learn the width of neural networks and are related in spirit to this work. The cascade correlation algorithm (Fahlman & Lebiere, 1989) alternates standard training with the creation of a new hidden unit minimizing the neural network’s residual error. Similarly, the firefly network descent (Wu et al., 2020) grows the width and depth of a network every $N$ training epochs via gradient descent on a dedicated loss. Yoon et al. (2018) propose an ad-hoc algorithm for lifelong learning that grows the network by splitting and duplicating units to learn new tasks. Wu et al. (2019) alternate training the network and then splitting existing neurons into offspring with equal weights. These works mostly focus on growing the neural network; Mitchell et al. (2023) propose natural gradient-based heuristics to grow/shrink layers and hidden units of MLPs and CNNs. The main difference from our work is that we grow and shrink the network by simply computing the gradient of the loss, without relying on human-defined heuristics. The unbounded depth network of Nazaret & Blei (2022), from which we draw inspiration, learns the number of layers of neural networks. Compared to that work, we focus our attention to the number of neurons, modifying the internals of the architecture rather than instantiating a multi-output one. Finally, we mention Bayesian nonparametrics approaches (Orbanz & Teh, 2010) that learn a potentially infinite number of clusters in an unsupervised fashion.
Orthogonal Methods
Neural Architecture Search (NAS) is an automated process that designs neural networks for a given task (Elsken et al., 2019b; White et al., 2023) and has been applied to different contexts (Zoph et al., 2018; Liu et al., 2019; So et al., 2019). Typically, neural network elements are added, removed, or modified based on validation performance (Elsken et al., 2019a; White et al., 2021, 2023), by means of reinforcement learning (Zoph & Le, 2016), evolutionary algorithms (Real et al., 2019), and gradient-based approaches (Liu et al., 2019). Typical NAS methods require enormous computational resources, sometimes reaching thousands of GPU days (Zoph & Le, 2016), due to the retraining of each new configuration. While recent advances on one-shot NAS models (Brock et al., 2018; Pham et al., 2018; Bender et al., 2018) have drastically reduced the computational costs, they mostly focus on CNNs, assume a bounded search space, and do not learn the width. As such, NAS methods are complementary to our approach.
Bi-level optimization algorithms have also been used for hyper-parameter tuning (Franceschi et al., 2018), where hyper-parameters are the variables of the outer objective and the model parameters those of the inner objective. The solution sets of the inner problem are usually not available in closed form, which has been partly addressed by repeated application of (stochastic) gradient descent (Domke, 2012; Maclaurin et al., 2015; Franceschi et al., 2017). These methods are restricted to continuous hyper-parameters’ optimization, and cannot therefore be applied to width optimization.
Finally, pruning (Blalock et al., 2020) and distillation (Hinton et al., 2015) are two methods that (only) reduce the size of neural networks by trading-off performances; the former deletes neural connections (Mishra et al., 2021) or entire neurons (Valerio et al., 2022; Dufort-Labbé et al., 2024), the latter trains a smaller network (student) to mimic a larger one (teacher) (Gou et al., 2021). Compared to most pruning approaches, our work can delete connections and reduce the model’s memory; compared to distillation, we do not necessarily need a new training. These techniques, however, can be easily combined with our approach.
3 Adaptive Width Learning
[…]
Figure 1: (Left) We depict the graphical model of AWNN, where observable random variables are dark and latent ones are white. (Middle) The distribution $f_{\ell}$ over hidden units’ importance at layer $\ell$ is parametrized by $\lambda_{\ell}$. The width of layer $\ell$ is chosen as the quantile function of the distribution $f_{\ell}$ evaluated at $k$ and denoted by $D_{\ell}$. (Right) The hidden units’ activations at layer $\ell$ are rescaled by their importance.
We introduce a general probabilistic framework, called Adaptive Width Neural Networks (AWNN) for convenience, showing how (ultimately simple) design choices arise from a variational inference treatment of a graphical model.
We are given a dataset of $N$ i.i.d. samples $(x_{i},y_{i})$, with input $x_{i}\in\mathbb{R}^{F},F\in\mathbb{N}^{+}$ and target $y_{i}$ whose domain depends on whether the task is regression or classification. The learning objective is to maximize
$\log\prod_{i=1}^{N}p(y_{i}|x_{i})=\sum_{i=1}^{N}\log p(y_{i}|x_{i})$ (1)
[…]
$p(y_{i},\boldsymbol{\lambda},\boldsymbol{\theta}|x_{i})=p(y_{i}|\boldsymbol{\lambda},\boldsymbol{\theta},x_{i})p(\boldsymbol{\lambda})p(\boldsymbol{\theta})$ (2)
$p(\boldsymbol{\lambda})=\prod^{L}_{\ell=1}p(\lambda_{\ell})=\prod^{L}_{\ell=1}\mathcal{N}(\lambda_{\ell};\mu^{\lambda}_{\ell},\sigma^{\lambda}_{\ell})$ (3)
$p(\boldsymbol{\theta})=\prod^{L}_{\ell=1}\prod^{\infty}_{n=1}p(\theta_{\ell n})\text{=}\prod^{L}_{\ell=1}\prod^{\infty}_{n=1}\mathcal{N}(\theta_{\ell n};\mathbf{0},\text{diag}(\sigma^{\theta}_{\ell}))$ (4)
[…]
Here, $\sigma^{\theta}_{\ell},\mu^{\lambda}_{\ell},\sigma^{\lambda}_{\ell}$ are hyper-parameters. The MLP is parametrized by realizations $\boldsymbol{\lambda},\boldsymbol{\theta}$, so it relies on a finite number of neurons and outputs either class probabilities (classification) or the mean of a Gaussian distribution (regression) to parametrize $p(y_{i}|\boldsymbol{\lambda},\boldsymbol{\theta},x_{i})$ depending on the task. Maximizing Equation (1), however, requires computing the above integral, which is intractable. Therefore, we turn to mean-field variational inference (Jordan et al., 1999; Blei et al., 2017) to maximize an expected lower bound (ELBO) instead. This requires to define a distribution over the latent variables $q(\boldsymbol{\lambda},\boldsymbol{\theta})$ and re-phrase the objective as:
$\sum_{i=1}^{N}\log p(y_{i}|x_{i})\geq\sum_{i=1}^{N}\mathbb{E}_{q(\boldsymbol{\lambda},\boldsymbol{\theta})}\left[\log\frac{p(y_{i},\boldsymbol{\lambda},\boldsymbol{\theta}|x_{i})}{q(\boldsymbol{\lambda},\boldsymbol{\theta})}\right],$ (6)
where $q(\boldsymbol{\lambda},\boldsymbol{\theta})$ is parametrized by learnable variational parameters. We factorize the variational distribution into:
$q(\boldsymbol{\lambda},\boldsymbol{\theta})=q(\boldsymbol{\lambda})q(\boldsymbol{\theta}|\boldsymbol{\lambda})$ (7)
$q(\boldsymbol{\lambda})=\prod_{\ell=1}^{L}q(\lambda_{\ell})=\prod_{\ell=1}^{L}\mathcal{N}(\lambda_{\ell};\nu_{\ell},1)$ (8)
$q(\boldsymbol{\theta}|\boldsymbol{\lambda})=\prod_{\ell=1}^{L}\prod_{n=1}^{D_{\ell}}q(\theta_{\ell n})\prod_{D_{\ell}+1}^{\infty}p(\theta_{\ell n})$ (9)
$q(\theta_{\ell n})=\mathcal{N}(\theta_{\ell n};\text{diag}(\rho_{\ell n}),\mathbf{I}).$ (10)
$D_{\ell}=\text{quantile function of }f_{\ell}(\cdot;\lambda_{\ell})\text{ evaluated at }k$ (11)
The value $k$ is a hyper-parameter, $\nu_{\ell},\rho_{\ell n}$ are variational parameters and, as before, we define $\boldsymbol{\rho}_{\ell}=\left\{\rho_{\ell n}\right\}_{n=1}^{D_{\ell}}$, $\boldsymbol{\rho}=\left\{\boldsymbol{\rho}_{\ell}\right\}_{\ell=1}^{L}$ and $\boldsymbol{\nu}=\left\{\nu_{\ell}\right\}_{\ell=1}^{L}$. Note that the set of variational parameters is finite. The truncated width $D_{\ell}$, that is the finite number of neurons at layer $\ell$, is computed as the quantile function evaluated at $k$ of a distribution [^2] General functions are allowed if a threshold can be computed. $f_{\ell}$ with infinite support over $\mathbb{N}^{+}$, parametrized by $\lambda_{\ell}$. W.l.o.g., we implement $f_{\ell}$ as a discretized exponential distribution, following the discretization strategy of Roy (2003): For every natural $x$, the discretized distribution relies on the cumulative distribution function (c.d.f.) of the exponential:
$f_{\ell}(x;\lambda_{\ell})=(1-e^{\lambda_{\ell}(x+1)})-(1-e^{\lambda_{\ell}(x)}).$ (12)
We choose the exponential because it is a monotonically decreasing function and allows us to impose an ordering of importance among neurons, as detailed in Section 3.1.
By expanding Equation 6 using the above definitions and approximating the expectations at the first order, i.e., $\mathbb{E}_{q(\boldsymbol{\lambda})}[f(\boldsymbol{\lambda})]$=$f(\boldsymbol{\nu})$ and $\mathbb{E}_{q(\boldsymbol{\theta}|\boldsymbol{\lambda})}[f(\boldsymbol{\theta})]=f(\boldsymbol{\rho})$ as in Nazaret & Blei (2022), we obtain the final form of the objective:
$\sum_{\ell}^{L}\log\frac{p(\nu_{\ell};\mu_{\ell}^{\lambda},\sigma_{\ell}^{\lambda})}{q(\nu_{\ell};\nu_{\ell})}+\sum_{\ell}^{L}\sum_{n=1}^{D_{\ell}}\log\frac{p(\rho_{\ell n};\sigma_{\ell}^{\theta})}{q(\rho_{\ell n};\rho_{\ell n})}+$
$+\sum_{i=1}^{N}\log p(y_{i}|\boldsymbol{\lambda}\text{=}\boldsymbol{\nu},\boldsymbol{\theta}\text{=}\boldsymbol{\rho},x_{i}),$ (13)
where distributions’ parameters are made explicit to distinguish them. The first two terms in the loss regularize the width of the layers and the magnitude of the parameters, respectively, whereas the third is is the predictive loss.
In practice, the finite variational parameters $\boldsymbol{\nu},\boldsymbol{\rho}$ are those used by the neural network in place of $\boldsymbol{\lambda},\boldsymbol{\theta}$, which enables easy optimization via backpropagation. Maximizing Equation (13) will change each variational parameter $\nu_{\ell}$, which in turn will change the value of $D_{\ell}$ during training. If $D_{\ell}$ increases we initialize new neurons and draw their weights from a standard normal distribution, otherwise we remove the excess ones. When implementing mini-batch training, the predictive loss needs to be rescaled by $N/M$, where $M$ is the mini-batch size. From a Bayesian perspective, this means regularizers weigh less if we have more data.
Compared to a fixed-width network with weight decay, we need to choose the values of $\mu^{\lambda}_{\ell},\sigma^{\lambda}_{\ell}$, as well as initialize $\nu_{\ell}$ (shared across layers as there is no particular reason to do otherwise). Therefore, we have two more hyper-parameters compared to the fixed-width network, but we make some considerations: i) it is always possible to use an uninformative prior over $\lambda_{\ell}$, removing the extra hyper-parameters letting the model freely adapt the width of each layer (as is typical of frequentist approaches); ii) the choice of higher level of hyper-parameters is known to be less stringent than that of hyper-parameters themselves (Goel & Degroot, 1981; Bernardo & Smith, 2009), so we do not need to explore many values of $\mu_{\ell}^{\lambda}$ and $\sigma_{\ell}^{\lambda}$; iii) our experiments suggest that AWNN converges to similar widths regardless of the starting point $\nu_{\ell}$, so that we may just need to perform model selection over one/two sensible initial values.
3.1 Imposing a Soft Ordering on Neurons’ Importance
Now that the learning objective has been formalized, the missing ingredient is the definition of the neural network $p(y_{i}|\boldsymbol{\lambda}\text{=}\boldsymbol{\nu},\boldsymbol{\theta}\text{=}\boldsymbol{\rho},x_{i})$. Compared to a standard MLP, we need to make use of the variational parameters $\boldsymbol{\nu}$ that affect the truncation width at each hidden layer, whereas $\boldsymbol{\rho}$ corresponds to the the usual MLP weights. We choose a monotonically decreasing function $f_{\ell}$ for the simple reason that when a new neuron is added, its relative importance is low and will not drastically impact the network. As a result, we are imposing a soft ordering of importance among neurons.
We modify the classical activation $h^{\ell}_{j}$ of a hidden neuron $j$ at layer $\ell$ as
$h^{\ell}_{j}=\sigma\left(\sum_{k=1}^{D_{\ell-1}}w^{\ell}_{jk}h^{\ell-1}_{k}\right)f_{\ell}(j;\nu_{\ell}),$ (14)
where $D_{\ell-1}$ is the truncated width of the previous layer, $\sigma$ is a non-linear activation function and $w^{\ell}_{jk}\in\rho_{\ell j}$. That is, we rescale the activation of each neuron $k$ by its “importance” $f_{\ell}(j;\nu_{\ell})$. Note that the bias parameter is taken into account by concatenating a dummy value 1 to $h^{\ell-1}_{k}$.
It is easy to see that, in theory, the optimization algorithm could rescale the weights of the next layer by a factor $1/f_{\ell}(j;\nu_{\ell})$ to compensate for the term $f_{\ell}(j;\nu_{\ell})$. This may lead to a degenerate situation, which is actually quite common in classical MLPs, where the activations of the first neurons are small relative to the others, thus breaking the soft-ordering and wasting neurons. There are two strategies to address this undesirable effect. The first is to regularize the magnitude of the weights thanks to the prior $p(\theta_{\ell+1n})$, so that it may be difficult to compensate for the least important neurons that have a high $1/f_{\ell}(j)$. The second and less obvious strategy is to prevent the units’ activations of the current layer to compensate for high values by bounding their range. We apply both strategies to our experiments.
3.2 Rescaled Weight Initialization for Deep AWNN
Rescaling the activations of hidden units using Equation 14 causes activations of deeper layers to quickly decay to zero, as shown in Figure 2 for a 5-layer AWNN MLP with ReLU nonlinearity initialized using the well known kaiming scheme (He et al., 2015). This affects convergence since gradients get close to zero and it becomes impractical to train deep AWNN MLPs. We therefore derive a rescaled kaiming weight that guarantees that the variance of activation across layers is constant at initialization.
Theorem 3.1.
Let us consider an MLP with activations as in Equation 14 and ReLU nonlinearity. At initialization, given $\alpha_{j}^{\ell}=\sigma\left(\sum_{k=1}^{D_{\ell-1}}w^{\ell}_{jk}h^{\ell-1}_{k}\right)$
$\mathrm{Var}[w_{j*}^{\ell}]=\frac{2}{\sum_{j=1}^{D_{\ell-1}}f^{2}_{\ell}(j)}\Rightarrow\mathrm{Var}[\alpha_{j}^{\ell}]\approx\mathrm{Var}[\alpha_{j}^{\ell-1}]$ (15)
Proof.
See Appendix A for an extended proof. ∎
The extended proof shows there is a connection between the variance of activations and the variance of gradients. In particular, the sufficient conditions over $\mathrm{Var}[w_{j*}^{\ell}]$ are identical if one initializes all $\nu_{\ell}$ in the same way for all layers. Therefore, if we initialize weights from a Gaussian distribution with standard deviation $\frac{\sqrt{2}}{\sqrt{\sum_{j=1}^{D_{\ell-1}}f^{2}_{\ell}(j)}}$, we guarantee that at initialization the variance of the deep network’s gradients will be constant. The effect of the new initialization (dubbed “kaiming+”) can be seen in Figure 2, where the distribution of activation values at initialization does not collapse in subsequent layers. This change drastically impacts overall convergence on the SpiralHard synthetic dataset (described in Section 4), where it appears it would be otherwise hard to converge using a standard kaiming initialization (please refer to Appendix A for a visual example).
Figure 2: Effect of the rescaled initialization scheme (“kaiming+”) on a ReLU-based AWNN MLP, where neurons’ activations are computed using Equation 14. Compared to the standard initialization, the variance of activations agrees with the theoretical result.
Algorithm 1 summarizes the main changes to the training procedure, namely the new initialization and the update of the model’s truncated width at each training step.
3.3 Future Directions and Limitations
MLPs are ubiquitous in modern deep architectures. They are used at the very end of CNNs, in each encoder and decoder layer of Transformer, and they process messages coming from neighbors in DGNs. Our experiments focus on MLPs to showcase AWNN’s broad applicability, but there are many other scenarios where one can apply AWNN’s principles. For instance, one could impose a soft ordering of importance on CNNs’ filters at each layer, therefore learning the number of filters during training. Alternatively, one could investigate the application of AWNN to timeseries, since recurrent networks (Hochreiter & Schmidhuber, 1997) can be seen as a special case of an MLP with weight sharing.
From a more theoretical perspective, we believe one could draw connections between our technique and the Information Bottleneck principle (Tishby et al., 2000), which seeks maximally representative (i.e., performance) and compact representations (e.g., width). Finally, the analysis of different functions $f$ to soft-order neurons, for instance a power-law distribution or a sigmoid-like function, may have a different impact on convergence and performance and could be empirically investigated.
The current limitation of AWNN is the relative overhead to change the network at training time. A minimal and naïve implementation has up to 10x cost compared to a fixed network (this relative overhead may reduce for more compute-bound tasks). Besides optimizing our code, other optimizations are conceivable: i) by updating the width every $M$ minibatches rather than at every step; ii) by allocating a larger network and dynamically grow it only if needed; iii) by creating ad-hoc support of AWNN on deep learning libraries like Pytorch. In any case, this overhead is relatively small compared to the exponential size of the hyper-parameter space when cross-validating the number of neurons of a multi-layered architecture.
Algorithm 1 AWNN Training Procedure
1:  Input: Dataset $\mathcal{D}$, initialized AWNN model $\mathcal{M}$ (Section 3.2)
2:  Output: Trained AWNN Model $\mathcal{M}$
3:  for each training epoch do
4:    for batch in $\mathcal{D}$ do
5:       $\mathcal{M}$)
6:       $\hat{y}\leftarrow\mathcal{M}(\text{batch})$
7:       $\text{loss}\leftarrow\text{ELBO}(\mathcal{M},\text{batch},\hat{y})$ // Eq. 13
8:       $\mathcal{M}\leftarrow\text{backpropagation}(\mathcal{M},\text{loss})$
9:    end for
10:  end for
11:  function $\mathcal{M}$):
[…]
13:       $D_{\ell}\leftarrow\text{quantile function of }f_{\ell}(\cdot;\nu_{\ell})\text{\ evaluated at }k$
14:       Use $D_{\ell}$ to update $\boldsymbol{\rho}_{\ell},\boldsymbol{\rho}_{\ell+1}$ // add/remove neurons
15:    end for
16:  end function
4 Experiments and Setup
The purpose of the empirical analysis is not to claim AWNN is generally better than the fixed-width baseline. Rather, we demonstrate how AWNN overcomes the problem of fixing the number of neurons by learning it, thus reducing the amount of configurations to test. As such, due to the nature of this work and similarly to Mitchell et al. (2023), we use the remaining space to thoroughly study the behavior of AWNN, so that it becomes clear how to use it in practice. We first quantitatively verify that AWNN does not harm the performance compared to baseline models and compare the chosen width by means of grid-search model selection with the learned width of AWNN. Secondly, we check that AWNN chooses a higher width for harder tasks, which can be seen as increasing the hypotheses space until the neural network finds a good path to convergence. Third, we verify that convergence speed is not significantly altered by AWNN, so that the main limitation lies in the extra overhead for adapting the network at each training step. As a sanity check, we study conditions under which AWNN’s learned width does not seem to depend on starting hyper-parameters, so that their choice does not matter much. Finally, we analyze other practical advantages of training a neural network under the AWNN framework: the ability to compress information during training or post training, and the resulting trade-offs. Further analyses are in the Appendix.
We now provide the details to reproduce our experiments on 9 datasets. We compare a baseline that undergoes proper hyper-parameter tuning against its AWNN version, where we replace any fixed MLP with an adaptive one. First, we train an MLP on 3 synthetic tabular tasks of increasing binary classification difficulty, namely a double moon, a spiral, and a double spiral that we call SpiralHard. A stratified hold-out split of 70% training/10% validation/20% test for risk assessment is chosen at random for these datasets. Similarly, we consider a ResNet-20 (He et al., 2016) trained on 3 image classification tasks, namely MNIST (LeCun, 1998), CIFAR10, and CIFAR100 (Krizhevsky, 2009), where data splits and preprocessing are taken from the original paper and AWNN is applied to the downstream classifier. In the graph domain, we train a Graph Isomorphism Network (Xu et al., 2019) on the NCI1 and REDDIT-B classification tasks using the same split and evaluation setup of Errica et al. (2020). These tasks are chosen as the structure was found to have a proper influence on the final performance. Here, the first 1 hidden layer MLP as well as the one used in each graph convolutional layer are replaced by adaptive AWNN versions. On all these tasks, the metric of interest is the accuracy. Finally, for the textual domain we train a Transformer architecture (Vaswani et al., 2017) on the Multi30k English-German translation task (Elliott et al., 2016), using a pretrained GPT-2 Tokenizer, and we evaluate the cross-entropy loss over the translated words. On tabular, image, and text-based tasks, an internal validation set (10%) for model selection is extracted from the union of outer training and validation sets, and the best configuration chosen according to the internal validation set is retrained 10 times on the outer train/validation/test splits, averaging test performances. Due to space reasons, we report datasets statistics and the hyper-parameter tried for the fixed and AWNN versions in Appendix B and C, respectively.
5 Results
We begin by discussing the quantitative results of our experiments: Table 1 reports means and standard deviations across the 10 final training runs. In terms of performance, we observe that AWNN is more stable or accurate than a fixed MLP on DoubleMoon, Spiral, and SpiralHard; all other things being equal, it seems that using more neurons and their soft ordering are the main contributing factors to these improvements. On the image datasets, performances of AWNN are comparable to those of the fixed baseline but for CIFAR100, due to an unlucky run that did not converge. In this case, AWNN learns a smaller total width compared to grid search.
Table 1: We compare performances and total width of MLP layers for the fixed and AWNN versions of the various models used. The exact width chosen by model selection on the graph datasets is unknown since we report published results. “Linear” means that the chosen downstream classifier is a linear model.
Fixed AWNN Width (Fixed) Width (AWNN)
Mean (Std) Mean (Std) Mean (Std)
DoubleMoon 100.0 (0.0) 100.0 (0.0) 8 8.1 (2.8)
Spiral 99.5 (0.5) 99.8 (0.1) 16 65.9 (8.7)
SpiralHard 98.0 (2.0) 100.0 (0.0) 32 227.4 (32.4)
MNIST 99.6 (0.1) 99.7 (0.0) Linear 19.4 (4.8)
CIFAR10 91.4 (0.2) 91.4 (0.2) Linear 80.1 (12.4)
CIFAR100 66.5 (0.4) 63.1 (4.0) 256 161.9 (57.8)
NCI1 80.0 (1.4) 80.0 (1.1) (96-320) 731.3 (128.2)
REDDIT-B 87.0 (4.4) 90.2 (1.3) (96-320) 793.6 (574.0)
Multi30k ($\downarrow$) 1.43 (0.4) 1.51 (0.2) 24576 123.2 (187.9)
Figure 3: (Left) The learned width adapts to the increasing difficulty of the task, from the DoubleMoon to SpiralHard. (Right) AWNN reaches perfect test accuracy with a comparable amount of epochs on DoubleMoon and Spiral, while it converges faster on SpiralHard.
Figure 4: Training converges to similar widths on SpiralHard for different batch sizes (left) and starting rates $\nu$, but the latter seems to require a bounded nonlinearity such as ReLU6 to converge in a reasonable amount of epochs (right).
Results on graph datasets are interesting in two respects: First, the performance on REDDIT-B is significantly improved by AWNN both in terms of average performance and stability of results; second, the total learned width is significantly higher than those tried in Xu et al. (2019); Errica et al. (2020), meaning that a biased choice of a good range of width has had a profound influence on the estimation of the risk for a specific family of DGN models (i.e., GIN). This result makes it evident that it is important to let the network decide how many neurons are necessary to solve the task. To check that the learned width is indeed conducive to good performances, Appendix D shows what happens when we retrain some fixed baselines using the total width as the width of each layer.
Finally, the results on the Multi30k show that the AWNN Transformer learns to use 200x parameters less than the fixed Transformer for the feed-forward networks, achieving a statistically comparable test loss. This result is appealing when read through the lenses of modern deep learning, as the power required by some neural networks such as Large Language Models (Brown et al., 2020) is so high that cannot be afforded by most institutions, and it demands future investigations.
5.1 Adaptation to Task Difficulty and Convergence
Intuitively, one would expect that AWNN learned larger widths for more difficult tasks. This is indeed what happens on the tabular datasets (and image datasets, see Appendix E) where some tasks are clearly harder than others. Figure 3 (left) shows that, given the same starting width per layer, the learned number of neurons grows according to the task’s difficulty. It is also interesting to observe that the total width for a multi-layer MLP on SpiralHard is significantly lower than that achieved by a single-layer MLP, which is consistent with the circuit complexity theory arguments put forward in Bengio et al. (2006). It also appears that convergence is not affected by the introduction of AWNN, as investigated in Figure 3 (right), which was not obvious considering the parametrization constraints encouraged by the rescaling of neurons’ activations.
Figure 5: It is possible to regularize the width at training time by increasing the magnitude of the loss term $\log\frac{p(\boldsymbol{\nu})}{q(\boldsymbol{\nu})}$. The total width is reduced by more than 50% (left) while preserving accuracy (right). The inset plot refers to the loss term that AWNN tries to maximize.
Figure 6: (Left) Thanks to the soft ordering imposed on the neurons, one can also truncate the neural network after training by simply removing the last neurons. (Right) The distribution of neurons’ activations for all Spiral test samples follows an exponential-like curve.
5.2 Training Stability Analysis
To support our argument that AWNN can reduce the time spent performing hyper-parameter selection, it is instructive to check whether AWNN learns a consistent amount of neurons across different training runs and hyper-parameter choices. Figure 4 reports the impact of the batch size and starting width averaged across the different configurations tried during model selection. Smaller batch sizes cause more instability, but in the long run we observe convergence to a similar width. Convergence with respect to different rates holds, instead, for the bounded ReLU6 activation; Appendix F shows that unbounded activations may cause the network to converge more slowly to the same width, which is in accord with the considerations about counterbalancing the rescaling effect of Section 3.1. Therefore, whenever possible, we recommend using bounded activations.
5.3 Online Network Compression via Regularization
So far, we have used an uninformative prior $p(\lambda)$ over the neural networks’ width. We demonstrate the effect of an informative prior by performing an annealing experiment on the SpiralHard dataset. We set an uninformative $p(\boldsymbol{\theta})$ and ReLU6 nonlinearity. At epoch 1000, we introduce $p(\lambda_{\ell})=\mathcal{N}(\lambda_{\ell};0.05,1)$, and gradually anneal the standard deviation up to $0.1$ at epoch $2500$. Figure 5 shows that the width of the network reduces from approximately 800 neurons to 300 without any test performance de gradation. We hypothesize that the least important neurons mostly carry negligible information, and therefore they can be safely removed without drastic changes in the output of the model. This technique might be particularly useful to compress large models with billions of parameters.
5.4 Post-hoc Trucation achieves a trade-off between performance and compute resources
To further investigate the advantages of imposing a soft ordering of importance among neurons, we show that it is possible to perform hard network’s truncation after training while still controlling the trade-off between performance and network size. Figure 6 shows an example for an MLP on the Spiral dataset, where the range of activation values (Equation 14) computed for all samples follows an exponential curve (right). Therefore, it is expected that removing the last neurons may have a negligible performance impact at the beginning and a drastic one as few neurons remain. This is what happens in the left plot, where we are able to cut an MLP with hidden width 83 by 30% without loss of accuracy, after which a smooth degradation happens. If one accepts such a trade-off, this technique may be used to distill a trained neural network at virtually zero additional cost while reducing the memory requirements.
6 Conclusions
We introduced a new methodology to learn an unbounded width of neural network layers within a single training, by imposing a soft ordering of importance among neurons. Our approach requires very few changes to the architecture, adapts the width to the task’s difficulty, and does not impact negatively convergence, although it introduces some overhead during training. We showed stability convergence to similar widths under bounded activations for different hyper-parameters configurations, advocating for a practical reduction of the width’s search space. A by-product of neurons’ ordering is the ability to easily compress the network during or after training, which is relevant in the context of foundational models trained on large data, which are believed to require billions of parameters. Finally, we have tested AWNN on different models and data domains to prove its broad scope of applicability: a Transformer architecture achieved the same loss with 200x less parameters.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Full derivation of the rescaled weight initialization
This Section derives the formulas for the rescaled weight initialization both in the case of ReLU and activations like tanh.
Background
We can rewrite Equation 14 as
$h^{\ell}_{i}=\alpha^{\ell}_{i}p^{\ell}_{i}$ (16)
$\alpha^{\ell}_{i}=\sigma\left(\sum_{j=1}^{D_{\ell-1}}w^{\ell}_{ij}\underbrace{\alpha^{\ell-1}_{j}p^{\ell-1}_{j}}_{h^{\ell-1}_{j}}\right)$ (17)
where $p^{\ell}_{i}=f_{\ell}(i)$.
As a refresher, the chain rule of calculus states that, given two differentiable functions $g:\mathbb{R}^{D}\rightarrow\mathbb{R}$ and $f=(f_{1},\dots,f_{D}):\mathbb{R}\rightarrow\mathbb{R}^{D}$, their composition $g\circ f:\mathbb{R}\rightarrow\mathbb{R}$ is differentiable and
$\left(g\circ f\right)^{\prime}(t)=\nabla g(f(t))^{T}f^{\prime}(t)$
$\nabla g(f(t))=\left(\frac{\partial g(f_{1}(t))}{\partial f_{1}(t)},\dots,\frac{\partial g(f_{D}(t))}{\partial f_{D}(t)}\right)\in\mathbb{R}^{1\times D}$
$f^{\prime}(t)=\left(f_{1}^{\prime}(t),\dots,f_{D}^{\prime}(t)\right)\in\mathbb{R}^{D\times 1}$
For reasons that will become clear later, we may want to compute the gradient of the loss function with respect to the intermediate activations $\boldsymbol{\alpha}^{\ell}$ at a given layer, that is $\nabla\mathcal{L}(\boldsymbol{\alpha}^{\ell})=\left(\frac{\partial\mathcal{L}(\alpha_{1}^{\ell})}{\partial\alpha_{1}^{\ell}},\dots,\frac{\partial\mathcal{L}(\alpha_{N^{\ell}}^{\ell})}{\partial\alpha_{N^{\ell}}^{\ell}}\right)$. We focus on the $i$-th partial derivative $\frac{\partial\mathcal{L}(\alpha_{i}^{\ell})}{\partial\alpha_{i}^{\ell}}$, where the only variable is $\alpha_{i}^{\ell}$. Then, we view the computation of the loss function starting from $\alpha_{i}^{\ell}$ as a composition of a function $\boldsymbol{\alpha}^{\ell+1}:\mathbb{R}\rightarrow\mathbb{R}^{N^{\ell+1}}=\left(\alpha_{1}^{\ell+1}(\alpha_{i}^{\ell})\dots,\alpha_{{N^{\ell+1}}}^{\ell+1}(\alpha_{i}^{\ell})\right)$ and another function (abusing the notation) $\mathcal{L}:\mathbb{R}^{N^{\ell+1}}\rightarrow\mathbb{R}$ that computes the loss value starting from $\boldsymbol{\alpha}^{\ell+1}$. By the chain rule:
$\underbrace{\frac{\partial\mathcal{L}(\alpha_{i}^{\ell})}{\partial\alpha_{i}^{\ell}}}_{\left(g\circ f\right)^{\prime}(t)}=\underbrace{\left(\frac{\partial\mathcal{L}(\alpha_{1}^{\ell+1})}{\partial\alpha_{1}^{\ell+1}},\dots,\frac{\partial\mathcal{L}(\alpha_{N^{\ell+1}}^{\ell+1})}{\partial\alpha_{N^{\ell+1}}^{\ell+1}}\right)^{T}}_{\nabla g(f(t))^{T}}\underbrace{\left(\frac{\partial\alpha_{1}^{\ell+1}(\alpha_{i}^{\ell})}{\partial\alpha_{i}^{\ell}},\dots,\frac{\partial\alpha_{N^{\ell+1}}^{\ell+1}(\alpha_{i}^{\ell})}{\partial\alpha_{i}^{\ell}}\right)}_{f^{\prime}(t)}$ (18)
Theorem 3.1
Let us consider an MLP with activations as in Equation 17. Let us also assume that the inputs and the parameters have been sampled independently from a Gaussian distribution with zero mean and variance $\sigma^{2}=\mathrm{Var}[w^{\ell}_{ij}]\ \forall i,j,\ell$. At initialization, the variance of the responses $\alpha_{i}^{\ell}$ across layers is constant if, $\forall\ell\in\{1,\dots,L\}$
$\mathrm{Var}[w^{\ell}]=\frac{1}{\sum_{j}^{D_{\ell-1}}\left(p^{\ell+1}_{j}\right)^{2}}\ \ \text{for activation $\sigma$ such that $\sigma^{\prime}(0)\approx 1$}$ (19)
$\mathrm{Var}[w^{\ell}]=\frac{2}{\sum_{j}^{D_{\ell-1}}\left(p^{\ell+1}_{j}\right)^{2}}\ \ \text{for the ReLU activation}.$ (20)
In addition, we provide closed form formulas to to preserve the variance of the gradient across layers.
Proof.
Let us start from the first case of $\sigma^{\prime}(0)\approx 1$. Using the Taylor expansions for the moments of functions of random variables as in Glorot & Bengio (2010)
$\mathrm{Var}[\alpha^{\ell}_{i}]=\mathrm{Var}\left[\sigma\left(\sum_{j=1}^{D_{\ell-1}}w^{\ell}_{ij}\alpha^{\ell-1}_{j}p^{\ell-1}_{j}\right)\right]$ (21)
$\approx\sigma^{\prime}\left(\mathbb{E}\left[\sum_{j=1}^{D_{\ell-1}}w^{\ell}_{ij}\alpha^{\ell-1}_{j}p^{\ell-1}_{j}\right]\right)\mathrm{Var}\left[\sum_{j=1}^{D_{\ell-1}}w^{\ell}_{ij}\alpha^{\ell-1}_{j}p^{\ell-1}_{j}\right]$ (22)
Using the fact that $p_{j}$ is a constant and that $w$ and $\alpha$ are independent from each other
$\mathbb{E}\left[\sum_{j=1}^{D_{\ell-1}}w^{\ell}_{ij}\alpha^{\ell-1}_{j}p^{\ell-1}_{j}\right]=\sum_{j=1}^{D_{\ell-1}}\underbrace{\mathbb{E}[w^{\ell}_{ij}]}_{0}\mathbb{E}[\alpha^{\ell-1}_{j}]p^{\ell-1}_{j}=0.$ (23)
Therefore, recalling that $\sigma^{\prime}(0)\approx 1$
$\sigma^{\prime}\left(\mathbb{E}\left[\sum_{j=1}^{D_{\ell-1}}w^{\ell}_{ij}\alpha^{\ell-1}_{j}p^{\ell-1}_{j}\right]\right)=1.$ (24)
As a result, we arrive at
$\mathrm{Var}[\alpha^{\ell}_{i}]\approx\mathrm{Var}\left[\sum_{j=1}^{D_{\ell-1}}w^{\ell}_{ij}\alpha^{\ell-1}_{j}p^{\ell-1}_{j}\right].$ (25)
Because $w$ and $\alpha$ are independent, they are also uncorrelated and their variances sum. Also, using the fact that $\mathrm{Var}[aX]=a^{2}\mathrm{Var}[X]$ for a constant $a$,
$\mathrm{Var}\left[\sum_{j=1}^{D_{\ell-1}}w^{\ell}_{ij}\alpha^{\ell-1}_{j}p^{\ell-1}_{j}\right]=\sum_{j=1}^{D_{\ell-1}}\mathrm{Var}\left[w^{\ell}_{ij}\alpha^{\ell-1}_{j}p^{\ell-1}_{j}\right]=\sum_{j=1}^{D_{\ell-1}}\mathrm{Var}\left[w^{\ell}_{ij}\alpha^{\ell-1}_{j}\right]\left(p^{\ell-1}_{j}\right)^{2}$ (26)
Finally, because the mean of the independent variables involved is zero by assumption, it holds that $\mathrm{Var}[w^{\ell}_{ij}\alpha^{\ell-1}_{j}]=\mathrm{Var}[w^{\ell}_{ij}]\mathrm{Var}[\alpha^{\ell-1}_{j}]$. We can also abstract from the indexes, since the weight variables are i.i.d. and from that it follows that $\mathrm{Var}[\alpha^{\ell}_{i}]=\mathrm{Var}[\alpha^{\ell}_{j}]\ \forall i,j,\ \ i,j\in\{1,\dots,D_{\ell}\}$, obtaining [^3] It is worth noting that, in standard MLPs, $p_{j}^{\ell}=1$ so we recover the derivation of Glorot & Bengio (2010), since $\sum_{j=1}^{D_{\ell-1}}\left(p^{\ell-1}_{j}\right)^{2}$ would be equal to $D_{\ell-1}$. In Glorot & Bengio (2010) our $D_{\ell-1}$ is denoted as “$n_{i^{\prime}}$” (see Equation 5).
$\mathrm{Var}[\alpha^{\ell}]\approx\mathrm{Var}[w^{\ell}]\mathrm{Var}[\alpha^{\ell-1}]\sum_{j=1}^{D_{\ell-1}}\left(p^{\ell-1}_{j}\right)^{2},$ (27)
noting again that the previous equation does not depend on $i$. We want to impose $\mathrm{Var}[\alpha^{\ell}]\approx\mathrm{Var}[\alpha^{\ell-1}]$, which can be achieved whenever
$\mathrm{Var}[w^{\ell}]\approx\frac{1}{\sum_{j=1}^{D_{\ell-1}}\left(p^{\ell-1}_{j}\right)^{2}}.$ (28)
Condition on the gradients
From a backpropagation perspective, a similar desideratum would be to ensure that $\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{i}^{\ell})}{\partial\alpha_{i}^{\ell}}\right]=\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{i}^{\ell+1})}{\partial\alpha_{i}^{\ell+1}}\right]$ [^4] Note that what we impose is different from Glorot & Bengio (2010), where the specific position $i$ is irrelevant. Here, we are asking that the variance of the gradients for neurons in the same position $i$, but at different layers, stays constant. Alternatively, we could impose an equivalence for all $i\neq i^{\prime}$..
Using Equation 18, and considering as in Glorot & Bengio (2010) that at initialization we are in a linear regime where $\sigma^{\prime}(x)\approx 1$,
$\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{i}^{\ell})}{\partial\alpha_{i}^{\ell}}\right]=\mathrm{Var}\left[\sum_{j=1}^{D_{\ell+1}}\frac{\partial\mathcal{L}(\alpha_{j}^{\ell+1})}{\partial\alpha_{j}^{\ell+1}}\frac{\partial\alpha_{j}^{\ell+1}(\alpha_{i}^{\ell})}{\partial\alpha_{i}^{\ell}}\right]=\mathrm{Var}\left[\sum_{j=1}^{D_{\ell+1}}\frac{\partial\mathcal{L}(\alpha_{j}^{\ell+1})}{\partial\alpha_{j}^{\ell+1}}\frac{\partial\alpha_{j}^{\ell+1}(\alpha_{i}^{\ell})}{\partial\alpha_{i}^{\ell}}\right]$ (29)
$=\mathrm{Var}\left[\sum_{j=1}^{D_{\ell+1}}\frac{\partial\mathcal{L}(\alpha_{j}^{\ell+1})}{\partial\alpha_{j}^{\ell+1}}\underbrace{\sigma^{\prime}\left(\sum_{k=1}^{D_{\ell}}w^{\ell+1}_{jk}\alpha^{\ell}_{k}p^{\ell}_{k}\right)}_{\approx 1}w^{\ell+1}_{ji}p^{\ell}_{i}\right].$ (30)
Using the same arguments as above one can write
$\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{i}^{\ell})}{\partial\alpha_{i}^{\ell}}\right]\approx\mathrm{Var}[w^{\ell+1}]\sum_{j=1}^{D_{\ell+1}}\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{j}^{\ell+1})}{\partial\alpha_{j}^{\ell+1}}p^{\ell}_{i}\right]=\mathrm{Var}[w^{\ell+1}](p^{\ell}_{i})^{2}\sum_{j=1}^{D_{\ell+1}}\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{j}^{\ell+1})}{\partial\alpha_{j}^{\ell+1}}\right].$ (31)
Expanding, we get
$\mathrm{Var}[w^{\ell+1}](p^{\ell}_{i})^{2}\sum_{j=1}^{D_{\ell+1}}\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{j}^{\ell+1})}{\partial\alpha_{j}^{\ell+1}}\right]=(p^{\ell}_{i})^{2}\left(\prod_{i=\ell+1}^{\ell+2}\mathrm{Var}[w^{\ell+1}]\right)\sum_{j=1}^{D_{\ell+1}}(p^{\ell+1}_{j})^{2}\underbrace{\left(\sum_{j^{\prime}=1}^{D_{\ell+2}}\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{j^{\prime}}^{\ell+2})}{\partial\alpha_{j^{\prime}}^{\ell+2}}\right]\right)}_{\text{constant w.r.t.\ }j}$ (32)
$=(p^{\ell}_{i})^{2}\left(\prod_{i=\ell+1}^{\ell+2}\mathrm{Var}[w^{\ell+1}]\right)\left(\sum_{j^{\prime}=1}^{D_{\ell+2}}\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{j^{\prime}}^{\ell+2})}{\partial\alpha_{j^{\prime}}^{\ell+2}}\right]\right)\left(\sum_{j=1}^{D_{\ell+1}}(p^{\ell+1}_{j})^{2}\right)$ (33)
Therefore, we can recursively expand these terms and obtain
$\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{i}^{\ell})}{\partial\alpha_{i}^{\ell}}\right]=(p^{\ell}_{i})^{2}\left(\prod_{k=\ell+1}^{L}\mathrm{Var}[w^{k}]\right)\left(\sum_{j^{\prime}=1}^{D_{L}}\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{j^{\prime}}^{D_{L}})}{\partial\alpha_{j^{\prime}}^{D_{L}}}\right]\right)\left(\prod_{k=\ell+1}^{L-1}\sum_{j_{k}=1}^{D_{k}}(p^{k}_{j_{k}})^{2}\right).$ (34)
In this case, the variance depends on $i$ just for the term $(p^{\ell}_{i})^{2}$. Finally, by imposing
$\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{i}^{\ell})}{\partial\alpha_{i}^{\ell}}\right]=\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{i}^{\ell+1})}{\partial\alpha_{i}^{\ell+1}}\right]$ (35)
and simplifying common terms we obtain
$(p^{\ell}_{i})^{2}\mathrm{Var}[w^{\ell+1}]\left(\sum_{j=1}^{D_{\ell+1}}(p^{\ell+1}_{j})^{2}\right)=(p^{\ell+1}_{i})^{2}$ (36)
$\mathrm{Var}[w^{\ell+1}]=\frac{(p^{\ell+1}_{i})^{2}}{(p^{\ell}_{i})^{2}\sum_{j=1}^{D_{\ell+1}}\left(p^{\ell+1}_{j}\right)^{2}}.$ (37)
Both Equations 28 and 37 are verified when we initialize the distributions $f_{\ell}$ in the same way for all layers, which implies $(p^{\ell}_{i})^{2}=(p^{\ell+1}_{i})^{2}$. Note that without this last requirement, Equation 37 would violate the i.i.d. assumption of the weights.
To show a similar initialization for ReLU activations, we need the following lemma.
Lemma A.1.
Consider the ReLU activation $y=\max(0,x)$ and a symmetric distribution $p(x)$ around zero. Then $\mathbb{E}[y^{2}]=\frac{1}{2}Var[x]$.
Proof.
$\mathbb{E}[y^{2}]=\int_{-\infty}^{{+\infty}}\max(0,x)^{2}p(x)dx=\int_{0}^{{+\infty}}x^{2}p(x)dx=\frac{1}{2}\int_{-\infty}^{{+\infty}}x^{2}p(x)dx$ (38)
Because $p(x)$ is symmetric, $\mathbb{E}[x]=0$. Then
[…]
Using Lemma A.1, we can rewrite Equation 22 as
$\mathrm{Var}[\alpha^{\ell}_{i}]=\mathrm{Var}\left[\sigma\left(\sum_{j=1}^{D_{\ell-1}}w^{\ell}_{ij}\alpha^{\ell-1}_{j}p^{\ell-1}_{j}\right)\right]=\frac{1}{2}\mathrm{Var}\left[\sum_{j=1}^{D_{\ell-1}}w^{\ell}_{ij}\alpha^{\ell-1}_{j}p^{\ell-1}_{j}\right]=\frac{1}{2}\mathrm{Var}[w^{\ell}]\mathrm{Var}[\alpha^{\ell-1}]\sum_{j=1}^{D_{\ell-1}}\left(p^{\ell-1}_{j}\right)^{2}$ (40)
where no approximation has been made. Similar to prior results, it follows that
$\mathrm{Var}[w^{\ell}]=\frac{2}{\sum_{j=1}^{D_{\ell-1}}\left(p^{\ell-1}_{j}\right)^{2}}.$ (41)
From a backpropagation perspective, since $\sigma^{\prime}(x)\not\approx 1$, we use similar arguments as in He et al. (2015): we assume that the pre-activations $\alpha$ have zero mean, then the derivative of the ReLU can have values zero or one with equal probability. One can show, by expanding the definition of variance, that $\mathrm{Var}[\sigma^{\prime}\left(\sum_{k=1}^{D_{\ell}}w^{\ell+1}_{jk}\alpha^{\ell}_{k}p^{\ell}_{k}\right)w^{\ell+1}_{ji}]=\frac{1}{2}\mathrm{Var}[w^{\ell+1}_{ji}]$. As a result, one can re-write Equation 22 as
$\mathrm{Var}\left[\sum_{j=1}^{D_{\ell+1}}\frac{\partial\mathcal{L}(\alpha_{j}^{\ell+1})}{\partial\alpha_{j}^{\ell+1}}\sigma^{\prime}\left(\sum_{k=1}^{D_{\ell}}w^{\ell+1}_{jk}\alpha^{\ell}_{k}p^{\ell}_{k}\right)w^{\ell+1}_{ji}p^{\ell}_{i}\right]$ (42)
$\approx\frac{1}{2}\mathrm{Var}[w^{\ell+1}](p^{\ell}_{i})^{2}\sum_{j=1}^{D_{\ell+1}}\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{j}^{\ell+1})}{\partial\alpha_{j}^{\ell+1}}\right],$ (43)
where we have made a rough assumption of independence between $\sigma^{\prime}\left(\sum_{k=1}^{D_{\ell}}w^{\ell+1}_{jk}\alpha^{\ell}_{k}p^{\ell}_{k}\right)$ and $w^{\ell+1}_{ji}$ since the former term is likely a constant. Following an identical derivation as above, we obtain
$\mathrm{Var}[w^{\ell+1}]=\frac{2(p^{\ell+1}_{i})^{2}}{(p^{\ell}_{i})^{2}\sum_{j=1}^{D_{\ell+1}}\left(p^{\ell+1}_{j}\right)^{2}}$ (44)
which is almost identical to Equation 41 and not dependent on $i$ when we initialize $f_{\ell}$ in the same way for all layers. ∎
Comparison of convergence for different initializations
Below, we provide a comparison of convergence between an AWNN ReLU MLP initialized with the standard kaiming scheme and one initialized according to our theoretical results (“kaiming+”).
Figure 7: (Left) Effect of the rescaled initialization scheme (“kaiming+”) on a ReLU-based MLP, where neurons’ activations are computed using Equation 14. Compared to the standard initialization, the variance of activations agrees with the theoretical result. (Right) Without the rescaled initialization, convergence is hard to attain on the SpiralHard dataset (please refer to the next section for details about the dataset).
Appendix B Dataset Info and Statistics
The synthetic datasets DoubleMoon, Spiral, and SpiralHard are shown in Figure 8. These binary classification datasets have been specifically created to analyze the behavior of AWNN in a controlled scenario where we are sure that the difficulty of the task increases. We provide the code to generate these datasets in the supplementary material.
Figure 8: The DoubleMoon, Spiral, and SpiralHard synthetic datasets are used to test AWNN’s inner workings, and are ordered difficulty. Orange and blue colors denote different classes.
Table 2 reports information on their base statistics and the data splits created to carry out model selection (inner split) and risk assessment (outer split), respectively.
Table 2: Dataset statistics and number of samples in each split are shown.
# Samples # Features # Classes Outer Split (TR/VL/TE) Inner Split (TR/VL)
DoubleMoon 2500 2 2 1750/250/500 1800/200
Spiral 2500 2 2 1750/250/500 1800/200
SpiralHard 5000 2 2 3500/500/1000 3600/400
MNIST 70000 28x28 10 53000/7000/10000 54000/6000
CIFAR10 60000 32x32x3 10 44000/6000/10000 45000/5000
CIFAR100 60000 32x32x3 100 44000/6000/10000 45000/5000
NCI1 4110 37 2 3288/411/411 3329/370
REDDIT-B 2000 1 2 1800/100/100 1710/190
Multi30k 31000 128x50257 Multilabel 29000/1000/1000 29000/1000
Note that for Multi30k, the maximum sequence length is 128 and the number of possible tokens is 50527. The classification loss is computed by comparing the predicted token against the expected one, ignoring any token beyond the length of the actual target. For a detailed description of the graph datasets, the reader can refer to Errica et al. (2020).
Appendix C Hyper-parameters for fixed and AWNN models
For each data domain, we discuss the set of hyper-parameters tried during model selection for the fixed baseline and for AWNN. We try our best to keep architectural choices identical with the exception of the width of the MLPs used by each model. In all experiments, we either run patience-based early stopping (Prechelt, 1998) or simply select the epoch with best validation score.
C.1 Synthetic Tabular Datasets
We report the hyper-parameter configurations tried for the fixed MLP and AWNN in Table 3. In particular, the starting width of AWNN under the quantile function evaluated at $k=0.9$ is approximately $256$ hidden neurons. We also noticed that a learning rate value of $0.1$ makes learning the width unstable, so we did not use it in our experiments. The total number of configurations for the fixed MLP is 180, for AWNN is 18. We also did not impose any prior on an expected size to see if AWNN recovers a similar width compared to the best fixed model (please refer to Table 1) starting from the largest configuration among the fixed models.
Table 3: Hyper-parameter configurations for standard MLP and AWNN versions on tabular datasets.
Hyper-Parameter DoubleMoon Spiral SpiralHard
Batch Size 32 $[32,128]$ (MLP), 128 (AWNN) $[32,128]$ (MLP), 128 (AWNN)
Epochs 500 1000 5000
Hidden Layers 1 1 $[1,2,4]$
Layer Width $[8,16,24,128,256]$
Non-linearity [ReLU, LeakyReLU, ReLU6]
Optimizer Adam, learning rate $\in[0.1,0.01]$ (MLP), $0.01$ (AWNN)
AWNN Specific
Quantile Fun. Threshold $k$ $0.9$
Exponential Distributions Rate $0.01$
$\sigma_{\ell}^{\theta}$ $[1.0,10.0]$
Prior over $\lambda$ Uninformative
C.2 Image Classification Datasets
We train from scratch a ResNet20 and focus our architectural choices on the MLP that performs classification using the flattened representation provided by the previous CNN layers. This ensures that any change is performance is only due to the changes we apply to the MLP. The configurations are shown in Table 4; note that a width of 0 implies a linear classifier instead of a 1 hidden layer MLP. After previous results on tabular datasets showed that a LeakyReLU performs very well, we fixed it in the rest of the experiments.
Table 4: Hyper-parameter configurations for standard MLP and AWNN versions on image classification datasets.
Hyper-Parameter MNIST / CIFAR10 / CIFAR100
Batch Size 128
Epochs 200
Layer Width (classification layer) $[0,32,128,256,512]$
Optimizer SGD, learning rate $0.1$, weight decay $0.0001$, momentum $0.9$
Scheduler Multiply learning rate by $0.1$ at epochs $50,100,150$
AWNN Specific
Quantile Fun. Threshold $k$ $0.9$
Exponential Distributions Rate $0.02$
$\sigma_{\ell}^{\theta}$ $1.0$
Prior over $\lambda$ Uninformative
In this case, we test 5 different configurations of the fixed baseline, whereas we do not have to perform any model selection for AWNN. The rate of the exponential distribution has been chosen to have a starting width of approximately $128$ neurons, which is the median value among the ones tried for the fixed model.
C.3 Text Translation Tasks
Due to the cost of training a Transformer architecture from scratch on a medium size dataset like Multi30k, we use most hyper-parameters’ configurations of the base Transformer in Vaswani et al. (2017), with the difference that AWNN is applied to the MLPs in each encoder and decoder layer. We train an architecture with 6 encoder and 6 decoder layers for 500 epochs, with a patience of 250 epochs applied to the validation loss. We use an Adam optimizer with learning rate 0.01, weight decay 5e-4, and epsilon 5e-9. We also introduce a scheduler that reduces the learning rate by a factor 0.9 when a plateau in the loss is reached. The number of attention heads is set to 8 and the dropout to 0.1. We try different widths for the MLPs, namely [128,256,512,1024,2048]. The embedding size is fixed to 512 and the batch size is 128. The AWNN version starts with an exponential distribution rate of 0.004, corresponding to approximately 512 neurons, and a $\sigma^{\alpha}_{\ell}$=1.
C.4 Graph Classification Tasks
We train the AWNN version of Graph Isomorphism Network, following the setup of Errica et al. (2020) and reusing the published results. We test 12 configurations as shown in the table below. Please note that an extra MLP with 1 hidden layer is always placed before the sequence of graph convolutional layers.
Table 5: Hyper-parameter configurations for AWNN versions on graph classification datasets.
Hyper-Parameter NCI1 / REDDIT-B
Batch Size [32,128]
Epochs 1000
Graph Conv. Layers $[1,2,4]$
Global Pooling [sum, mean]
Optimizer Adam, learning rate $0.01$
AWNN Specific
Patience 500 epochs
Quantile Fun. Threshold $k$ $0.9$
Exponential Distributions Rate $0.02$
$\sigma_{\ell}^{\theta}$ $10.0$
Prior over $\lambda$ Uninformative
Appendix D Re-training a fixed network with the learned width
While on DoubleMoon AWNN learns a similar number of neurons chosen by the model selection procedure and the convergence to the perfect solution over 10 runs seem much more stable, this is not the case on Spiral, SpiralHard and REDDIT-B, where AWNN learns a higher total width over layers. To investigate whether the width learned by AWNN on these datasets is indeed the reason behind the better performances (either a more stable training or better accuracy), we run again the experiments on the baseline networks, this time by fixing the width of each layer to the average width over hidden layers learned by AWNN and reported in Table 1. We call this experiment “Fixed+”.
Spiral SpiralHard REDDIT-B
Fixed 99.5(0.5) 98.0(2.0) 87.0(4.4)
AWNN 99.8(0.1) 100.0(0.0) 90.2(1.3)
Fixed+ 100.0(0.1) 83.2(17.0) 90.7(1.4)
The results suggest that by using more neurons than those selected by grid search – possibly because the validation results were identical between different configurations – it is actually possible to improve both accuracy and stability of results on Spiral and REDDIT-B, while we had a convergence issue on SpiralHard. Notice that the number of neurons learned by AWNN on REDDIT-B was well outside of the range investigated by Xu et al. (2019) and Errica et al. (2020), demonstrating that the practitioner’s bias can play an important role on the final performance.
Appendix E Further qualitative results on image classification datasets
Because it is well known that MNIST, CIFAR10, and CIFAR100 are datasets of increasing difficulty, we attach a figure depicting the learned number of neurons and the convergence of AWNN on these datasets, following the practices highlighted in He et al. (2016) and described in Table 4 to bring the ResNet20 to convergence. Akin to the tabular datasets, we see a clear pattern where the learned number of neurons, averaged over the 10 final training runs, increases together with the task difficulty. The AWNN version of the ResNet20 converges in the same amount of epochs as the fixed one.
Figure 9: We present an analysis similar to Figure 3 but for image classification datasets of increasing difficulty. One can observe the jumps caused by the learning rate’s scheduling strategy.
Appendix F Non-linearity ablation study
Below, we report an ablation study on the impact of non-linear activation functions. We analyze the convergence to the same width across different batch sizes and exponential starting rates $\lambda$ on the SpiralHard dataset.
Figure 10: Impact of different batch sizes and starting exponential rates $\lambda$, organized by type of non-linear activation function.
These qualitative analyses support our claims that using a bounded activation is one way to encourage the network not to counterbalance the learned rescaling of each neuron. In fact, using a ReLU6 shows a distinctive convergence to the same amount of neurons among all configurations tried. The decreasing trend for LeakyReLU and ReLU activations may suggest that these configurations are also converging to a similar value than ReLU6, but they are taking a much longer time.
Appendix G Learned width for each layer on Graph Datasets
Figure 11: Learned neurons per layer (from 1 to 5), averaged over 10 final runs for each of the 10 best configurations selected in the outer folds. The first and second rows refer to NCI1, the third and fourth refer to REDDIT-B. One can observe similar trends in most cases.

### v3

@@
 Abstract
-For almost 70 years, researchers have mostly relied on hyper-parameter tuning to pick the width of neural networks’ layers out of many possible choices.
+For almost 70 years, researchers have mostly relied on hyper-parameter tuning to select the width of neural networks’ layers.
 This paper challenges the status quo by introducing an easy-to-use technique to learn an unbounded width of a neural network’s layer during training.
 The technique does not rely on alternate optimization nor hand-crafted gradient heuristics; rather, it jointly optimizes the width and the parameters of each layer via simple backpropagation.
-We apply the technique to a broad range of data domains such as tables, images, texts, and graphs, showing how the width adapts to the task’s difficulty.
-By imposing a soft ordering of importance among neurons, it is possible to truncate the trained network at virtually zero cost, achieving a smooth trade-off between performance and compute resources in a structured way.
+We apply the technique to a broad range of data domains such as tables, images, text, sequences, and graphs, showing how the width adapts to the task’s difficulty.
+The method imposes a soft ordering of importance among neurons, by which it also is possible to truncate the trained network at virtually zero cost, achieving a smooth trade-off between performance and compute resources in a structured way.
 Alternatively, one can dynamically compress the network with no performance degradation.
-In light of recent foundation models trained on large datasets, believed to require billions of parameters and where hyper-parameter tuning is unfeasible due to huge training costs, our approach stands as a viable alternative for width learning.
-Neural Networks, Learning the Number of Neurons, Adaptive Width Learning, Dynamic Architectures, Information Compression, Variational Inference
+In light of recent foundation models trained on large datasets, believed to require billions of parameters and where hyper-parameter tuning is unfeasible due to humongous training costs, our approach stands as a viable alternative for width learning.
 1 Introduction
-Since the construction of the Mark I Perceptron machine (Rosenblatt, 1958) the effective training of neural networks has remained an open research problem of great academic and practical value.
-The Mark I solved image recognition tasks by exploiting a layer of 512 fixed “association units” that in modern language correspond to the hidden units of a (possibly) Multi-Layer Perceptron (MLP).
-MLPs possess universal approximation capabilities when assuming arbitrary width (Cybenko, 1989) and sigmoidal activations, and their convergence to good solutions was studied, for instance, in Rumelhart et al.
-(1986) where the backpropagation algorithm was described as “simple, easy to implement on parallel hardware”, and improvable by other techniques such as momentum that preserve locality of weight updates.
-Yet, after almost 70 years of progress (LeCun et al., 2015), the vast majority of neural networks, be they shallow or deep, still rely on a fixed choice of the number of neurons in their hidden layers.
-This property is typically treated as an architectural design choice, one of the many hyper-parameters that have to be carefully tuned whenever we approach a new task (Wolpert, 1996).
-The tuning process has many names, such as model selection, hyper-parameter tuning, and cross-validation, and it is associated with non-negligible costs: Different architectural configurations are to be trained until one that performs best on a validation set is selected (Mitchell, 1997).
+Since the construction of the Mark I Perceptron machine  [48] the effective training of neural networks has remained an open research problem of great academic and practical value.
+The Mark I solved image recognition tasks by exploiting a layer of 512 fixed “association units” that in modern language correspond to the hidden units of a Multi-Layer Perceptron (MLP).
+MLPs possess universal approximation capabilities when assuming arbitrary width  [10] and sigmoidal activations, and their convergence to good solutions was studied, for instance, in  Rumelhart et al.
+[50] where the backpropagation algorithm was introduced as “simple, easy to implement on parallel hardware”, and improvable by other techniques such as momentum that preserve locality of weight updates.
+Yet, after almost 70 years of progress  [34], the vast majority of neural networks, be they shallow or deep, still rely on a fixed choice of the number of neurons in their hidden layers.
+The width is typically treated as one of the many hyper-parameters that have to be carefully tuned whenever we approach a new task  [62].
+The tuning process has many names, such as model selection, hyper-parameter tuning, and cross-validation, and it is associated with non-negligible costs: Different architectural configurations are trained until one that performs best on a validation set is selected  [41].
 The configurations’ space grows exponentially in the number of layers, so practitioners often resort to shortcuts such as picking a specific number of hidden units for all layers, which greatly reduces the search space together with the chances of selecting a better architecture for the task.
-Other techniques to tune hyper-parameters include constructive approaches (Fahlman & Lebiere, 1989; Wu et al., 2020), which alternate parameter optimization and creation of new neurons, natural gradient-based heuristics that dynamically modify the network (Mitchell et al., 2023), bi-level optimization (Franceschi et al., 2018), and neural architecture search (White et al., 2023), which often requires separate training runs for each configuration.
-The hyper-parameters’ space exploration problem is exacerbated by the steep increase in size of recent neural architectures for language (Brown et al., 2020) and vision (Zhai et al., 2022), for example, where parameters are in the order of billions to (supposedly) accommodate for a huge dataset.
+Other techniques to tune hyper-parameters include constructive approaches  [17, 64], which alternate parameter optimization and creation of new neurons, natural gradient-based heuristics that dynamically modify the network  [40], bi-level optimization  [19], and neural architecture search  [61], which often requires separate training runs for each configuration.
+The hyper-parameters’ space exploration problem is exacerbated by the steep increase in size of recent neural architectures for language  [9] and vision  [68], for example, where parameters are in the order of billions to (supposedly) accommodate for a huge dataset.
 Training these models requires an amount of time, compute power, and energy that currently makes it unfeasible for most institutions to perform a thorough model selection and find good width parameters; the commonly accepted compromise is to stick to previously successful hyper-parameter choices.
-This may also explain why network pruning (Blalock et al., 2020; Mishra et al., 2021), distillation (Zhang et al., 2019) and quantization (Mishra et al., 2021) techniques have recently been in the spotlight, as they trade-off hardware requirements and performance.
+This may also explain why network pruning  [6, 39], distillation  [69] and quantization  [39] techniques have recently been in the spotlight, as they trade-off hardware requirements and performance.
 This work introduces a simple and easy to use technique to learn the width of each neural network’s layer without imposing upper bounds (we refer to it as unbounded width).
-The width of each layer is dynamically adjusted during backpropagation using the latest deep learning libraries (Paszke et al., 2017), and it only requires a slight modification to the neural activations that does not alter the ability to parallelize computation.
+The width of each layer is dynamically adjusted during backpropagation using the latest deep learning libraries  [44], and it only requires a slight modification to the neural activations that does not alter the ability to parallelize computation.
 The technical strategy is to impose a soft ordering of hidden units by exploiting any monotonically decreasing function with unbounded support on natural numbers, which has a number of benefits.
-For instance, we do not need to fix a maximum number of neurons, which is typical of orthogonal approaches like supernetworks (White et al., 2023).
-As a by-product, we can achieve a straightforward trade-off between parametrization and performance by removing the last rows/columns of weight matrices, which corresponds to ignoring the “least important” neurons in the computation.
-Finally, we break symmetries in the parametrization of neural networks: it is not possible anymore to obtain an equivalent neural network behavior by permuting the weight matrices, which should in principle reduce the “jostling” effect where symmetric parametrizations compete at the beginning of training (Barber, 2012).
-We test the proposed method on an MLP for tabular data, a Convolutional Neural Networks (CNN) (LeCun et al., 1989) for images, a Transformer architecture (Vaswani et al., 2017) for text, and a Deep Graph Network (DGN) (Micheli, 2009; Scarselli et al., 2009) for graphs, to showcase a broad scope of applicability.
-Empirical results suggest, as may be expected, that the width adapts to the task’s difficulty with the same performance of a fixed-width baseline.
-We investigate compressing of the network at training time while preserving accuracy, as well as a post-hoc truncation inducing a controlled trade-off at zero additional cost.
-Ablations suggest that the learned width is not influenced by the starting width (under bounded activations) nor by the batch size, advocating for a potentially large reduction of the hyper-parameter configuration space at model selection time.
+For instance, we do not need to fix a maximum number of neurons, which is typical of orthogonal approaches like supernetworks  [61].
+As a by-product, we can achieve a straightforward trade-off between parametrization and performance by removing the last rows/columns of weight matrices, which correspond to the “least important” neurons in the computation.
+Finally, we break symmetries in the parametrization of neural networks: it is not possible anymore to obtain an equivalent neural network behavior by permuting the weight matrices, which should in principle reduce the “jostling” effect where symmetric parametrizations compete when training starts  [1].
+We test our method on MLPs for tabular data, a Convolutional Neural Network (CNN)  [33] for images, a Transformer architecture  [58] for text, a Recurrent Neural Network (RNN) for sequences, for and a Deep Graph Network (DGN)  [38, 52] for graphs, to showcase a broad scope of applicability.
+Empirical results show, as may be intuitively expected, that the width adapts to the task’s difficulty with the same performance of a fixed-width baseline.
+We also investigate compressing of the network at training time while preserving accuracy, as well as a post-hoc truncation inducing a controlled trade-off at zero additional cost.
+Ablations suggest that the learned width is not influenced by the starting width (under bounded activations) nor by the batch size, advocating for a potentially large reduction of the hyper-parameter configuration space.
 2 Related Work
 Constructive methods dynamically learn the width of neural networks and are related in spirit to this work.
-The cascade correlation algorithm (Fahlman & Lebiere, 1989) alternates standard training with the creation of a new hidden unit minimizing the neural network’s residual error.
-Similarly, the firefly network descent (Wu et al., 2020) grows the width and depth of a network every $N$ training epochs via gradient descent on a dedicated loss.
+The cascade correlation algorithm  [17] alternates standard training with the creation of a new hidden unit minimizing the neural network’s residual error.
+Similarly, the firefly network descent  [64] grows the width and depth of a network every $N$ training epochs via gradient descent on a dedicated loss.
 Yoon et al.
-(2018) propose an ad-hoc algorithm for lifelong learning that grows the network by splitting and duplicating units to learn new tasks.
+[66] propose an ad-hoc algorithm for lifelong learning that grows the network by splitting and duplicating units to learn new tasks.
 Wu et al.
-(2019) alternate training the network and then splitting existing neurons into offspring with equal weights.
+[63] alternate training the network and then splitting existing neurons into offspring with equal weights.
 These works mostly focus on growing the neural network;  Mitchell et al.
-(2023) propose natural gradient-based heuristics to grow/shrink layers and hidden units of MLPs and CNNs.
+[40] propose natural gradient-based heuristics to grow/shrink layers and hidden units of MLPs and CNNs.
 The main difference from our work is that we grow and shrink the network by simply computing the gradient of the loss, without relying on human-defined heuristics.
-The unbounded depth network of Nazaret & Blei (2022), from which we draw inspiration, learns the number of layers of neural networks.
+The unbounded depth network of  Nazaret and Blei [42], from which we draw inspiration, learns the number of layers of neural networks.
 Compared to that work, we focus our attention to the number of neurons, modifying the internals of the architecture rather than instantiating a multi-output one.
-Finally, we mention Bayesian nonparametrics approaches (Orbanz & Teh, 2010) that learn a potentially infinite number of clusters in an unsupervised fashion.
+Finally, we mention Bayesian nonparametrics approaches  [43] that learn a potentially infinite number of clusters in an unsupervised fashion, as well as dynamic neural networks  [24] that condition the architecture on input properties.
 Orthogonal Methods
-Neural Architecture Search (NAS) is an automated process that designs neural networks for a given task (Elsken et al., 2019b; White et al., 2023) and has been applied to different contexts (Zoph et al., 2018; Liu et al., 2019; So et al., 2019).
-Typically, neural network elements are added, removed, or modified based on validation performance (Elsken et al., 2019a; White et al., 2021, 2023), by means of reinforcement learning (Zoph & Le, 2016), evolutionary algorithms (Real et al., 2019), and gradient-based approaches (Liu et al., 2019).
-Typical NAS methods require enormous computational resources, sometimes reaching thousands of GPU days (Zoph & Le, 2016), due to the retraining of each new configuration.
-While recent advances on one-shot NAS models (Brock et al., 2018; Pham et al., 2018; Bender et al., 2018) have drastically reduced the computational costs, they mostly focus on CNNs, assume a bounded search space, and do not learn the width.
+Neural Architecture Search (NAS) is an automated process that designs neural networks for a given task  [15, 61] and has been applied to different contexts  [71, 35, 53].
+Typically, neural network elements are added, removed, or modified based on validation performance  [14, 60, 61], by means of reinforcement learning  [70], evolutionary algorithms  [47], and gradient-based approaches  [35].
+Typical NAS methods require enormous computational resources, sometimes reaching thousands of GPU days  [70], due to the retraining of each new configuration.
+While recent advances on one-shot NAS models  [8, 45, 2, 4, 55, 54] have drastically reduced the computational costs, they mostly focus on CNNs, assume a bounded search space, and do not learn the width.
 As such, NAS methods are complementary to our approach.
-Bi-level optimization algorithms have also been used for hyper-parameter tuning (Franceschi et al., 2018), where hyper-parameters are the variables of the outer objective and the model parameters those of the inner objective.
-The solution sets of the inner problem are usually not available in closed form, which has been partly addressed by repeated application of (stochastic) gradient descent (Domke, 2012; Maclaurin et al., 2015; Franceschi et al., 2017).
+Bi-level optimization algorithms have also been used for hyper-parameter tuning  [19], where hyper-parameters are the variables of the outer objective and the model parameters those of the inner objective.
+The solution sets of the inner problem are usually not available in closed form, which has been partly addressed by repeated application of (stochastic) gradient descent  [11, 36, 18].
 These methods are restricted to continuous hyper-parameters’ optimization, and cannot therefore be applied to width optimization.
-Finally, pruning (Blalock et al., 2020) and distillation (Hinton et al., 2015) are two methods that (only) reduce the size of neural networks by trading-off performances; the former deletes neural connections (Mishra et al., 2021) or entire neurons (Valerio et al., 2022; Dufort-Labbé et al., 2024), the latter trains a smaller network (student) to mimic a larger one (teacher) (Gou et al., 2021).
-Compared to most pruning approaches, our work can delete connections and reduce the model’s memory; compared to distillation, we do not necessarily need a new training.
+Finally, pruning  [6] and distillation  [29] are two methods that reduce the size of neural networks by trading-off performances; the former deletes neural connections  [39] or entire neurons  [57, 12], the latter trains a smaller network (student) to mimic a larger one (teacher)  [22].
+In particular, dynamic pruning techniques can compress the network at training time  [23], by applying hard or soft masks  [28]; for a comprehensive survey on pruning, please refer to  [27].
+Compared to most pruning approaches, our work can delete connections and reduce the model’s memory, but also grow it indefinitely; compared to distillation, we do not necessarily need a new training to compress the network.
 These techniques, however, can be easily combined with our approach.
 3 Adaptive Width Learning
+We now introduce Adaptive Width Neural Networks (AWNN), a probabilistic framework that maximizes a simple variational objective via backpropagation over a neural network’s parameters.
+We are given a dataset of $N$ i.i.d.
+samples, with input $x\in\mathbb{R}^{F},F\in\mathbb{N}^{+}$ and target $y$ whose domain depends on whether the task is regression or classification.
+For samples $X\in\mathbb{R}^{N\times F}$ and targets $Y$, the learning objective is to maximize the likelihood
+$\log p(Y|X)=\log\prod_{i=1}^{N}p(y_{i}|x_{i})=\sum_{i=1}^{N}\log p(y_{i}|x_{i})$ (1)
 […]
-Figure 1: (Left) We depict the graphical model of AWNN, where observable random variables are dark and latent ones are white.
+$p(Y,\boldsymbol{\lambda},\boldsymbol{\theta}|X)=\prod_{i=1}^{N}p(y_{i},\boldsymbol{\lambda},\boldsymbol{\theta}|x_{i})\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ p(y_{i},\boldsymbol{\lambda},\boldsymbol{\theta}|x_{i})=p(y_{i}|\boldsymbol{\lambda},\boldsymbol{\theta},x_{i})p(\boldsymbol{\lambda})p(\boldsymbol{\theta})$ (2)
+$p(\boldsymbol{\lambda})=\prod^{L}_{\ell=1}p(\lambda_{\ell})=\prod^{L}_{\ell=1}\mathcal{N}(\lambda_{\ell};\mu^{\lambda}_{\ell},\sigma^{\lambda}_{\ell})\ \ \ \ \ \ p(\boldsymbol{\theta})=\prod^{L}_{\ell=1}\prod^{\infty}_{n=1}p(\theta_{\ell n})\text{=}\prod^{L}_{\ell=1}\prod^{\infty}_{n=1}\mathcal{N}(\theta_{\ell n};\mathbf{0},\text{diag}(\sigma^{\theta}_{\ell}))$ (3)
+[…]
+[…]
+Figure 1: (Left) The graphical model of AWNN, with dark observable random variables and white latent ones.
 (Middle) The distribution $f_{\ell}$ over hidden units’ importance at layer $\ell$ is parametrized by $\lambda_{\ell}$.
@@
 (Right) The hidden units’ activations at layer $\ell$ are rescaled by their importance.
-We introduce a general probabilistic framework, called Adaptive Width Neural Networks (AWNN) for convenience, showing how (ultimately simple) design choices arise from a variational inference treatment of a graphical model.
-We are given a dataset of $N$ i.i.d.
-samples $(x_{i},y_{i})$, with input $x_{i}\in\mathbb{R}^{F},F\in\mathbb{N}^{+}$ and target $y_{i}$ whose domain depends on whether the task is regression or classification.
-The learning objective is to maximize
-$\log\prod_{i=1}^{N}p(y_{i}|x_{i})=\sum_{i=1}^{N}\log p(y_{i}|x_{i})$ (1)
 […]
-$p(y_{i},\boldsymbol{\lambda},\boldsymbol{\theta}|x_{i})=p(y_{i}|\boldsymbol{\lambda},\boldsymbol{\theta},x_{i})p(\boldsymbol{\lambda})p(\boldsymbol{\theta})$ (2)
-$p(\boldsymbol{\lambda})=\prod^{L}_{\ell=1}p(\lambda_{\ell})=\prod^{L}_{\ell=1}\mathcal{N}(\lambda_{\ell};\mu^{\lambda}_{\ell},\sigma^{\lambda}_{\ell})$ (3)
-$p(\boldsymbol{\theta})=\prod^{L}_{\ell=1}\prod^{\infty}_{n=1}p(\theta_{\ell n})\text{=}\prod^{L}_{\ell=1}\prod^{\infty}_{n=1}\mathcal{N}(\theta_{\ell n};\mathbf{0},\text{diag}(\sigma^{\theta}_{\ell}))$ (4)
-[…]
-Here, $\sigma^{\theta}_{\ell},\mu^{\lambda}_{\ell},\sigma^{\lambda}_{\ell}$ are hyper-parameters.
-The MLP is parametrized by realizations $\boldsymbol{\lambda},\boldsymbol{\theta}$, so it relies on a finite number of neurons and outputs either class probabilities (classification) or the mean of a Gaussian distribution (regression) to parametrize $p(y_{i}|\boldsymbol{\lambda},\boldsymbol{\theta},x_{i})$ depending on the task.
-Maximizing Equation (1), however, requires computing the above integral, which is intractable.
-Therefore, we turn to mean-field variational inference (Jordan et al., 1999; Blei et al., 2017) to maximize an expected lower bound (ELBO) instead.
-This requires to define a distribution over the latent variables $q(\boldsymbol{\lambda},\boldsymbol{\theta})$ and re-phrase the objective as:
-$\sum_{i=1}^{N}\log p(y_{i}|x_{i})\geq\sum_{i=1}^{N}\mathbb{E}_{q(\boldsymbol{\lambda},\boldsymbol{\theta})}\left[\log\frac{p(y_{i},\boldsymbol{\lambda},\boldsymbol{\theta}|x_{i})}{q(\boldsymbol{\lambda},\boldsymbol{\theta})}\right],$ (6)
-where $q(\boldsymbol{\lambda},\boldsymbol{\theta})$ is parametrized by learnable variational parameters. We factorize the variational distribution into:
-$q(\boldsymbol{\lambda},\boldsymbol{\theta})=q(\boldsymbol{\lambda})q(\boldsymbol{\theta}|\boldsymbol{\lambda})$ (7)
-$q(\boldsymbol{\lambda})=\prod_{\ell=1}^{L}q(\lambda_{\ell})=\prod_{\ell=1}^{L}\mathcal{N}(\lambda_{\ell};\nu_{\ell},1)$ (8)
-$q(\boldsymbol{\theta}|\boldsymbol{\lambda})=\prod_{\ell=1}^{L}\prod_{n=1}^{D_{\ell}}q(\theta_{\ell n})\prod_{D_{\ell}+1}^{\infty}p(\theta_{\ell n})$ (9)
-$q(\theta_{\ell n})=\mathcal{N}(\theta_{\ell n};\text{diag}(\rho_{\ell n}),\mathbf{I}).$ (10)
-$D_{\ell}=\text{quantile function of }f_{\ell}(\cdot;\lambda_{\ell})\text{ evaluated at }k$ (11)
-The value $k$ is a hyper-parameter, $\nu_{\ell},\rho_{\ell n}$ are variational parameters and, as before, we define $\boldsymbol{\rho}_{\ell}=\left\{\rho_{\ell n}\right\}_{n=1}^{D_{\ell}}$, $\boldsymbol{\rho}=\left\{\boldsymbol{\rho}_{\ell}\right\}_{\ell=1}^{L}$ and $\boldsymbol{\nu}=\left\{\nu_{\ell}\right\}_{\ell=1}^{L}$.
-Note that the set of variational parameters is finite.
-The truncated width $D_{\ell}$, that is the finite number of neurons at layer $\ell$, is computed as the quantile function evaluated at $k$ of a distribution [^2] General functions are allowed if a threshold can be computed.
-$f_{\ell}$ with infinite support over $\mathbb{N}^{+}$, parametrized by $\lambda_{\ell}$.
-W.l.o.g., we implement $f_{\ell}$ as a discretized exponential distribution, following the discretization strategy of Roy (2003): For every natural $x$, the discretized distribution relies on the cumulative distribution function (c.d.f.) of the exponential:
+$\log p(Y|X)\geq\mathbb{E}_{q(\boldsymbol{\lambda},\boldsymbol{\theta})}\left[\log\frac{p(Y,\boldsymbol{\lambda},\boldsymbol{\theta}|X)}{q(\boldsymbol{\lambda},\boldsymbol{\theta})}\right],$ (5)
+where $q(\boldsymbol{\lambda},\boldsymbol{\theta})$ is parametrized by learnable variational parameters.
+Before continuing, we define the truncated width $D_{\ell}$, that is the finite number of neurons at layer $\ell$, as the quantile function evaluated at $k$, with $k$ a hyper-parameter, of a distribution [^1] General functions are allowed if a threshold can be computed.
+$f_{\ell}$ with infinite support over $\mathbb{N}^{+}$ and parametrized by $\lambda_{\ell}$; Appendix A and B provide formal requirements about $f_{\ell}$.
+In other words, we find the integer such that the cumulative mass function of $f_{\ell}$ takes value $k$, and that integer $D_{\ell}$ is the truncated width of layer $\ell$.
+W.l.o.g., we implement $f_{\ell}$ as a discretized exponential distribution adhering to Def.
+A.2, following the discretization strategy of  Roy [49]: For every natural $x$, the discretized distribution relies on the exponential’s cumulative distribution function:
 $f_{\ell}(x;\lambda_{\ell})=(1-e^{\lambda_{\ell}(x+1)})-(1-e^{\lambda_{\ell}(x)}).$ (6)
 We choose the exponential because it is a monotonically decreasing function and allows us to impose an ordering of importance among neurons, as detailed in Section 3.1.
-By expanding Equation 6 using the above definitions and approximating the expectations at the first order, i.e., $\mathbb{E}_{q(\boldsymbol{\lambda})}[f(\boldsymbol{\lambda})]$=$f(\boldsymbol{\nu})$ and $\mathbb{E}_{q(\boldsymbol{\theta}|\boldsymbol{\lambda})}[f(\boldsymbol{\theta})]=f(\boldsymbol{\rho})$ as in Nazaret & Blei (2022), we obtain the final form of the objective:
-$\sum_{\ell}^{L}\log\frac{p(\nu_{\ell};\mu_{\ell}^{\lambda},\sigma_{\ell}^{\lambda})}{q(\nu_{\ell};\nu_{\ell})}+\sum_{\ell}^{L}\sum_{n=1}^{D_{\ell}}\log\frac{p(\rho_{\ell n};\sigma_{\ell}^{\theta})}{q(\rho_{\ell n};\rho_{\ell n})}+$
-$+\sum_{i=1}^{N}\log p(y_{i}|\boldsymbol{\lambda}\text{=}\boldsymbol{\nu},\boldsymbol{\theta}\text{=}\boldsymbol{\rho},x_{i}),$ (13)
+Then, we can factorize the variational distribution $q(\boldsymbol{\lambda},\boldsymbol{\theta})$ into:
+$q(\boldsymbol{\lambda},\boldsymbol{\theta})=q(\boldsymbol{\lambda})q(\boldsymbol{\theta}|\boldsymbol{\lambda})$ $q(\boldsymbol{\lambda})=\prod_{\ell=1}^{L}q(\lambda_{\ell})=\prod_{\ell=1}^{L}\mathcal{N}(\lambda_{\ell};\nu_{\ell},1)$ (7)
+$q(\boldsymbol{\theta}|\boldsymbol{\lambda})=\prod_{\ell=1}^{L}\prod_{n=1}^{D_{\ell}}q(\theta_{\ell n})\prod_{D_{\ell}+1}^{\infty}p(\theta_{\ell n})$ $q(\theta_{\ell n})=\mathcal{N}(\theta_{\ell n};\text{diag}(\rho_{\ell n}),\mathbf{I}).$ (8)
+Here, $\nu_{\ell},\rho_{\ell n}$ are learnable variational parameters and, as before, we define $\boldsymbol{\rho}_{\ell}=\left\{\rho_{\ell n}\right\}_{n=1}^{D_{\ell}}$, $\boldsymbol{\rho}=\left\{\boldsymbol{\rho}_{\ell}\right\}_{\ell=1}^{L}$ and $\boldsymbol{\nu}=\left\{\nu_{\ell}\right\}_{\ell=1}^{L}$.
+Note that the set of variational parameters is finite as it depends on $D_{\ell}$.
+By expanding Equation 5 using the above definitions and approximating the expectations at the first order, i.e., $\mathbb{E}_{q(\boldsymbol{\lambda})}[f(\boldsymbol{\lambda})]$=$f(\boldsymbol{\nu})$ and $\mathbb{E}_{q(\boldsymbol{\theta}|\boldsymbol{\lambda})}[f(\boldsymbol{\theta})]=f(\boldsymbol{\rho})$ as in  Nazaret and Blei [42], we obtain the final form of the ELBO [^2] The higher $k$ the better the ELBO approximation.
+(the full derivation is in Appendix C):
+$\sum_{\ell}^{L}\log\frac{p(\nu_{\ell};\mu_{\ell}^{\lambda},\sigma_{\ell}^{\lambda})}{q(\nu_{\ell};\nu_{\ell})}+\sum_{\ell}^{L}\sum_{n=1}^{D_{\ell}}\log\frac{p(\rho_{\ell n};\sigma_{\ell}^{\theta})}{q(\rho_{\ell n};\rho_{\ell n})}+\sum_{i=1}^{N}\log p(y_{i}|\boldsymbol{\lambda}\text{=}\boldsymbol{\nu},\boldsymbol{\theta}\text{=}\boldsymbol{\rho},x_{i}),$ (9)
 where distributions’ parameters are made explicit to distinguish them.
-The first two terms in the loss regularize the width of the layers and the magnitude of the parameters, respectively, whereas the third is is the predictive loss.
-In practice, the finite variational parameters $\boldsymbol{\nu},\boldsymbol{\rho}$ are those used by the neural network in place of $\boldsymbol{\lambda},\boldsymbol{\theta}$, which enables easy optimization via backpropagation.
-Maximizing Equation (13) will change each variational parameter $\nu_{\ell}$, which in turn will change the value of $D_{\ell}$ during training.
-If $D_{\ell}$ increases we initialize new neurons and draw their weights from a standard normal distribution, otherwise we remove the excess ones.
+The first two (optional) terms in the ELBO regularize the width of the layers and the magnitude of the parameters, respectively, whereas the third term accounts for the predictive performance.
+In practice, the finite variational parameters $\boldsymbol{\nu},\boldsymbol{\rho}$ are those used to parametrize the neural network rather than sampling $\boldsymbol{\lambda},\boldsymbol{\theta}$, which enables easy optimization via backpropagation.
+Maximizing Equation (9) will update each variational parameter $\nu_{\ell}$, which in turn will change the value of $D_{\ell}$ during training.
+If $D_{\ell}$ increases we initialize new neurons and draw their weights from a standard normal distribution, otherwise we discard the weights of the extra neurons.
 When implementing mini-batch training, the predictive loss needs to be rescaled by $N/M$, where $M$ is the mini-batch size.
-From a Bayesian perspective, this means regularizers weigh less if we have more data.
-Compared to a fixed-width network with weight decay, we need to choose the values of $\mu^{\lambda}_{\ell},\sigma^{\lambda}_{\ell}$, as well as initialize $\nu_{\ell}$ (shared across layers as there is no particular reason to do otherwise).
-Therefore, we have two more hyper-parameters compared to the fixed-width network, but we make some considerations: i) it is always possible to use an uninformative prior over $\lambda_{\ell}$, removing the extra hyper-parameters letting the model freely adapt the width of each layer (as is typical of frequentist approaches); ii) the choice of higher level of hyper-parameters is known to be less stringent than that of hyper-parameters themselves (Goel & Degroot, 1981; Bernardo & Smith, 2009), so we do not need to explore many values of $\mu_{\ell}^{\lambda}$ and $\sigma_{\ell}^{\lambda}$; iii) our experiments suggest that AWNN converges to similar widths regardless of the starting point $\nu_{\ell}$, so that we may just need to perform model selection over one/two sensible initial values.
+From a Bayesian perspective, this is necessary as regularizers should weigh less if we have more data.
+Finally, note that the first-order approximation leaves $\boldsymbol{\nu}$ and $\boldsymbol{\rho}$ completely free to change if not for the regularizers, that is, there is potentially no restriction imposed by the Gaussian assumptions.
+Compared to a fixed-width network with weight decay, we need to choose the values of $\mu^{\lambda}_{\ell},\sigma^{\lambda}_{\ell}$, as well as initialize the learnable $\nu_{\ell}$.
+The latter can be initially set to same value across layers since they can freely adapt later, or it can be sampled from the prior $p(\lambda_{\ell})$.
+Therefore, we have two more hyper-parameters compared to the fixed-width network, but we make some considerations: i) it is always possible to use an uninformative prior over $\lambda_{\ell}$, removing the extra hyper-parameters letting the model freely adapt the width of each layer (as is typical of frequentist approaches); ii) the choice of higher level of hyper-parameters is known to be less stringent than that of hyper-parameters themselves  [21, 5], so we do not need to explore many values of $\mu_{\ell}^{\lambda}$ and $\sigma_{\ell}^{\lambda}$; iii) our experiments suggest that AWNN can converge to similar widths regardless of the starting point $\nu_{\ell}$, so that we may just need to perform model selection over one/two sensible initial values.
 3.1 Imposing a Soft Ordering on Neurons’ Importance
-Now that the learning objective has been formalized, the missing ingredient is the definition of the neural network $p(y_{i}|\boldsymbol{\lambda}\text{=}\boldsymbol{\nu},\boldsymbol{\theta}\text{=}\boldsymbol{\rho},x_{i})$.
-Compared to a standard MLP, we need to make use of the variational parameters $\boldsymbol{\nu}$ that affect the truncation width at each hidden layer, whereas $\boldsymbol{\rho}$ corresponds to the the usual MLP weights.
-We choose a monotonically decreasing function $f_{\ell}$ for the simple reason that when a new neuron is added, its relative importance is low and will not drastically impact the network.
-As a result, we are imposing a soft ordering of importance among neurons.
-We modify the classical activation $h^{\ell}_{j}$ of a hidden neuron $j$ at layer $\ell$ as
+Now that the learning objective has been formalized, the missing ingredient is the definition of the neural network $p(y_{i}|\boldsymbol{\lambda}\text{=}\boldsymbol{\nu},\boldsymbol{\theta}\text{=}\boldsymbol{\rho},x_{i})$ of Equation 4 as a modified MLP.
+Compared to a standard MLP, we need to make use of the variational parameters $\boldsymbol{\nu}$ that affect the truncation width at each hidden layer, whereas $\boldsymbol{\rho}$ are the weights.
+We chose a monotonically decreasing function $f_{\ell}$, thus when a new neuron is added its relative importance is low and will not drastically impact the network output and hidden activations.
+In other words, we impose a soft ordering of importance among neurons.
+We simply modify the classical activation $h^{\ell}_{j}$ of a hidden neuron $j$ at layer $\ell$ as
 $h^{\ell}_{j}=\sigma\left(\sum_{k=1}^{D_{\ell-1}}w^{\ell}_{jk}h^{\ell-1}_{k}\right)f_{\ell}(j;\nu_{\ell}),$ (10)
 where $D_{\ell-1}$ is the truncated width of the previous layer, $\sigma$ is a non-linear activation function and $w^{\ell}_{jk}\in\rho_{\ell j}$.
-That is, we rescale the activation of each neuron $k$ by its “importance” $f_{\ell}(j;\nu_{\ell})$.
-Note that the bias parameter is taken into account by concatenating a dummy value 1 to $h^{\ell-1}_{k}$.
+That is, we rescale the activation of each neuron $k$ by its “importance" $f_{\ell}(j;\nu_{\ell})$.
+Note that the bias parameter can be part of the weight vector as usual.
 It is easy to see that, in theory, the optimization algorithm could rescale the weights of the next layer by a factor $1/f_{\ell}(j;\nu_{\ell})$ to compensate for the term $f_{\ell}(j;\nu_{\ell})$.
-This may lead to a degenerate situation, which is actually quite common in classical MLPs, where the activations of the first neurons are small relative to the others, thus breaking the soft-ordering and wasting neurons.
+This could lead to a degenerate situation where the activations of the first neurons are small relative to the others, thus breaking the soft-ordering and potentially wasting neurons.
 There are two strategies to address this undesirable effect.
 The first is to regularize the magnitude of the weights thanks to the prior $p(\theta_{\ell+1n})$, so that it may be difficult to compensate for the least important neurons that have a high $1/f_{\ell}(j)$.
-The second and less obvious strategy is to prevent the units’ activations of the current layer to compensate for high values by bounding their range.
-We apply both strategies to our experiments.
+The second and less obvious strategy is to prevent the units’ activations of the current layer to compensate for high values by bounding their range, e.g., using a ReLU6 activation  [51].
+We apply both strategies to our experiments, although we noted that they do not seem strictly necessary in practice.
 3.2 Rescaled Weight Initialization for Deep AWNN
-Rescaling the activations of hidden units using Equation 14 causes activations of deeper layers to quickly decay to zero, as shown in Figure 2 for a 5-layer AWNN MLP with ReLU nonlinearity initialized using the well known kaiming scheme (He et al., 2015).
+Rescaling the activations of hidden units using Equation 10 causes activations of deeper layers to quickly decay to zero for an AWNN MLP with ReLU nonlinearity initialized using the well known Kaiming scheme  [25].
 This affects convergence since gradients get close to zero and it becomes impractical to train deep AWNN MLPs.
@@
 Theorem 3.1.
-Let us consider an MLP with activations as in Equation 14 and ReLU nonlinearity. At initialization, given $\alpha_{j}^{\ell}=\sigma\left(\sum_{k=1}^{D_{\ell-1}}w^{\ell}_{jk}h^{\ell-1}_{k}\right)$
-$\mathrm{Var}[w_{j*}^{\ell}]=\frac{2}{\sum_{j=1}^{D_{\ell-1}}f^{2}_{\ell}(j)}\Rightarrow\mathrm{Var}[\alpha_{j}^{\ell}]\approx\mathrm{Var}[\alpha_{j}^{\ell-1}]$ (15)
+Consider an MLP with activations as in Equation 10 and ReLU nonlinearity.
+At initialization, given $\alpha_{j}^{\ell}=\sigma\left(\sum_{k=1}^{D_{\ell-1}}w^{\ell}_{jk}h^{\ell-1}_{k}\right),\mathrm{Var}[w_{j*}^{\ell}]=\frac{2}{\sum_{j=1}^{D_{\ell-1}}f^{2}_{\ell}(j)}\Rightarrow\mathrm{Var}[\alpha_{j}^{\ell}]\approx\mathrm{Var}[\alpha_{j}^{\ell-1}]$.
 Proof.
-See Appendix A for an extended proof. ∎
+See Appendix E for an extended proof. ∎
 The extended proof shows there is a connection between the variance of activations and the variance of gradients.
-In particular, the sufficient conditions over $\mathrm{Var}[w_{j*}^{\ell}]$ are identical if one initializes all $\nu_{\ell}$ in the same way for all layers.
+In particular, the sufficient conditions over $\mathrm{Var}[w_{j*}^{\ell}]$ are identical if one initializes $\nu_{\ell}$ in the same way for all layers.
 Therefore, if we initialize weights from a Gaussian distribution with standard deviation $\frac{\sqrt{2}}{\sqrt{\sum_{j=1}^{D_{\ell-1}}f^{2}_{\ell}(j)}}$, we guarantee that at initialization the variance of the deep network’s gradients will be constant.
-The effect of the new initialization (dubbed “kaiming+”) can be seen in Figure 2, where the distribution of activation values at initialization does not collapse in subsequent layers.
-This change drastically impacts overall convergence on the SpiralHard synthetic dataset (described in Section 4), where it appears it would be otherwise hard to converge using a standard kaiming initialization (please refer to Appendix A for a visual example).
-Figure 2: Effect of the rescaled initialization scheme (“kaiming+”) on a ReLU-based AWNN MLP, where neurons’ activations are computed using Equation 14.
-Compared to the standard initialization, the variance of activations agrees with the theoretical result.
-Algorithm 1 summarizes the main changes to the training procedure, namely the new initialization and the update of the model’s truncated width at each training step.
-3.3 Future Directions and Limitations
-MLPs are ubiquitous in modern deep architectures.
-They are used at the very end of CNNs, in each encoder and decoder layer of Transformer, and they process messages coming from neighbors in DGNs.
-Our experiments focus on MLPs to showcase AWNN’s broad applicability, but there are many other scenarios where one can apply AWNN’s principles.
-For instance, one could impose a soft ordering of importance on CNNs’ filters at each layer, therefore learning the number of filters during training.
-Alternatively, one could investigate the application of AWNN to timeseries, since recurrent networks (Hochreiter & Schmidhuber, 1997) can be seen as a special case of an MLP with weight sharing.
-From a more theoretical perspective, we believe one could draw connections between our technique and the Information Bottleneck principle (Tishby et al., 2000), which seeks maximally representative (i.e., performance) and compact representations (e.g., width).
-Finally, the analysis of different functions $f$ to soft-order neurons, for instance a power-law distribution or a sigmoid-like function, may have a different impact on convergence and performance and could be empirically investigated.
-The current limitation of AWNN is the relative overhead to change the network at training time.
-A minimal and naïve implementation has up to 10x cost compared to a fixed network (this relative overhead may reduce for more compute-bound tasks).
-Besides optimizing our code, other optimizations are conceivable: i) by updating the width every $M$ minibatches rather than at every step; ii) by allocating a larger network and dynamically grow it only if needed; iii) by creating ad-hoc support of AWNN on deep learning libraries like Pytorch.
-In any case, this overhead is relatively small compared to the exponential size of the hyper-parameter space when cross-validating the number of neurons of a multi-layered architecture.
+The effect of the new initialization (dubbed “Kaiming+”) can be seen in Figure
 Algorithm 1 AWNN Training Procedure
@@
 8:       $\mathcal{M}\leftarrow\text{backpropagation}(\mathcal{M},\text{loss})$
-9:    end for
-10:  end for
-11:  function $\mathcal{M}$):
+9:  function $\mathcal{M}$):
 […]
-13:       $D_{\ell}\leftarrow\text{quantile function of }f_{\ell}(\cdot;\nu_{\ell})\text{\ evaluated at }k$
-14:       Use $D_{\ell}$ to update $\boldsymbol{\rho}_{\ell},\boldsymbol{\rho}_{\ell+1}$ // add/remove neurons
-15:    end for
-16:  end function
+11:       $D_{\ell}\leftarrow\text{quantile function of }f_{\ell}(\cdot;\nu_{\ell})\text{\ evaluated at }k$
+12:       Use $D_{\ell}$ to update $\boldsymbol{\rho}_{\ell},\boldsymbol{\rho}_{\ell+1}$ // add/remove neurons
+7, where the distribution of activation values at initialization does not collapse in subsequent layers.
+This change drastically impacts overall convergence on the SpiralHard synthetic dataset (described in Section 4), where it appears it would be otherwise hard to converge using a standard Kaiming initialization.
+Algorithm 1 summarizes the main changes to the training procedure, namely the new initialization and the update of the model’s truncated width at each training step.
+3.3 Future Directions and Limitations
+MLPs are ubiquitous in modern deep architectures.
+They are used at the very end of CNNs, in each Transformer layer, and they process messages coming from neighbors in DGNs.
+Our experiments focus on MLPs to showcase AWNN’s broad applicability, but there are many other scenarios where one can apply AWNN’s principles.
+For instance, one could impose a soft ordering of importance on CNNs’ filters at each layer, therefore learning the number of filters during training.
+From a more theoretical perspective, we believe one could draw connections between our technique and the Information Bottleneck principle  [56], which seeks maximally representative (i.e., performance) and compact representations (e.g., width).
+Finally, the analysis of different functions $f$ to soft-order neurons, for instance a power-law distribution or a sigmoid-like function, may have a different impact on convergence and performance and could be empirically investigated.
+AWNN needs to change the network at training time with an associated overhead.
+However, our current implementation has up to 3-4$\times$ overhead on the smallest datasets, where the inference cost is negligible compared to the network update.
+Although this should not be an issue with more realistic tasks, especially if one considers the time spent testing different widths during grid search, further optimizations are conceivable: i) by updating the width every $M$ mini-batches rather than at every step; ii) by allocating a larger network and dynamically grow it only if needed; iii) by creating ad-hoc support of AWNN on deep learning libraries like PyTorch.
 4 Experiments and Setup
@@
 As such, due to the nature of this work and similarly to  Mitchell et al.
-(2023), we use the remaining space to thoroughly study the behavior of AWNN, so that it becomes clear how to use it in practice.
+[40], we use the remaining space to thoroughly study the behavior of AWNN, so that it becomes clear how to use it in practice.
 We first quantitatively verify that AWNN does not harm the performance compared to baseline models and compare the chosen width by means of grid-search model selection with the learned width of AWNN.
-Secondly, we check that AWNN chooses a higher width for harder tasks, which can be seen as increasing the hypotheses space until the neural network finds a good path to convergence.
+Second, we check that AWNN chooses a larger width for harder tasks, which can be seen as increasing the hypotheses space until the neural network finds a good path to convergence.
 Third, we verify that convergence speed is not significantly altered by AWNN, so that the main limitation lies in the extra overhead for adapting the network at each training step.
@@
 Further analyses are in the Appendix.
-We now provide the details to reproduce our experiments on 9 datasets.
-We compare a baseline that undergoes proper hyper-parameter tuning against its AWNN version, where we replace any fixed MLP with an adaptive one.
+We compare baselines that undergo proper hyper-parameter tuning (called “Fixed”) against its AWNN version, where we replace any fixed MLP with an adaptive one.
 First, we train an MLP on 3 synthetic tabular tasks of increasing binary classification difficulty, namely a double moon, a spiral, and a double spiral that we call SpiralHard.
 A stratified hold-out split of 70% training/10% validation/20% test for risk assessment is chosen at random for these datasets.
-Similarly, we consider a ResNet-20 (He et al., 2016) trained on 3 image classification tasks, namely MNIST (LeCun, 1998), CIFAR10, and CIFAR100 (Krizhevsky, 2009), where data splits and preprocessing are taken from the original paper and AWNN is applied to the downstream classifier.
-In the graph domain, we train a Graph Isomorphism Network (Xu et al., 2019) on the NCI1 and REDDIT-B classification tasks using the same split and evaluation setup of Errica et al.
-(2020).
-These tasks are chosen as the structure was found to have a proper influence on the final performance.
+Similarly, we consider a ResNet-20  [26] trained on 3 image classification tasks, namely MNIST  [32], CIFAR10, and CIFAR100  [31], where data splits and preprocessing are taken from the original paper and AWNN is applied to the downstream classifier.
+In the sequential domain, we implemented a basic adaptive Recurrent Neural Network (RNN) and evaluated on the PMNIST dataset  [67] with the same data split as MNIST.
+In the graph domain, we train a Graph Isomorphism Network  [65] on the NCI1 and REDDIT-B classification tasks, where topological information matters, using the same split and evaluation setup of  Errica et al.
+[16].
 Here, the first 1 hidden layer MLP as well as the one used in each graph convolutional layer are replaced by adaptive AWNN versions.
 On all these tasks, the metric of interest is the accuracy.
-Finally, for the textual domain we train a Transformer architecture (Vaswani et al., 2017) on the Multi30k English-German translation task (Elliott et al., 2016), using a pretrained GPT-2 Tokenizer, and we evaluate the cross-entropy loss over the translated words.
-On tabular, image, and text-based tasks, an internal validation set (10%) for model selection is extracted from the union of outer training and validation sets, and the best configuration chosen according to the internal validation set is retrained 10 times on the outer train/validation/test splits, averaging test performances.
-Due to space reasons, we report datasets statistics and the hyper-parameter tried for the fixed and AWNN versions in Appendix B and C, respectively.
+Finally, for the textual domain we train a Transformer architecture  [58] on the Multi30k English-German translation task  [13], using a pretrained GPT-2 Tokenizer, and we evaluate the cross-entropy loss over the translated words.
+On tabular, image, and text-based tasks, an internal validation set (10%) for model selection is extracted from the union of outer training and validation sets, and the best configuration chosen according to the internal validation set is retrained 10 times on the outer train/validation/test splits, averaging test performances after an early stopping strategy on the validation set.
+Due to space reasons, we report datasets statistics and the hyper-parameter tried for the fixed and AWNN versions in Appendix F and G, respectively [^3] Code to reproduce results is in the supplementary material..
+We ran the experiments on a server with 128 cores, 1.5TB of RAM, and 4 NVIDIA A40 with 46GB of memory.
 5 Results
-We begin by discussing the quantitative results of our experiments: Table 1 reports means and standard deviations across the 10 final training runs.
-In terms of performance, we observe that AWNN is more stable or accurate than a fixed MLP on DoubleMoon, Spiral, and SpiralHard; all other things being equal, it seems that using more neurons and their soft ordering are the main contributing factors to these improvements.
-On the image datasets, performances of AWNN are comparable to those of the fixed baseline but for CIFAR100, due to an unlucky run that did not converge.
-In this case, AWNN learns a smaller total width compared to grid search.
-Table 1: We compare performances and total width of MLP layers for the fixed and AWNN versions of the various models used.
+Table 1: Performances and total width of MLP layers for the fixed and AWNN versions of the various models used.
 The exact width chosen by model selection on the graph datasets is unknown since we report published results.
-“Linear” means that the chosen downstream classifier is a linear model.
+“Linear" means the chosen downstream classifier is a linear model.
 Fixed AWNN Width (Fixed) Width (AWNN)
@@
 CIFAR100 66.5 (0.4) 63.1 (4.0) 256 161.9 (57.8)
+PMNIST 91.1 (0.4) 95.7 (0.2) 24 806.3 (44.5)
 NCI1 80.0 (1.4) 80.0 (1.1) (96-320) 731.3 (128.2)
@@
 Multi30k ($\downarrow$) 1.43 (0.4) 1.51 (0.2) 24576 123.2 (187.9)
+We begin by discussing the quantitative results of our experiments: Table 1 reports means and standard deviations across the 10 final training runs.
+In terms of performance, we observe that AWNN is more stable or accurate than a fixed MLP on DoubleMoon, Spiral, and SpiralHard; all other things being equal, it seems that using more neurons and their soft ordering are the main contributing factors to these improvements.
+On the image datasets, performances of AWNN are comparable to those of the fixed baseline but for CIFAR100, due to an unlucky run that did not converge.
+There, AWNN learns a smaller total width compared to grid search.
+Results on graph datasets are interesting in two respects: First, the performance on REDDIT-B is significantly improved by AWNN both in terms of average performance and stability of results; second, the total learned width is significantly higher than those tried in  Xu et al.
+[65], Errica et al.
+[16], meaning that a biased choice of possible widths had a profound influence on risk estimation for DGN models (i.e., GIN).
+This result makes it evident that it is important to let the network decide how many neurons are necessary to solve the task.
+Appendix H shows what happens when we retrain some fixed baselines using the total width as the width of each layer.
+Finally, the results on the Multi30k show that the AWNN Transformer learns to use 200x parameters less than the fixed Transformer for the feed-forward networks, achieving a statistically comparable test loss.
+This result is appealing when read through the lenses of modern deep learning, as the power required by some neural networks such as Large Language Models  [9] is so high that cannot be afforded by most institutions, and it demands future investigations.
+Adaptation to Task Difficulty and Convergence
+Intuitively, one would expect that AWNN learned larger widths for more difficult tasks.
+This is indeed what happens on the tabular datasets (and image datasets, see Appendix I) where some tasks are clearly harder than others.
+Figure 2 (left) shows that, given the same starting width per layer, the learned number of neurons grows according to the task’s difficulty.
+It is also interesting to observe that the total width for a multi-layer MLP on SpiralHard is significantly lower than that achieved by a single-layer MLP, which is consistent with the circuit complexity theory arguments put forward in  Bengio et al.
+[3], Mhaskar et al.
+[37].
+It also appears that convergence is not affected by the introduction of AWNN, as investigated in Figure 2 (right), which was not obvious considering the parametrization constraints encouraged by the rescaling of neurons’ activations.
 Figure 2: (Left) The learned width adapts to the increasing difficulty of the task, from the DoubleMoon to SpiralHard.
 (Right) AWNN reaches perfect test accuracy with a comparable amount of epochs on DoubleMoon and Spiral, while it converges faster on SpiralHard.
+Training Stability Analysis
+To support our argument that AWNN can reduce the time spent performing hyper-parameter selection, we check whether AWNN learns a consistent amount of neurons across different training runs and hyper-parameter choices.
+Figure 3 reports the impact of the batch size and starting width averaged across the different configurations tried during model selection.
+Smaller batch sizes cause more instability, but in the long run we observe convergence to a similar width.
+Convergence with respect to different rates holds, instead, for the bounded ReLU6 activation; Appendix J shows that unbounded activations may cause the network to converge more slowly to the same width, which is in accord with the considerations about counterbalancing the rescaling effect of Section 3.1.
+Therefore, whenever possible, we recommend using bounded activations.
 Figure 3: Training converges to similar widths on SpiralHard for different batch sizes (left) and starting rates $\nu$, but the latter seems to require a bounded nonlinearity such as ReLU6 to converge in a reasonable amount of epochs (right).
-Results on graph datasets are interesting in two respects: First, the performance on REDDIT-B is significantly improved by AWNN both in terms of average performance and stability of results; second, the total learned width is significantly higher than those tried in Xu et al.
-(2019); Errica et al.
-(2020), meaning that a biased choice of a good range of width has had a profound influence on the estimation of the risk for a specific family of DGN models (i.e., GIN).
-This result makes it evident that it is important to let the network decide how many neurons are necessary to solve the task.
-To check that the learned width is indeed conducive to good performances, Appendix D shows what happens when we retrain some fixed baselines using the total width as the width of each layer.
-Finally, the results on the Multi30k show that the AWNN Transformer learns to use 200x parameters less than the fixed Transformer for the feed-forward networks, achieving a statistically comparable test loss.
-This result is appealing when read through the lenses of modern deep learning, as the power required by some neural networks such as Large Language Models (Brown et al., 2020) is so high that cannot be afforded by most institutions, and it demands future investigations.
-5.1 Adaptation to Task Difficulty and Convergence
-Intuitively, one would expect that AWNN learned larger widths for more difficult tasks.
-This is indeed what happens on the tabular datasets (and image datasets, see Appendix E) where some tasks are clearly harder than others.
-Figure 3 (left) shows that, given the same starting width per layer, the learned number of neurons grows according to the task’s difficulty.
-It is also interesting to observe that the total width for a multi-layer MLP on SpiralHard is significantly lower than that achieved by a single-layer MLP, which is consistent with the circuit complexity theory arguments put forward in Bengio et al.
-(2006).
-It also appears that convergence is not affected by the introduction of AWNN, as investigated in Figure 3 (right), which was not obvious considering the parametrization constraints encouraged by the rescaling of neurons’ activations.
+Online Network Compression via Regularization
+So far, we have used an uninformative prior $p(\lambda)$ over the neural networks’ width.
+We demonstrate the effect of an informative prior by performing a width-annealing experiment on the SpiralHard dataset.
+We set an uninformative $p(\boldsymbol{\theta})$ and ReLU6 nonlinearity.
+At epoch 1000, we introduce $p(\lambda_{\ell})=\mathcal{N}(\lambda_{\ell};0.05,1)$, and gradually anneal the standard deviation up to $0.1$ at epoch $2500$.
+Figure 4 shows that the width of the network reduces from approximately 800 neurons to 300 without any test performance degradation.
+We hypothesize that the least important neurons carry negligible information, therefore they can be safely removed without drastic changes in the output of the model.
+This technique might be particularly useful to compress large models with billions of parameters.
 Figure 4: It is possible to regularize the width at training time by increasing the magnitude of the loss term $\log\frac{p(\boldsymbol{\nu})}{q(\boldsymbol{\nu})}$.
@@
 The inset plot refers to the loss term that AWNN tries to maximize.
+Post-hoc Truncation Achieves a Trade-off between Performance and Compute Resources
+To further investigate the consequences of imposing a soft ordering among neurons, we show that it is possible to perform straightforward post-training truncation while still controlling the trade-off between performance and network size.
+Figure 5 shows an example for an MLP on the Spiral dataset, where the range of activation values (Equation 10) computed for all samples follows an exponential curve (right).
+Intuitively, removing the last neurons may have a negligible performance impact at the beginning and a drastic one as few neurons remain.
+This is what happens, where we are able to cut an MLP with hidden width 83 by 30% without loss of accuracy, after which a smooth degradation happens.
+If one accepts such a trade-off, this technique may be used to “distill” a trained neural network at virtually zero additional cost while reducing the memory requirements.
+Note that truncation heuristics that are either random or based on the magnitude of neurons’ activations (i.e., excluding the rescaling term) do not perform as well.
 Figure 5: (Left) Thanks to the soft ordering imposed on the neurons, one can also truncate the neural network after training by simply removing the last neurons.
 (Right) The distribution of neurons’ activations for all Spiral test samples follows an exponential-like curve.
-5.2 Training Stability Analysis
-To support our argument that AWNN can reduce the time spent performing hyper-parameter selection, it is instructive to check whether AWNN learns a consistent amount of neurons across different training runs and hyper-parameter choices.
-Figure 4 reports the impact of the batch size and starting width averaged across the different configurations tried during model selection.
-Smaller batch sizes cause more instability, but in the long run we observe convergence to a similar width.
-Convergence with respect to different rates holds, instead, for the bounded ReLU6 activation; Appendix F shows that unbounded activations may cause the network to converge more slowly to the same width, which is in accord with the considerations about counterbalancing the rescaling effect of Section 3.1.
-Therefore, whenever possible, we recommend using bounded activations.
-5.3 Online Network Compression via Regularization
-So far, we have used an uninformative prior $p(\lambda)$ over the neural networks’ width.
-We demonstrate the effect of an informative prior by performing an annealing experiment on the SpiralHard dataset.
-We set an uninformative $p(\boldsymbol{\theta})$ and ReLU6 nonlinearity.
-At epoch 1000, we introduce $p(\lambda_{\ell})=\mathcal{N}(\lambda_{\ell};0.05,1)$, and gradually anneal the standard deviation up to $0.1$ at epoch $2500$.
-Figure 5 shows that the width of the network reduces from approximately 800 neurons to 300 without any test performance de gradation.
-We hypothesize that the least important neurons mostly carry negligible information, and therefore they can be safely removed without drastic changes in the output of the model.
-This technique might be particularly useful to compress large models with billions of parameters.
-5.4 Post-hoc Trucation achieves a trade-off between performance and compute resources
-To further investigate the advantages of imposing a soft ordering of importance among neurons, we show that it is possible to perform hard network’s truncation after training while still controlling the trade-off between performance and network size.
-Figure 6 shows an example for an MLP on the Spiral dataset, where the range of activation values (Equation 14) computed for all samples follows an exponential curve (right).
-Therefore, it is expected that removing the last neurons may have a negligible performance impact at the beginning and a drastic one as few neurons remain.
-This is what happens in the left plot, where we are able to cut an MLP with hidden width 83 by 30% without loss of accuracy, after which a smooth degradation happens.
-If one accepts such a trade-off, this technique may be used to distill a trained neural network at virtually zero additional cost while reducing the memory requirements.
 6 Conclusions
 We introduced a new methodology to learn an unbounded width of neural network layers within a single training, by imposing a soft ordering of importance among neurons.
-Our approach requires very few changes to the architecture, adapts the width to the task’s difficulty, and does not impact negatively convergence, although it introduces some overhead during training.
-We showed stability convergence to similar widths under bounded activations for different hyper-parameters configurations, advocating for a practical reduction of the width’s search space.
+Our approach requires very few changes to the architecture, adapts the width to the task’s difficulty, and does not impact negatively convergence.
+We showed stability of convergence to similar widths under bounded activations for different hyper-parameters configurations, advocating for a practical reduction of the width’s search space.
 A by-product of neurons’ ordering is the ability to easily compress the network during or after training, which is relevant in the context of foundational models trained on large data, which are believed to require billions of parameters.
-Finally, we have tested AWNN on different models and data domains to prove its broad scope of applicability: a Transformer architecture achieved the same loss with 200x less parameters.
-Impact Statement
-This paper presents work whose goal is to advance the field of Machine Learning.
-There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
-Appendix A Full derivation of the rescaled weight initialization
+Finally, we have tested AWNN on different models and data domains to prove its broad scope of applicability: a Transformer architecture achieved a similar loss with 200x less parameters.
+Appendix A Background Notions
+Since we were mostly inspired by the work of  Nazaret and Blei [42], we complement the main text with an introduction to truncated distributions.
+In particular, by truncating a distribution at its quantile function evaluated at $k$ (see Appendix B for a visual explanation), the support of the distribution becomes finite and countable, hence we can compute expectations in a finite number of steps.
+A truncated distribution should ideally satisfy some requirements defined in  Nazaret and Blei [42] and illustrated below.
+Definition A.1.
+A variational family $Q={q(\boldsymbol{\omega})}$ over $\mathbb{N}^{+}$ is unbounded with connected and bounded members if
+  1. 1.
+    $\forall q\in Q$, $\operatorname{support}(q)$ is bounded
+  2. 2.
+    $\forall L\in\mathbb{N}^{+},\exists q\in Q$ such that $L\in\operatorname{argmax}(q)$
+  3. 3.
+    Each parameter in the set $\boldsymbol{\omega}$ is a continuous variable.
+The first condition allows us to compute the expectation over any $q\in Q$ in finite time, condition 2 ensures there is a parametrization that assigns enough probability mass to each point in the support of $q$, and condition 3 is necessary for learning via backpropagation.
+An important distinction with  Nazaret and Blei [42] is that we do not need nor want condition 2 to be satisfied.
+As a matter of fact, we want to enforce a strict ordering of neurons where the first is always the most important one.
+Condition 2 is useful, as done in  Nazaret and Blei [42], to assign enough ‘importance” to a specific layer of an adaptive-depth architecture, while previous layers are still functional to the final result.
+In the context of adaptive-width networks, however, this mechanism may leave a lot of neurons completely unutilized, letting the network grow indefinitely.
+That is why we require that each distribution $q$ is monotonically decreasing and all possible widths should have non-zero importance.
+We change the definition for the distributions we are interested in as follows:
+Definition A.2.
+A variational family $Q={q(\boldsymbol{\omega})}$ over $\mathbb{N}^{+}$ is unbounded with connected, bounded, and decreasing members if
+  1. 1.
+    $\forall q\in Q$, $\operatorname{support}(q)$ is bounded
+  2. 2.
+    $\forall q\in Q$, $q$ is monotonically decreasing and $\forall x\in\mathbb{N}^{+},q(x)>0$
+  3. 3.
+    Each parameter in the set $\boldsymbol{\omega}$ is a continuous variable.
+Appendix B How to Compute $D_{\ell}$ in Practice
+To aid the understanding of how one computes $D_{\ell}$, we provide a visualization of the exponential distribution, its discretized version (Equation 6) and the quantile function evaluated at $k=0.9$ in Figure 6 below.
+Figure 6: On the left, we provide a visualization of the exponential density with $\lambda=0.5$, its cumulative density, and the quantile function evaluated at $k=0.9$, which returns a value close to $5$.
+On the right, we show the discretized exponential and its cumulative mass function; we can see that the quantile function of the continuous density provides an upper bound for the natural where the cumulative mass function is greater than $k=0.9$.
+Therefore, one can safety use the ceiling of the output of the quantile function to determine the width $D_{\ell}$ to use for a given value of $\lambda$.
+Appendix C Full ELBO Derivation
+The full ELBO derivation is as follows:
+$\mathbb{E}_{q(\boldsymbol{\lambda},\boldsymbol{\theta})}\left[\log\frac{p(Y,\boldsymbol{\lambda},\boldsymbol{\theta}|X)}{q(\boldsymbol{\lambda},\boldsymbol{\theta})}\right]=$ (11)
+$\mathbb{E}_{q(\boldsymbol{\lambda})}\mathbb{E}_{q(\boldsymbol{\theta}|\boldsymbol{\lambda})}[\log p(Y,\boldsymbol{\lambda},\boldsymbol{\theta}|X)-\log q(\boldsymbol{\lambda},\boldsymbol{\theta})]=$ (12)
+$\mathbb{E}_{q(\boldsymbol{\lambda})}\mathbb{E}_{q(\boldsymbol{\boldsymbol{\theta}|\lambda})}[\log(p(Y|\boldsymbol{\lambda},\boldsymbol{\theta},X)p(\boldsymbol{\theta})p(\boldsymbol{\lambda}))-\log(q(\boldsymbol{\lambda})q(\boldsymbol{\theta}|\boldsymbol{\lambda}))]$ (13)
+where we factorized probabilities according to the independence assumptions of the graphical model. Then
+$\mathbb{E}_{q(\boldsymbol{\lambda})}\mathbb{E}_{q(\boldsymbol{\boldsymbol{\theta}|\lambda})}[\log p(Y|\boldsymbol{\lambda},\boldsymbol{\theta},X)+\log p(\boldsymbol{\lambda})+\log p(\boldsymbol{\theta})-\log q(\boldsymbol{\lambda})-\log q(\boldsymbol{\theta}|\boldsymbol{\lambda})]=$ (14)
+$\mathbb{E}_{q(\boldsymbol{\lambda})}\mathbb{E}_{q(\boldsymbol{\boldsymbol{\theta}|\lambda})}[\log\frac{p(\boldsymbol{\lambda})}{q(\boldsymbol{\lambda})}+\log\frac{p(\boldsymbol{\theta})}{q(\boldsymbol{\theta}|\boldsymbol{\lambda})}+\log p(Y|\boldsymbol{\lambda},\boldsymbol{\theta},X)]\approx$ (15)
+$\log\frac{p(\boldsymbol{\nu})}{q(\boldsymbol{\nu})}+\log\frac{p(\boldsymbol{\rho})}{q(\boldsymbol{\rho}|\boldsymbol{\nu})}+\log p(Y|\boldsymbol{\nu},\boldsymbol{\rho},X)]$ (16)
+where we use the first-order approximation in the last step.
+Note that when computing $\frac{p(\theta)}{q(\theta|\lambda)}$, the products from $D_{\ell+1}$ to infinity cancel out.
+Equation 13 follows by simply applying definitions, the iid assumption $p(Y|\boldsymbol{\nu},\boldsymbol{\rho},X)=\prod_{i}^{N}p(y_{i}|\boldsymbol{\nu},\boldsymbol{\rho},x_{i})$, and by transforming products into summations using the logarithm.
+An important note about the first-order (Taylor) approximation $E_{q(\boldsymbol{\lambda};\boldsymbol{\nu})q(\boldsymbol{\theta}|\boldsymbol{\lambda};\boldsymbol{\rho})}[f(\boldsymbol{\lambda},\boldsymbol{\theta})]\approx E_{q(\boldsymbol{\lambda};\boldsymbol{\nu})}[f(\boldsymbol{\lambda},\boldsymbol{\rho})]\approx f(\boldsymbol{\nu},\boldsymbol{\rho})$ is that the function $f(\boldsymbol{\lambda},\boldsymbol{\theta})=\log\frac{p(\boldsymbol{\lambda})}{q(\boldsymbol{\lambda})}+\log\frac{p(\boldsymbol{\theta})}{q(\boldsymbol{\theta}|\boldsymbol{\lambda})}+\log p(Y|\boldsymbol{\lambda},\boldsymbol{\theta},X)$ is not differentiable with respect to $D_{\ell}$ when there is a jump in the quantile evaluated at $k$, so it would seem we cannot apply it.
+This also happens in  Nazaret and Blei [42], Equation 11, when the function is a neural network containing non-differentiable activations like ReLU, but we could not find a discussion about it.
+Importantly, the function is not differentiable at a countable set of points (one for each possible width), which has a measure null.
+Since the Lebesgue integral of the expectation is over reals, we should be able to remove this set from the integral while ensuring that the approximation still holds mathematically.
+Appendix D Lipschitz Continuity of the ELBO
+Whenever the quantile of the finite distribution changes, we need to change the number of neurons of the neural network.
+Here, we want to show that the change in the ELBO after a change in the number of neurons is bounded.
+Therefore, we provide a theoretical result (Proposition D.1) that shows that the ELBO satisfies the Lipschitz continuity.
+Proposition D.1 (AWNN ELBO Lipschitz continuity).
+The ELBO loss of Equation 9, with respect to the change in the depth $D_{\ell}$ for the layer $l$, is Lipschitz continuous.
+Proof.
+We focus on the term involving $D_{\ell}$ of the ELBO, we write Equation 9 as
+$\log\frac{p(\boldsymbol{\nu})}{q(\boldsymbol{\nu})}+\log\frac{p(\boldsymbol{\rho})}{q(\boldsymbol{\rho}|\boldsymbol{\nu})}+\log p(Y|\boldsymbol{\nu},\boldsymbol{\rho},X)$
+where only the second and last terms depend on $\mathbf{D}=\{D_{\ell}\}_{\ell=1}^{L}$ Let’s first define
+$\log\frac{p(\boldsymbol{\rho})}{q(\boldsymbol{\rho}|\boldsymbol{\nu})}=\sum_{{\ell}=1}^{L}\sum_{n=1}^{D_{\ell}}\log\frac{p(\rho^{\ell}_{n})}{q(\rho^{\ell}_{n}|\boldsymbol{\nu})}=\sum_{{\ell}=1}^{L}f_{1}(D_{\ell})$
+We have that $f_{1}(D_{\ell})$ is Lipschitz continuous, indeed, when $D_{\ell}$ changes to $D_{l^{\prime}}$, we have
+$|f_{1}(D_{\ell^{\prime}})-f_{1}(D_{\ell})|$ $=|\sum_{n=D_{\ell}}^{D_{\ell^{\prime}}}\log\frac{p(\boldsymbol{\rho}_{n})}{q(\boldsymbol{\rho}_{n}|\boldsymbol{\nu})}|$ (17)
+$\leq~\sum_{n=D_{l}}^{D_{\ell^{\prime}}}|\log\frac{p(\boldsymbol{\rho}_{n})}{q(\boldsymbol{\rho}_{n}|\boldsymbol{\nu})}|$ (18)
+$\leq\max_{n}|\log\frac{p(\boldsymbol{\rho}_{n})}{q(\boldsymbol{\rho}_{n}|\boldsymbol{\nu})}||D_{\ell^{\prime}}-D_{\ell}|$ (19)
+Therefore
+$|f_{1}(D_{\ell^{\prime}})-f_{1}(D_{\ell})|\leq M|D_{\ell^{\prime}}-D_{\ell}|$
+with $M=\max_{n}|\log\frac{p(\boldsymbol{\rho}_{n})}{q(\boldsymbol{\rho}_{n}|\boldsymbol{\nu})}|$. We now look at the last term,
+$f_{2}(\mathbf{D})=\log p(Y|\boldsymbol{\nu},\boldsymbol{\rho},X)$
+This function is a neural network followed by the squared norm.
+Therefore, $f_{2}(\mathbf{D})$ is continuous almost everywhere  [59], therefore Lipschitz on both parameters and the input.
+When we change $D_{\ell}$ to $D_{\ell^{\prime}}$, we are adding network parameters, in particular, if we look at the two-layer network, we have
+$y=V\sigma(Wx+b)+c,V\in\mathbb{R}^{m\times n},W\in\mathbb{R}^{n\times d}$
+we then have new parameters, which can be shown to be
+$y^{\prime}=V\sigma(Wx+b)+c+\underbrace{V^{\prime}\sigma(W^{\prime}x+b^{\prime})+c^{\prime}}_{\text{new neurons}},V^{\prime}\in\mathbb{R}^{m\times n^{\prime}},W^{\prime}\in\mathbb{R}^{n^{\prime}\times d},b^{\prime}\in\mathbb{R}^{n^{\prime}},c^{\prime}\in\mathbb{R}^{m}$
+the second term is also Lipschitz, therefore the function $f_{2}(\mathbf{D})$ is also Lipschitz.
+Indeed the network is composed of a linear operation and the element-wise activation functions.
+If the activation functions are continuous and of bounded gradient, then the whole network is Lipschitz continuous.
+∎
+Appendix E Full derivation of the rescaled weight initialization
 This Section derives the formulas for the rescaled weight initialization both in the case of ReLU and activations like tanh.
@@
 Proof.
-Let us start from the first case of $\sigma^{\prime}(0)\approx 1$. Using the Taylor expansions for the moments of functions of random variables as in Glorot & Bengio (2010)
+Let us start from the first case of $\sigma^{\prime}(0)\approx 1$. Using the Taylor expansions for the moments of functions of random variables as in  Glorot and Bengio [20]
 $\mathrm{Var}[\alpha^{\ell}_{i}]=\mathrm{Var}\left[\sigma\left(\sum_{j=1}^{D_{\ell-1}}w^{\ell}_{ij}\alpha^{\ell-1}_{j}p^{\ell-1}_{j}\right)\right]$ (25)
@@
 We can also abstract from the indexes, since the weight variables are i.i.d.
-and from that it follows that $\mathrm{Var}[\alpha^{\ell}_{i}]=\mathrm{Var}[\alpha^{\ell}_{j}]\ \forall i,j,\ \ i,j\in\{1,\dots,D_{\ell}\}$, obtaining [^3] It is worth noting that, in standard MLPs, $p_{j}^{\ell}=1$ so we recover the derivation of Glorot & Bengio (2010), since $\sum_{j=1}^{D_{\ell-1}}\left(p^{\ell-1}_{j}\right)^{2}$ would be equal to $D_{\ell-1}$.
-In Glorot & Bengio (2010) our $D_{\ell-1}$ is denoted as “$n_{i^{\prime}}$” (see Equation 5).
+and from that it follows that $\mathrm{Var}[\alpha^{\ell}_{i}]=\mathrm{Var}[\alpha^{\ell}_{j}]\ \forall i,j,\ \ i,j\in\{1,\dots,D_{\ell}\}$, obtaining [^4] It is worth noting that, in standard MLPs, $p_{j}^{\ell}=1$ so we recover the derivation of  Glorot and Bengio [20], since $\sum_{j=1}^{D_{\ell-1}}\left(p^{\ell-1}_{j}\right)^{2}$ would be equal to $D_{\ell-1}$.
+In  Glorot and Bengio [20] our $D_{\ell-1}$ is denoted as “$n_{i^{\prime}}$” (see Equation 5).
 $\mathrm{Var}[\alpha^{\ell}]\approx\mathrm{Var}[w^{\ell}]\mathrm{Var}[\alpha^{\ell-1}]\sum_{j=1}^{D_{\ell-1}}\left(p^{\ell-1}_{j}\right)^{2},$ (31)
@@
 Condition on the gradients
-From a backpropagation perspective, a similar desideratum would be to ensure that $\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{i}^{\ell})}{\partial\alpha_{i}^{\ell}}\right]=\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{i}^{\ell+1})}{\partial\alpha_{i}^{\ell+1}}\right]$ [^4] Note that what we impose is different from Glorot & Bengio (2010), where the specific position $i$ is irrelevant.
+From a backpropagation perspective, a similar desideratum would be to ensure that $\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{i}^{\ell})}{\partial\alpha_{i}^{\ell}}\right]=\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{i}^{\ell+1})}{\partial\alpha_{i}^{\ell+1}}\right]$ [^5] Note that what we impose is different from  Glorot and Bengio [20], where the specific position $i$ is irrelevant.
 Here, we are asking that the variance of the gradients for neurons in the same position $i$, but at different layers, stays constant.
 Alternatively, we could impose an equivalence for all $i\neq i^{\prime}$..
-Using Equation 18, and considering as in Glorot & Bengio (2010) that at initialization we are in a linear regime where $\sigma^{\prime}(x)\approx 1$,
+Using Equation 22, and considering as in  Glorot and Bengio [20] that at initialization we are in a linear regime where $\sigma^{\prime}(x)\approx 1$,
 $\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{i}^{\ell})}{\partial\alpha_{i}^{\ell}}\right]=\mathrm{Var}\left[\sum_{j=1}^{D_{\ell+1}}\frac{\partial\mathcal{L}(\alpha_{j}^{\ell+1})}{\partial\alpha_{j}^{\ell+1}}\frac{\partial\alpha_{j}^{\ell+1}(\alpha_{i}^{\ell})}{\partial\alpha_{i}^{\ell}}\right]=\mathrm{Var}\left[\sum_{j=1}^{D_{\ell+1}}\frac{\partial\mathcal{L}(\alpha_{j}^{\ell+1})}{\partial\alpha_{j}^{\ell+1}}\frac{\partial\alpha_{j}^{\ell+1}(\alpha_{i}^{\ell})}{\partial\alpha_{i}^{\ell}}\right]$ (33)
@@
 Expanding, we get
-$\mathrm{Var}[w^{\ell+1}](p^{\ell}_{i})^{2}\sum_{j=1}^{D_{\ell+1}}\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{j}^{\ell+1})}{\partial\alpha_{j}^{\ell+1}}\right]=(p^{\ell}_{i})^{2}\left(\prod_{i=\ell+1}^{\ell+2}\mathrm{Var}[w^{\ell+1}]\right)\sum_{j=1}^{D_{\ell+1}}(p^{\ell+1}_{j})^{2}\underbrace{\left(\sum_{j^{\prime}=1}^{D_{\ell+2}}\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{j^{\prime}}^{\ell+2})}{\partial\alpha_{j^{\prime}}^{\ell+2}}\right]\right)}_{\text{constant w.r.t.\ }j}$ (32)
+$\mathrm{Var}[w^{\ell+1}](p^{\ell}_{i})^{2}\sum_{j=1}^{D_{\ell+1}}\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{j}^{\ell+1})}{\partial\alpha_{j}^{\ell+1}}\right]$ (36)
+$=(p^{\ell}_{i})^{2}\left(\prod_{i=\ell+1}^{\ell+2}\mathrm{Var}[w^{\ell+1}]\right)\sum_{j=1}^{D_{\ell+1}}(p^{\ell+1}_{j})^{2}\underbrace{\left(\sum_{j^{\prime}=1}^{D_{\ell+2}}\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{j^{\prime}}^{\ell+2})}{\partial\alpha_{j^{\prime}}^{\ell+2}}\right]\right)}_{\text{constant w.r.t.\ }j}$ (37)
 $=(p^{\ell}_{i})^{2}\left(\prod_{i=\ell+1}^{\ell+2}\mathrm{Var}[w^{\ell+1}]\right)\left(\sum_{j^{\prime}=1}^{D_{\ell+2}}\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{j^{\prime}}^{\ell+2})}{\partial\alpha_{j^{\prime}}^{\ell+2}}\right]\right)\left(\sum_{j=1}^{D_{\ell+1}}(p^{\ell+1}_{j})^{2}\right)$ (38)
@@
 $\mathrm{Var}[w^{\ell+1}]=\frac{(p^{\ell+1}_{i})^{2}}{(p^{\ell}_{i})^{2}\sum_{j=1}^{D_{\ell+1}}\left(p^{\ell+1}_{j}\right)^{2}}.$ (42)
-Both Equations 28 and 37 are verified when we initialize the distributions $f_{\ell}$ in the same way for all layers, which implies $(p^{\ell}_{i})^{2}=(p^{\ell+1}_{i})^{2}$.
+Both Equations 32 and 42 are verified when we initialize the distributions $f_{\ell}$ in the same way for all layers, which implies $(p^{\ell}_{i})^{2}=(p^{\ell+1}_{i})^{2}$.
 Note that without this last requirement, Equation 42 would violate the i.i.d.
@@
 To show a similar initialization for ReLU activations, we need the following lemma.
-Lemma A.1.
+Lemma E.1.
 Consider the ReLU activation $y=\max(0,x)$ and a symmetric distribution $p(x)$ around zero. Then $\mathbb{E}[y^{2}]=\frac{1}{2}Var[x]$.
@@
 […]
-Using Lemma A.1, we can rewrite Equation 22 as
-$\mathrm{Var}[\alpha^{\ell}_{i}]=\mathrm{Var}\left[\sigma\left(\sum_{j=1}^{D_{\ell-1}}w^{\ell}_{ij}\alpha^{\ell-1}_{j}p^{\ell-1}_{j}\right)\right]=\frac{1}{2}\mathrm{Var}\left[\sum_{j=1}^{D_{\ell-1}}w^{\ell}_{ij}\alpha^{\ell-1}_{j}p^{\ell-1}_{j}\right]=\frac{1}{2}\mathrm{Var}[w^{\ell}]\mathrm{Var}[\alpha^{\ell-1}]\sum_{j=1}^{D_{\ell-1}}\left(p^{\ell-1}_{j}\right)^{2}$ (40)
+Using Lemma E.1, we can rewrite Equation 26 as
+$\mathrm{Var}[\alpha^{\ell}_{i}]=\mathrm{Var}\left[\sigma\left(\sum_{j=1}^{D_{\ell-1}}w^{\ell}_{ij}\alpha^{\ell-1}_{j}p^{\ell-1}_{j}\right)\right]=\frac{1}{2}\mathrm{Var}\left[\sum_{j=1}^{D_{\ell-1}}w^{\ell}_{ij}\alpha^{\ell-1}_{j}p^{\ell-1}_{j}\right]$ (45)
+$=\frac{1}{2}\mathrm{Var}[w^{\ell}]\mathrm{Var}[\alpha^{\ell-1}]\sum_{j=1}^{D_{\ell-1}}\left(p^{\ell-1}_{j}\right)^{2}$ (46)
 where no approximation has been made. Similar to prior results, it follows that
@@
 From a backpropagation perspective, since $\sigma^{\prime}(x)\not\approx 1$, we use similar arguments as in  He et al.
-(2015): we assume that the pre-activations $\alpha$ have zero mean, then the derivative of the ReLU can have values zero or one with equal probability.
+[25]: we assume that the pre-activations $\alpha$ have zero mean, then the derivative of the ReLU can have values zero or one with equal probability.
 One can show, by expanding the definition of variance, that $\mathrm{Var}[\sigma^{\prime}\left(\sum_{k=1}^{D_{\ell}}w^{\ell+1}_{jk}\alpha^{\ell}_{k}p^{\ell}_{k}\right)w^{\ell+1}_{ji}]=\frac{1}{2}\mathrm{Var}[w^{\ell+1}_{ji}]$.
@@
 $\approx\frac{1}{2}\mathrm{Var}[w^{\ell+1}](p^{\ell}_{i})^{2}\sum_{j=1}^{D_{\ell+1}}\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{j}^{\ell+1})}{\partial\alpha_{j}^{\ell+1}}\right],$ (49)
-where we have made a rough assumption of independence between $\sigma^{\prime}\left(\sum_{k=1}^{D_{\ell}}w^{\ell+1}_{jk}\alpha^{\ell}_{k}p^{\ell}_{k}\right)$ and $w^{\ell+1}_{ji}$ since the former term is likely a constant.
+where we have made the assumption of independence between $\sigma^{\prime}\left(\sum_{k=1}^{D_{\ell}}w^{\ell+1}_{jk}\alpha^{\ell}_{k}p^{\ell}_{k}\right)$ and $w^{\ell+1}_{ji}$ since the former term is likely a constant.
 Following an identical derivation as above, we obtain
@@
 Comparison of convergence for different initializations
-Below, we provide a comparison of convergence between an AWNN ReLU MLP initialized with the standard kaiming scheme and one initialized according to our theoretical results (“kaiming+”).
+Below, we provide a comparison of convergence between an AWNN ReLU MLP initialized with the standard Kaiming scheme and one initialized according to our theoretical results (“Kaiming+").
 Figure 7: (Left) Effect of the rescaled initialization scheme (“Kaiming+”) on a ReLU-based MLP, where neurons’ activations are computed using Equation 10.
@@
 (Right) Without the rescaled initialization, convergence is hard to attain on the SpiralHard dataset (please refer to the next section for details about the dataset).
-Appendix B Dataset Info and Statistics
+Appendix F Dataset Info and Statistics
 The synthetic datasets DoubleMoon, Spiral, and SpiralHard are shown in Figure 8.
@@
 For a detailed description of the graph datasets, the reader can refer to  Errica et al.
-(2020).
-Appendix C Hyper-parameters for fixed and AWNN models
+[16].
+Appendix G Hyper-parameters for fixed and AWNN models
 For each data domain, we discuss the set of hyper-parameters tried during model selection for the fixed baseline and for AWNN.
 We try our best to keep architectural choices identical with the exception of the width of the MLPs used by each model.
-In all experiments, we either run patience-based early stopping (Prechelt, 1998) or simply select the epoch with best validation score.
-C.1 Synthetic Tabular Datasets
-We report the hyper-parameter configurations tried for the fixed MLP and AWNN in Table 3.
+In all experiments, we either run patience-based early stopping  [46] or simply select the epoch with best validation score.
+G.1 Synthetic Tabular Datasets and Permuted MNIST
+We report the hyper-parameter configurations tried for the fixed MLP/RNN and AWNN in Table 3.
 In particular, the starting width of AWNN under the quantile function evaluated at $k=0.9$ is approximately $256$ hidden neurons.
 We also noticed that a learning rate value of $0.1$ makes learning the width unstable, so we did not use it in our experiments.
-The total number of configurations for the fixed MLP is 180, for AWNN is 18.
+The total number of configurations for the fixed models is 180, for AWNN is 18.
 We also did not impose any prior on an expected size to see if AWNN recovers a similar width compared to the best fixed model (please refer to Table 1) starting from the largest configuration among the fixed models.
-Table 3: Hyper-parameter configurations for standard MLP and AWNN versions on tabular datasets.
+Table 3: Hyper-parameter configurations for standard MLP/RNN and AWNN versions on tabular datasets.
 Hyper-Parameter DoubleMoon Spiral SpiralHard
-Batch Size 32 $[32,128]$ (MLP), 128 (AWNN) $[32,128]$ (MLP), 128 (AWNN)
+Batch Size 32 $[32,128]$ (MLP/RNN), 128 (AWNN) $[32,128]$ (MLP/RNN), 128 (AWNN)
 Epochs 500 1000 5000
@@
 Non-linearity [ReLU, LeakyReLU, ReLU6]
-Optimizer Adam, learning rate $\in[0.1,0.01]$ (MLP), $0.01$ (AWNN)
+Optimizer Adam, learning rate $\in[0.1,0.01]$ (MLP/RNN), $0.01$ (AWNN)
 AWNN Specific
@@
 Prior over $\lambda$ Uninformative
-C.2 Image Classification Datasets
+G.2 Image Classification Datasets
 We train from scratch a ResNet20 and focus our architectural choices on the MLP that performs classification using the flattened representation provided by the previous CNN layers.
@@
 The rate of the exponential distribution has been chosen to have a starting width of approximately $128$ neurons, which is the median value among the ones tried for the fixed model.
-C.3 Text Translation Tasks
+G.3 Text Translation Tasks
 Due to the cost of training a Transformer architecture from scratch on a medium size dataset like Multi30k, we use most hyper-parameters’ configurations of the base Transformer in  Vaswani et al.
-(2017), with the difference that AWNN is applied to the MLPs in each encoder and decoder layer.
+[58], with the difference that AWNN is applied to the MLPs in each encoder and decoder layer.
 We train an architecture with 6 encoder and 6 decoder layers for 500 epochs, with a patience of 250 epochs applied to the validation loss.
@@
 The AWNN version starts with an exponential distribution rate of 0.004, corresponding to approximately 512 neurons, and a $\sigma^{\alpha}_{\ell}$=1.
-C.4 Graph Classification Tasks
+G.4 Graph Classification Tasks
 We train the AWNN version of Graph Isomorphism Network, following the setup of  Errica et al.
-(2020) and reusing the published results.
+[16] and reusing the published results.
 We test 12 configurations as shown in the table below.
@@
 Prior over $\lambda$ Uninformative
-Appendix D Re-training a fixed network with the learned width
+Appendix H Re-training a fixed network with the learned width
 While on DoubleMoon AWNN learns a similar number of neurons chosen by the model selection procedure and the convergence to the perfect solution over 10 runs seem much more stable, this is not the case on Spiral, SpiralHard and REDDIT-B, where AWNN learns a higher total width over layers.
@@
 Notice that the number of neurons learned by AWNN on REDDIT-B was well outside of the range investigated by  Xu et al.
-(2019) and Errica et al.
-(2020), demonstrating that the practitioner’s bias can play an important role on the final performance.
-Appendix E Further qualitative results on image classification datasets
+[65] and  Errica et al.
+[16], demonstrating that the practitioner’s bias can play an important role on the final performance.
+Appendix I Further qualitative results on image classification datasets
 Because it is well known that MNIST, CIFAR10, and CIFAR100 are datasets of increasing difficulty, we attach a figure depicting the learned number of neurons and the convergence of AWNN on these datasets, following the practices highlighted in  He et al.
-(2016) and described in Table 4 to bring the ResNet20 to convergence.
+[26] and described in Table 4 to bring the ResNet20 to convergence.
 Akin to the tabular datasets, we see a clear pattern where the learned number of neurons, averaged over the 10 final training runs, increases together with the task difficulty.
@@
 Figure 9: We present an analysis similar to Figure 2 but for image classification datasets of increasing difficulty. One can observe the jumps caused by the learning rate’s scheduling strategy.
-Appendix F Non-linearity ablation study
+Appendix J Non-linearity ablation study
 Below, we report an ablation study on the impact of non-linear activation functions.
@@
 The decreasing trend for LeakyReLU and ReLU activations may suggest that these configurations are also converging to a similar value than ReLU6, but they are taking a much longer time.
-Appendix G Learned width for each layer on Graph Datasets
+Appendix K Learned width for each layer on Graph Datasets
 Figure 11: Learned neurons per layer (from 1 to 5), averaged over 10 final runs for each of the 10 best configurations selected in the outer folds.
@@
 One can observe similar trends in most cases.
+Appendix L Comparison with Neural Architecture Search (NAS)
+Although complementary to our approach, we thought instructive to run a some basic NAS methods on the image classification datasets.
+Most NAS methods fix an upper bound on the maximum width, whereas AWNN’s methodology does not.
+The truncation ability of AWNN is not intended as a substitute for pruning but rather as a confirmation of our claims.
+In fact, NAS and pruning can actually be combined with AWNN and do not necessarily be seen as “competing” strategies.
+The results are shown in Table 6.
+That said, we compared the performance of AWNN against simple NAS methods on the vision datasets.
+We tested grid search, random search (Following  Wu et al.
+[64]), local search  [60], and Bayesian Optimization  [1], fixing the same budget of 5 for all methods and width values between 0 (Linear classifier) and 512 neurons.
+Results are comparable to those of AWNN.
+Table 6: Comparison between different NAS approaches and AWNN. We remind the reader that NAS is an orthogonal and complementary research direction to AWNN.
+Grid Search Random Search Local Search Bayes. Optim. AWNN
+MNIST 99.6(0.1) 99.6(0.0) 99.5 99.4 99.7(0.0)
+CIFAR10 91.4(0.2) 90.1(0.5) 90.6 91.2 91.4(0.2)
+CIFAR100 66.5(0.4) 64.9(1.1) 64.9 65.9 63.1(4.0)
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: Abstract and introduction reflect our results.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: Section 3.3 discusses limitations and future extensions.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification: All theorems in the appendix come with a complete proof with the necessary assumptions.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19.
+Justification: The paper provides experimental details about the datasets, data splits, and hyper-parameters tried in the appendix.
+Also, the experiments are easy to reproduce thanks to the adoption of ad-hoc libraries in the supplementary material.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: Yes, see also our previous answer.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: Yes, see also our previous answers.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: See Table 1.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: See Section 4.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: Our approach is a general machine learning methodology that could be used in unforeseen ways. We did not experiment on human subjects and the data is publicly available.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [N/A]
+  49. Justification: No societal impact of the work performed.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: No such risks.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [No]
+  59. Justification: We do rely on Python packages but did not acknowledge each of them in our paper.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64. Justification: See the supplementary material.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: No research on human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: No research on human subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [N/A]
+  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v4

@@
 Abstract
-For almost 70 years, researchers have mostly relied on hyper-parameter tuning to select the width of neural networks’ layers.
+For almost 70 years, researchers have primarily relied on hyper-parameter tuning to select the width of neural networks’ layers.
 This paper challenges the status quo by introducing an easy-to-use technique to learn an unbounded width of a neural network’s layer during training.
-The technique does not rely on alternate optimization nor hand-crafted gradient heuristics; rather, it jointly optimizes the width and the parameters of each layer via simple backpropagation.
+The method jointly optimizes the width and the parameters of each layer via standard backpropagation.
 We apply the technique to a broad range of data domains such as tables, images, text, sequences, and graphs, showing how the width adapts to the task’s difficulty.
-The method imposes a soft ordering of importance among neurons, by which it also is possible to truncate the trained network at virtually zero cost, achieving a smooth trade-off between performance and compute resources in a structured way.
-Alternatively, one can dynamically compress the network with no performance degradation.
-In light of recent foundation models trained on large datasets, believed to require billions of parameters and where hyper-parameter tuning is unfeasible due to humongous training costs, our approach stands as a viable alternative for width learning.
+A by product of our width learning approach is the easy truncation of the trained network at virtually zero cost, achieving a smooth trade-off between performance and compute resources.
+Alternatively, one can dynamically compress the network until performances do not degrade.
+In light of recent foundation models trained on large datasets, requiring billions of parameters and where hyper-parameter tuning is unfeasible due to huge training costs, our approach introduces a viable alternative for width learning.
 1 Introduction
@@
 Other techniques to tune hyper-parameters include constructive approaches [17, 64], which alternate parameter optimization and creation of new neurons, natural gradient-based heuristics that dynamically modify the network [40], bi-level optimization [19], and neural architecture search [61], which often requires separate training runs for each configuration.
-The hyper-parameters’ space exploration problem is exacerbated by the steep increase in size of recent neural architectures for language  [9] and vision  [68], for example, where parameters are in the order of billions to (supposedly) accommodate for a huge dataset.
+The hyper-parameters’ space exploration problem is exacerbated by the steep increase in size of recent neural architectures for language [9] and vision [68], for example, where parameters are in the order of billions to accommodate for a huge dataset.
 Training these models requires an amount of time, compute power, and energy that currently makes it unfeasible for most institutions to perform a thorough model selection and find good width parameters; the commonly accepted compromise is to stick to previously successful hyper-parameter choices.
@@
 This work introduces a simple and easy to use technique to learn the width of each neural network’s layer without imposing upper bounds (we refer to it as unbounded width).
-The width of each layer is dynamically adjusted during backpropagation using the latest deep learning libraries  [44], and it only requires a slight modification to the neural activations that does not alter the ability to parallelize computation.
-The technical strategy is to impose a soft ordering of hidden units by exploiting any monotonically decreasing function with unbounded support on natural numbers, which has a number of benefits.
-For instance, we do not need to fix a maximum number of neurons, which is typical of orthogonal approaches like supernetworks  [61].
-As a by-product, we can achieve a straightforward trade-off between parametrization and performance by removing the last rows/columns of weight matrices, which correspond to the “least important” neurons in the computation.
-Finally, we break symmetries in the parametrization of neural networks: it is not possible anymore to obtain an equivalent neural network behavior by permuting the weight matrices, which should in principle reduce the “jostling” effect where symmetric parametrizations compete when training starts  [1].
+The width of each layer is dynamically adjusted during backpropagation [44], and it only requires a slight modification to the neural activations that does not alter the ability to parallelize computation.
+The technical strategy is to impose a soft ordering of hidden units by exploiting any monotonically decreasing function with unbounded support on natural numbers.
+With this, we do not need to fix a maximum number of neurons, which is typical of orthogonal approaches like supernetworks [61].
+As a by-product, we can achieve a straightforward trade-off between parametrization and performance by deleting the last rows/columns of weight matrices, i.e., removing the “least important” neurons from the computation.
+Finally, we break symmetries in the parametrization of neural networks: it is not possible anymore to obtain an equivalent neural network behavior by permuting the weight matrices, which reduces the “jostling” effect where symmetric parametrizations compete when training starts [1].
 We test our method on MLPs for tabular data, a Convolutional Neural Network (CNN) [33] for images, a Transformer architecture [58] for text, a Recurrent Neural Network (RNN) for sequences, for and a Deep Graph Network (DGN) [38, 52] for graphs, to showcase a broad scope of applicability.
@@
 samples, with input $x\in\mathbb{R}^{F},F\in\mathbb{N}^{+}$ and target $y$ whose domain depends on whether the task is regression or classification.
-For samples $X\in\mathbb{R}^{N\times F}$ and targets $Y$, the learning objective is to maximize the likelihood
+For samples $X\in\mathbb{R}^{N\times F}$ and targets $Y$, the learning objective is to maximize the log-likelihood
 $\log p(Y|X)=\log\prod_{i=1}^{N}p(y_{i}|x_{i})=\sum_{i=1}^{N}\log p(y_{i}|x_{i})$ (1)
+with respect to the learnable parameters of $p(y|x)$.
 […]
+We then decompose the joint distribution using the independence assumptions of the graphical model:
 $p(Y,\boldsymbol{\lambda},\boldsymbol{\theta}|X)=\prod_{i=1}^{N}p(y_{i},\boldsymbol{\lambda},\boldsymbol{\theta}|x_{i})\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ p(y_{i},\boldsymbol{\lambda},\boldsymbol{\theta}|x_{i})=p(y_{i}|\boldsymbol{\lambda},\boldsymbol{\theta},x_{i})p(\boldsymbol{\lambda})p(\boldsymbol{\theta})$ (2)
@@
 In other words, we find the integer such that the cumulative mass function of $f_{\ell}$ takes value $k$, and that integer $D_{\ell}$ is the truncated width of layer $\ell$.
-W.l.o.g., we implement $f_{\ell}$ as a discretized exponential distribution adhering to Def.
+We implement $f_{\ell}$ as a discretized exponential distribution adhering to Def.
 A.2, following the discretization strategy of Roy [49]: For every natural $x$, the discretized distribution relies on the exponential’s cumulative distribution function:
@@
 where distributions’ parameters are made explicit to distinguish them.
-The first two (optional) terms in the ELBO regularize the width of the layers and the magnitude of the parameters, respectively, whereas the third term accounts for the predictive performance.
+The first two terms in the ELBO regularize the width of the layers and the magnitude of the parameters when priors are informative, whereas the third term accounts for the predictive performance.
 In practice, the finite variational parameters $\boldsymbol{\nu},\boldsymbol{\rho}$ are those used to parametrize the neural network rather than sampling $\boldsymbol{\lambda},\boldsymbol{\theta}$, which enables easy optimization via backpropagation.
@@
 From a Bayesian perspective, this is necessary as regularizers should weigh less if we have more data.
-Finally, note that the first-order approximation leaves $\boldsymbol{\nu}$ and $\boldsymbol{\rho}$ completely free to change if not for the regularizers, that is, there is potentially no restriction imposed by the Gaussian assumptions.
-Compared to a fixed-width network with weight decay, we need to choose the values of $\mu^{\lambda}_{\ell},\sigma^{\lambda}_{\ell}$, as well as initialize the learnable $\nu_{\ell}$.
+Compared to a fixed-width network with weight decay, we need to choose the priors’ values of $\mu^{\lambda}_{\ell},\sigma^{\lambda}_{\ell}$, as well as initialize the learnable $\nu_{\ell}$.
 The latter can be initially set to same value across layers since they can freely adapt later, or it can be sampled from the prior $p(\lambda_{\ell})$.
-Therefore, we have two more hyper-parameters compared to the fixed-width network, but we make some considerations: i) it is always possible to use an uninformative prior over $\lambda_{\ell}$, removing the extra hyper-parameters letting the model freely adapt the width of each layer (as is typical of frequentist approaches); ii) the choice of higher level of hyper-parameters is known to be less stringent than that of hyper-parameters themselves  [21, 5], so we do not need to explore many values of $\mu_{\ell}^{\lambda}$ and $\sigma_{\ell}^{\lambda}$; iii) our experiments suggest that AWNN can converge to similar widths regardless of the starting point $\nu_{\ell}$, so that we may just need to perform model selection over one/two sensible initial values.
+Therefore, we have two more hyper-parameters compared to the fixed-width network, but we make some considerations: i) it is always possible to use an uninformative prior over $\lambda_{\ell}$, removing the extra hyper-parameters letting the model freely adapt the width of each layer (as is typical of frequentist approaches); ii) the choice of higher level of hyper-parameters is known to be less stringent than that of hyper-parameters themselves [21, 5], so we do not need to explore many values of $\mu_{\ell}^{\lambda}$ and $\sigma_{\ell}^{\lambda}$; iii) our experiments suggest that AWNN can converge to similar widths regardless of the starting point $\nu_{\ell}$, so that we may just need to perform model selection over one/two sensible initial values; iv) the more data, the less the priors will matter.
 3.1 Imposing a Soft Ordering on Neurons’ Importance
 Now that the learning objective has been formalized, the missing ingredient is the definition of the neural network $p(y_{i}|\boldsymbol{\lambda}\text{=}\boldsymbol{\nu},\boldsymbol{\theta}\text{=}\boldsymbol{\rho},x_{i})$ of Equation 4 as a modified MLP.
-Compared to a standard MLP, we need to make use of the variational parameters $\boldsymbol{\nu}$ that affect the truncation width at each hidden layer, whereas $\boldsymbol{\rho}$ are the weights.
-We chose a monotonically decreasing function $f_{\ell}$, thus when a new neuron is added its relative importance is low and will not drastically impact the network output and hidden activations.
+Compared to a standard MLP, we use the variational parameters $\boldsymbol{\nu}$ that affect the truncation width at each hidden layer, whereas $\boldsymbol{\rho}$ are the weights.
+We choose a monotonically decreasing function $f_{\ell}$, thus when a new neuron is added its relative importance is low and will not drastically impact the network output and hidden activations.
 In other words, we impose a soft ordering of importance among neurons.
@@
 That is, we rescale the activation of each neuron $k$ by its “importance" $f_{\ell}(j;\nu_{\ell})$.
-Note that the bias parameter can be part of the weight vector as usual.
+Note that the bias parameter is part of the weight vector as usual.
 It is easy to see that, in theory, the optimization algorithm could rescale the weights of the next layer by a factor $1/f_{\ell}(j;\nu_{\ell})$ to compensate for the term $f_{\ell}(j;\nu_{\ell})$.
@@
 The first is to regularize the magnitude of the weights thanks to the prior $p(\theta_{\ell+1n})$, so that it may be difficult to compensate for the least important neurons that have a high $1/f_{\ell}(j)$.
-The second and less obvious strategy is to prevent the units’ activations of the current layer to compensate for high values by bounding their range, e.g., using a ReLU6 activation  [51].
+The second and less obvious strategy is to prevent the units’ activations of the current layer to compensate for high values by bounding their range, e.g., using a ReLU6 or tanh activation [51].
 We apply both strategies to our experiments, although we noted that they do not seem strictly necessary in practice.
@@
 Rescaling the activations of hidden units using Equation 10 causes activations of deeper layers to quickly decay to zero for an AWNN MLP with ReLU nonlinearity initialized using the well known Kaiming scheme [25].
-This affects convergence since gradients get close to zero and it becomes impractical to train deep AWNN MLPs.
+This affects convergence since gradients get close to zero and it becomes slow to train deep AWNN MLPs.
 We therefore derive a rescaled Kaiming weight that guarantees that the variance of activation across layers is constant at initialization.
@@
 Therefore, if we initialize weights from a Gaussian distribution with standard deviation $\frac{\sqrt{2}}{\sqrt{\sum_{j=1}^{D_{\ell-1}}f^{2}_{\ell}(j)}}$, we guarantee that at initialization the variance of the deep network’s gradients will be constant.
-The effect of the new initialization (dubbed “Kaiming+”) can be seen in Figure
+The effect of the new initialization (dubbed “Kaiming+”) can be seen in Figure 7, where the distribution of activation values at initialization does not collapse in subsequent layers.
+This change drastically impacts overall convergence on the SpiralHard synthetic dataset (described in Section 4), where it appears it would be otherwise hard to converge using a standard Kaiming initialization.
 Algorithm 1 AWNN Training Procedure
@@
 12:       Use $D_{\ell}$ to update $\boldsymbol{\rho}_{\ell},\boldsymbol{\rho}_{\ell+1}$ // add/remove neurons
-7, where the distribution of activation values at initialization does not collapse in subsequent layers.
-This change drastically impacts overall convergence on the SpiralHard synthetic dataset (described in Section 4), where it appears it would be otherwise hard to converge using a standard Kaiming initialization.
 Algorithm 1 summarizes the main changes to the training procedure, namely the new initialization and the update of the model’s truncated width at each training step.
@@
 AWNN needs to change the network at training time with an associated overhead.
-However, our current implementation has up to 3-4$\times$ overhead on the smallest datasets, where the inference cost is negligible compared to the network update.
-Although this should not be an issue with more realistic tasks, especially if one considers the time spent testing different widths during grid search, further optimizations are conceivable: i) by updating the width every $M$ mini-batches rather than at every step; ii) by allocating a larger network and dynamically grow it only if needed; iii) by creating ad-hoc support of AWNN on deep learning libraries like PyTorch.
+Our current implementation has only a 3-4$\times$ overhead on the smallest datasets, the worst case scenario for our approach, where the inference cost is negligible compared to the network update.
+For more realistic tasks, we believe our method is advantageous compared to the cost of a grid-search or alternating training/updating approaches.
+In addition, further optimizations are possible: i) by updating the width every $M$ mini-batches rather than at every step; ii) by allocating a larger network and dynamically grow it only if needed; iii) by creating ad-hoc support of AWNN on deep learning libraries.
 4 Experiments and Setup
 The purpose of the empirical analysis is not to claim AWNN is generally better than the fixed-width baseline.
-Rather, we demonstrate how AWNN overcomes the problem of fixing the number of neurons by learning it, thus reducing the amount of configurations to test.
+Rather, we demonstrate how AWNN overcomes the problem of fixing the number of neurons by learning it end-to-end, thus reducing the amount of hyper-parameter configurations to test.
 As such, due to the nature of this work and similarly to Mitchell et al.
@@
 Due to space reasons, we report datasets statistics and the hyper-parameter tried for the fixed and AWNN versions in Appendix F and G, respectively [^3] Code to reproduce results is in the supplementary material..
-We ran the experiments on a server with 128 cores, 1.5TB of RAM, and 4 NVIDIA A40 with 46GB of memory.
+We ran the experiments on a server with 64 cores, 1.5TB of RAM, and 4 NVIDIA A40 with 48GB of memory.
 5 Results
@@
 SpiralHard 98.0 (2.0) 100.0 (0.0) 32 227.4 (32.4)
+PMNIST 91.1 (0.4) 95.7 (0.2) 24 806.3 (44.5)
 MNIST 99.6 (0.1) 99.7 (0.0) Linear 19.4 (4.8)
@@
 CIFAR100 66.5 (0.4) 63.1 (4.0) 256 161.9 (57.8)
-PMNIST 91.1 (0.4) 95.7 (0.2) 24 806.3 (44.5)
 NCI1 80.0 (1.4) 80.0 (1.1) (96-320) 731.3 (128.2)
@@
 We begin by discussing the quantitative results of our experiments: Table 1 reports means and standard deviations across the 10 final training runs.
-In terms of performance, we observe that AWNN is more stable or accurate than a fixed MLP on DoubleMoon, Spiral, and SpiralHard; all other things being equal, it seems that using more neurons and their soft ordering are the main contributing factors to these improvements.
+In terms of performance, we observe that AWNN is more stable or accurate than a fixed MLP on DoubleMoon, Spiral, SpiralHard, and PMNIST; all other things being equal, it seems that using more neurons and their soft ordering are the main contributing factors to these improvements.
 On the image datasets, performances of AWNN are comparable to those of the fixed baseline but for CIFAR100, due to an unlucky run that did not converge.
 There, AWNN learns a smaller total width compared to grid search.
-Results on graph datasets are interesting in two respects: First, the performance on REDDIT-B is significantly improved by AWNN both in terms of average performance and stability of results; second, the total learned width is significantly higher than those tried in  Xu et al.
+Results on graph datasets are interesting in two respects: First, the performance on REDDIT-B is significantly improved by AWNN both in terms of average performance and stability of results; second, and akin to PMNIST, the total learned width is significantly higher than those tried in Xu et al.
 [65], Errica et al.
-[16], meaning that a biased choice of possible widths had a profound influence on risk estimation for DGN models (i.e., GIN).
+[16], meaning a biased choice of possible widths had a profound influence on risk estimation for DGN models (i.e., GIN).
 This result makes it evident that it is important to let the network decide how many neurons are necessary to solve the task.
-Appendix H shows what happens when we retrain some fixed baselines using the total width as the width of each layer.
+Appendix H shows what happens when we retrain Fixed baselines using the total width as the width of each layer.
 Finally, the results on the Multi30k show that the AWNN Transformer learns to use 200x parameters less than the fixed Transformer for the feed-forward networks, achieving a statistically comparable test loss.
@@
 Proposition D.1 (AWNN ELBO Lipschitz continuity).
-The ELBO loss of Equation 9, with respect to the change in the depth $D_{\ell}$ for the layer $l$, is Lipschitz continuous.
+The ELBO loss of Equation 9, with respect to the change in the depth $D_{\ell}$ for the layer $\ell$, is Lipschitz continuous.
 Proof.
@@
 $|f_{1}(D_{\ell^{\prime}})-f_{1}(D_{\ell})|$ $=|\sum_{n=D_{\ell}}^{D_{\ell^{\prime}}}\log\frac{p(\boldsymbol{\rho}_{n})}{q(\boldsymbol{\rho}_{n}|\boldsymbol{\nu})}|$ (17)
-$\leq~\sum_{n=D_{l}}^{D_{\ell^{\prime}}}|\log\frac{p(\boldsymbol{\rho}_{n})}{q(\boldsymbol{\rho}_{n}|\boldsymbol{\nu})}|$ (18)
+$\leq~\sum_{n=D_{\ell}}^{D_{\ell^{\prime}}}|\log\frac{p(\boldsymbol{\rho}_{n})}{q(\boldsymbol{\rho}_{n}|\boldsymbol{\nu})}|$ (18)
 $\leq\max_{n}|\log\frac{p(\boldsymbol{\rho}_{n})}{q(\boldsymbol{\rho}_{n}|\boldsymbol{\nu})}||D_{\ell^{\prime}}-D_{\ell}|$ (19)
@@
 CIFAR100 66.5(0.4) 64.9(1.1) 64.9 65.9 63.1(4.0)
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: Abstract and introduction reflect our results.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: Section 3.3 discusses limitations and future extensions.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory assumptions and proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [Yes]
-  14. Justification: All theorems in the appendix come with a complete proof with the necessary assumptions.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental result reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19.
-Justification: The paper provides experimental details about the datasets, data splits, and hyper-parameters tried in the appendix.
-Also, the experiments are easy to reproduce thanks to the adoption of ad-hoc libraries in the supplementary material.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24. Justification: Yes, see also our previous answer.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental setting/details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: Yes, see also our previous answers.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment statistical significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes]
-  34. Justification: See Table 1.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments compute resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: See Section 4.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code of ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: Our approach is a general machine learning methodology that could be used in unforeseen ways. We did not experiment on human subjects and the data is publicly available.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [N/A]
-  49. Justification: No societal impact of the work performed.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54. Justification: No such risks.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [No]
-  59. Justification: We do rely on Python packages but did not acknowledge each of them in our paper.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [Yes]
-  64. Justification: See the supplementary material.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and research with human subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: No research on human subjects.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: No research on human subjects.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
-  76. 16.
-    Declaration of LLM usage
-  77.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-  78. Answer: [N/A]
-  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
-  80. Guidelines:
-    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
-    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v5

@@
 Abstract
-For almost 70 years, researchers have primarily relied on hyper-parameter tuning to select the width of neural networks’ layers.
+For almost 70 years, researchers have typically selected the width of neural networks’ layers either manually or through automated hyperparameter tuning methods such as grid search and, more recently, neural architecture search.
 This paper challenges the status quo by introducing an easy-to-use technique to learn an unbounded width of a neural network’s layer during training.
@@
 1 Introduction
-Since the construction of the Mark I Perceptron machine [48] the effective training of neural networks has remained an open research problem of great academic and practical value.
-The Mark I solved image recognition tasks by exploiting a layer of 512 fixed “association units” that in modern language correspond to the hidden units of a Multi-Layer Perceptron (MLP).
-MLPs possess universal approximation capabilities when assuming arbitrary width [10] and sigmoidal activations, and their convergence to good solutions was studied, for instance, in Rumelhart et al.
-[50] where the backpropagation algorithm was introduced as “simple, easy to implement on parallel hardware”, and improvable by other techniques such as momentum that preserve locality of weight updates.
-Yet, after almost 70 years of progress [34], the vast majority of neural networks, be they shallow or deep, still rely on a fixed choice of the number of neurons in their hidden layers.
-The width is typically treated as one of the many hyper-parameters that have to be carefully tuned whenever we approach a new task [62].
-The tuning process has many names, such as model selection, hyper-parameter tuning, and cross-validation, and it is associated with non-negligible costs: Different architectural configurations are trained until one that performs best on a validation set is selected [41].
-The configurations’ space grows exponentially in the number of layers, so practitioners often resort to shortcuts such as picking a specific number of hidden units for all layers, which greatly reduces the search space together with the chances of selecting a better architecture for the task.
-Other techniques to tune hyper-parameters include constructive approaches [17, 64], which alternate parameter optimization and creation of new neurons, natural gradient-based heuristics that dynamically modify the network [40], bi-level optimization [19], and neural architecture search [61], which often requires separate training runs for each configuration.
-The hyper-parameters’ space exploration problem is exacerbated by the steep increase in size of recent neural architectures for language [9] and vision [68], for example, where parameters are in the order of billions to accommodate for a huge dataset.
+Since the construction of the Mark I Perceptron machine (Rosenblatt, 1958) the effective training of neural networks has remained an open research problem of great academic and practical value.
+The Mark I solved image recognition tasks with a layer of 512 fixed “association units” that in modern language are the hidden units of a Multi-Layer Perceptron (MLP).
+MLPs possess universal approximation capabilities when assuming arbitrary width (Cybenko, 1989) and sigmoidal activations, and their convergence to good solutions was studied, for instance, in Rumelhart et al.
+(1986) where the backpropagation algorithm was introduced as “simple, easy to implement on parallel hardware”, and improvable by other techniques such as momentum that preserve locality of weight updates.
+Yet, after almost 70 years of progress (LeCun et al., 2015), the vast majority of neural networks, be they shallow or deep, still rely on a fixed choice of the number of neurons in their hidden layers.
+The width is typically treated as one of the many hyper-parameters that have to be carefully tuned whenever we approach a new task (Wolpert, 1996).
+The tuning process has many names, such as model selection, hyper-parameter tuning, and cross-validation, and it is associated with non-negligible costs: Different architectural configurations are trained until one that performs best on a validation set is selected (Mitchell, 1997).
+The configurations’ space grows exponentially in the number of layers, so practitioners often resort to shortcuts such as picking a specific number of hidden units for all layers and trying a few values, which greatly reduces the search space together with the chances of selecting a better architecture for the task.
+Other techniques to tune hyper-parameters include constructive approaches (Fahlman & Lebiere, 1989; Wu et al., 2020), which alternate parameter optimization and creation of new neurons, natural gradient-based heuristics that dynamically modify the network (Mitchell et al., 2023), bi-level optimization (Franceschi et al., 2018), and neural architecture search (White et al., 2023), which often requires separate training runs for each configuration.
+The hyper-parameters’ space exploration problem is exacerbated by the steep increase in size of recent neural architectures for language (Brown et al., 2020) and vision (Zhai et al., 2022), for example, where parameters are in the order of billions to accommodate for a huge dataset.
 Training these models requires an amount of time, compute power, and energy that currently makes it unfeasible for most institutions to perform a thorough model selection and find good width parameters; the commonly accepted compromise is to stick to previously successful hyper-parameter choices.
-This may also explain why network pruning [6, 39], distillation [69] and quantization [39] techniques have recently been in the spotlight, as they trade-off hardware requirements and performance.
+This may also explain why network pruning (Blalock et al., 2020; Mishra et al., 2021), distillation (Zhang et al., 2019) and quantization (Mishra et al., 2021) techniques have recently been in the spotlight, as they trade-off hardware requirements and performance.
 This work introduces a simple and easy to use technique to learn the width of each neural network’s layer without imposing upper bounds (we refer to it as unbounded width).
-The width of each layer is dynamically adjusted during backpropagation [44], and it only requires a slight modification to the neural activations that does not alter the ability to parallelize computation.
+The width of each layer is dynamically adjusted during backpropagation (Paszke et al., 2017), and it only requires a slight modification to the neural activations that does not alter the ability to parallelize computation.
 The technical strategy is to impose a soft ordering of hidden units by exploiting any monotonically decreasing function with unbounded support on natural numbers.
-With this, we do not need to fix a maximum number of neurons, which is typical of orthogonal approaches like supernetworks [61].
+With this, we do not need to fix a maximum number of neurons, which is typical of orthogonal approaches like supernetworks (White et al., 2023).
 As a by-product, we can achieve a straightforward trade-off between parametrization and performance by deleting the last rows/columns of weight matrices, i.e., removing the “least important” neurons from the computation.
-Finally, we break symmetries in the parametrization of neural networks: it is not possible anymore to obtain an equivalent neural network behavior by permuting the weight matrices, which reduces the “jostling” effect where symmetric parametrizations compete when training starts [1].
-We test our method on MLPs for tabular data, a Convolutional Neural Network (CNN) [33] for images, a Transformer architecture [58] for text, a Recurrent Neural Network (RNN) for sequences, for and a Deep Graph Network (DGN) [38, 52] for graphs, to showcase a broad scope of applicability.
-Empirical results show, as may be intuitively expected, that the width adapts to the task’s difficulty with the same performance of a fixed-width baseline.
-We also investigate compressing of the network at training time while preserving accuracy, as well as a post-hoc truncation inducing a controlled trade-off at zero additional cost.
+Finally, we break symmetries in the parametrization of neural networks: it is not possible anymore to obtain an equivalent neural network behavior by permuting the weight matrices, which reduces the “jostling” effect where symmetric parametrizations compete when training starts (Barber, 2012).
+We test our method on MLPs for tabular data, a Convolutional Neural Network (CNN) (LeCun et al., 1989) for images, a Transformer (Vaswani et al., 2017) for text, a Recurrent Neural Network (RNN) for sequences, and a Deep Graph Network (DGN) (Micheli, 2009; Scarselli et al., 2009) for graphs, showcasing a broad scope of applicability as MLPs are ubiquitous in modern architectures.
+Empirical results show, as may be intuitively expected, that the width adapts to the task’s difficulty with performances comparable to fixed-width baseline.
+We also encourage compression of the network at training time while preserving accuracy, as well as a post-hoc truncation inducing a controlled trade-off at zero additional cost.
 Ablations suggest that the learned width is not influenced by the starting width (under bounded activations) nor by the batch size, advocating for a potentially large reduction of the hyper-parameter configuration space.
@@
 Constructive methods dynamically learn the width of neural networks and are related in spirit to this work.
-The cascade correlation algorithm [17] alternates standard training with the creation of a new hidden unit minimizing the neural network’s residual error.
-Similarly, the firefly network descent [64] grows the width and depth of a network every $N$ training epochs via gradient descent on a dedicated loss.
+The cascade correlation algorithm (Fahlman & Lebiere, 1989) alternates standard training with the creation of a new hidden unit minimizing the neural network’s residual error.
+Similarly, the firefly network descent (Wu et al., 2020) grows the width and depth of a network every $N$ training epochs via gradient descent on a dedicated loss.
 Yoon et al.
-[66] propose an ad-hoc algorithm for lifelong learning that grows the network by splitting and duplicating units to learn new tasks.
+(2018) propose an ad-hoc algorithm for lifelong learning that grows the network by splitting and duplicating units to learn new tasks.
 Wu et al.
-[63] alternate training the network and then splitting existing neurons into offspring with equal weights.
+(2019) alternate training the network and then splitting existing neurons into offspring with equal weights.
 These works mostly focus on growing the neural network; Mitchell et al.
-[40] propose natural gradient-based heuristics to grow/shrink layers and hidden units of MLPs and CNNs.
+(2023) propose natural gradient-based heuristics to grow/shrink layers and hidden units of MLPs and CNNs.
 The main difference from our work is that we grow and shrink the network by simply computing the gradient of the loss, without relying on human-defined heuristics.
-The unbounded depth network of Nazaret and Blei [42], from which we draw inspiration, learns the number of layers of neural networks.
+The unbounded depth network of Nazaret & Blei (2022), from which we draw inspiration, learns the number of layers of neural networks.
 Compared to that work, we focus our attention to the number of neurons, modifying the internals of the architecture rather than instantiating a multi-output one.
-Finally, we mention Bayesian nonparametrics approaches [43] that learn a potentially infinite number of clusters in an unsupervised fashion, as well as dynamic neural networks [24] that condition the architecture on input properties.
+In particular: i) learning the width requires a different formulation of the evidence lower-bound; and ii) the importance distribution needs to be monotonically decreasing, which does not encode the right inductive bias for the depth of a neural network.
+We also mention Bayesian nonparametric approaches (Orbanz & Teh, 2010) that learn a potentially infinite number of clusters in an unsupervised fashion, as well as dynamic neural networks (Han et al., 2021) that condition the architecture on input properties.
+Finally, the work of Caron et al.
+(2025) investigates, in a similar spirit to this work, how to introduce decreasing rescaling of the pre-activations of neurons in the context of very wide networks; the authors provide interesting global convergence results in the Neural Tangent Kernel (NTK) regime.
 Orthogonal Methods
-Neural Architecture Search (NAS) is an automated process that designs neural networks for a given task [15, 61] and has been applied to different contexts [71, 35, 53].
-Typically, neural network elements are added, removed, or modified based on validation performance [14, 60, 61], by means of reinforcement learning [70], evolutionary algorithms [47], and gradient-based approaches [35].
-Typical NAS methods require enormous computational resources, sometimes reaching thousands of GPU days [70], due to the retraining of each new configuration.
-While recent advances on one-shot NAS models [8, 45, 2, 4, 55, 54] have drastically reduced the computational costs, they mostly focus on CNNs, assume a bounded search space, and do not learn the width.
+Neural Architecture Search (NAS) is an automated process that designs neural networks for a given task (Elsken et al., 2019b; White et al., 2023) and has been applied to different contexts (Zoph et al., 2018; Liu et al., 2019; So et al., 2019).
+Typically, neural network elements are added, removed, or modified based on validation performance (Elsken et al., 2019a; White et al., 2021; 2023), by means of reinforcement learning (Zoph & Le, 2016), evolutionary algorithms (Real et al., 2019), and gradient-based approaches (Liu et al., 2019).
+Typical NAS methods require enormous computational resources, sometimes reaching thousands of GPU days (Zoph & Le, 2016), due to the retraining of each new configuration.
+While recent advances on one-shot NAS models (Brock et al., 2018; Pham et al., 2018; Bender et al., 2018; Berman et al., 2020; Su et al., 2021b; a) have drastically reduced the computational costs, they mostly focus on CNNs, assume a bounded search space, and do not learn the width.
 As such, NAS methods are complementary to our approach.
-Bi-level optimization algorithms have also been used for hyper-parameter tuning [19], where hyper-parameters are the variables of the outer objective and the model parameters those of the inner objective.
-The solution sets of the inner problem are usually not available in closed form, which has been partly addressed by repeated application of (stochastic) gradient descent [11, 36, 18].
-These methods are restricted to continuous hyper-parameters’ optimization, and cannot therefore be applied to width optimization.
-Finally, pruning [6] and distillation [29] are two methods that reduce the size of neural networks by trading-off performances; the former deletes neural connections [39] or entire neurons [57, 12], the latter trains a smaller network (student) to mimic a larger one (teacher) [22].
-In particular, dynamic pruning techniques can compress the network at training time [23], by applying hard or soft masks [28]; for a comprehensive survey on pruning, please refer to [27].
+Bi-level optimization algorithms have also been used for hyper-parameter tuning (Franceschi et al., 2018), where hyper-parameters are the variables of the outer objective and the model parameters those of the inner objective.
+The solution sets of the inner problem are usually not available in closed form, which has been partly addressed by repeated application of (stochastic) gradient descent (Domke, 2012; Maclaurin et al., 2015; Franceschi et al., 2017).
+These methods are restricted to continuous hyper-parameters’ optimization and cannot be applied to width optimization.
+Finally, pruning (Blalock et al., 2020) and distillation (Hinton et al., 2015) are two methods that reduce the size of neural networks by trading-off performances; the former deletes neural connections (Mishra et al., 2021) or entire neurons (Valerio et al., 2022; Dufort-Labbé et al., 2024), the latter trains a smaller network (student) to mimic a larger one (teacher) (Gou et al., 2021).
+In particular, dynamic pruning techniques can compress the network at training time (Guo et al., 2016), by applying hard or soft masks (He et al., 2018); for a comprehensive survey on pruning, please refer to (He & Xiao, 2023).
+It is worth mentioning the pruning strategy of Wolinski et al.
+(2020), which promotes learning in some neurons and penalizes others by appropriate rescaling of a potentially (infinite-dimensional) input; this work shares conceptual similarities with the our approach, though it does not consider learning the rescaling factor as a proxy for the width.
 Compared to most pruning approaches, our work can delete connections and reduce the model’s memory, but also grow it indefinitely; compared to distillation, we do not necessarily need a new training to compress the network.
-These techniques, however, can be easily combined with our approach.
+These techniques can be easily combined with our approach, whose main goal is not compression but rather the automatic adaptation of a neural network’s width.
 3 Adaptive Width Learning
-We now introduce Adaptive Width Neural Networks (AWNN), a probabilistic framework that maximizes a simple variational objective via backpropagation over a neural network’s parameters.
+We now introduce Adaptive Width Neural Networks (AWN), a probabilistic framework that maximizes a simple variational objective via backpropagation over a neural network’s parameters.
 We are given a dataset of $N$ i.i.d.
@@
 with respect to the learnable parameters of $p(y|x)$.
-[…]
+We define $p(y|x)$ according to the graphical model of Figure 1 (left), to learn a neural network with unbounded width for each hidden layer $\ell$.
+To do so, we assume the existence of an infinite sequence of i.i.d.
+latent variables $\bm{\theta}_{\ell}=\left\{\theta_{\ell n}\right\}_{n=1}^{\infty}$, where $\theta_{\ell n}$ is a multivariate variable over the learnable weights of neuron $n$ at layer $\ell.$ However, since working with an infinite-width layer is not possible in practice, we also introduce a latent variable $\lambda_{\ell}$ that samples how many neurons to use for layer $\ell$.
+That is, it truncates an infinite width to a finite value so that we can feasibly perform inference with the neural network.
+For a neural network of $L$ layers, we define $\bm{\theta}=\left\{\bm{\theta}_{\ell}\right\}_{\ell=1}^{L}$ and $\bm{\lambda}=\left\{\lambda_{\ell}\right\}_{\ell=1}^{L}$, assuming independence across layers.
+Therefore, by marginalization one can write $p(Y|X)=\int p(Y,\bm{\lambda},\bm{\theta}|X)d\bm{\lambda}d\bm{\theta}$.
 We then decompose the joint distribution using the independence assumptions of the graphical model:
-$p(Y,\boldsymbol{\lambda},\boldsymbol{\theta}|X)=\prod_{i=1}^{N}p(y_{i},\boldsymbol{\lambda},\boldsymbol{\theta}|x_{i})\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ p(y_{i},\boldsymbol{\lambda},\boldsymbol{\theta}|x_{i})=p(y_{i}|\boldsymbol{\lambda},\boldsymbol{\theta},x_{i})p(\boldsymbol{\lambda})p(\boldsymbol{\theta})$ (2)
-$p(\boldsymbol{\lambda})=\prod^{L}_{\ell=1}p(\lambda_{\ell})=\prod^{L}_{\ell=1}\mathcal{N}(\lambda_{\ell};\mu^{\lambda}_{\ell},\sigma^{\lambda}_{\ell})\ \ \ \ \ \ p(\boldsymbol{\theta})=\prod^{L}_{\ell=1}\prod^{\infty}_{n=1}p(\theta_{\ell n})\text{=}\prod^{L}_{\ell=1}\prod^{\infty}_{n=1}\mathcal{N}(\theta_{\ell n};\mathbf{0},\text{diag}(\sigma^{\theta}_{\ell}))$ (3)
+$p(Y,\bm{\lambda},\bm{\theta}|X)=\prod_{i=1}^{N}p(y_{i},\bm{\lambda},\bm{\theta}|x_{i})\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ p(y_{i},\bm{\lambda},\bm{\theta}|x_{i})=p(y_{i}|\bm{\lambda},\bm{\theta},x_{i})p(\bm{\lambda})p(\bm{\theta})$ (2)
+$p(\bm{\lambda})=\prod^{L}_{\ell=1}p(\lambda_{\ell})=\prod^{L}_{\ell=1}\mathcal{N}(\lambda_{\ell};\mu^{\lambda}_{\ell},\sigma^{\lambda}_{\ell})\ \ \ \ \ \ p(\bm{\theta})=\prod^{L}_{\ell=1}\prod^{\infty}_{n=1}p(\theta_{\ell n})\text{=}\prod^{L}_{\ell=1}\prod^{\infty}_{n=1}\mathcal{N}(\theta_{\ell n};\mathbf{0},\text{diag}(\sigma^{\theta}_{\ell}))$ (3)
 […]
 […]
-Figure 1: (Left) The graphical model of AWNN, with dark observable random variables and white latent ones.
+Figure 1: (Left) The graphical model of AWN, with dark observable random variables and white latent ones.
 (Middle) The distribution $f_{\ell}$ over hidden units’ importance at layer $\ell$ is parametrized by $\lambda_{\ell}$.
@@
 (Right) The hidden units’ activations at layer $\ell$ are rescaled by their importance.
-[…]
-$\log p(Y|X)\geq\mathbb{E}_{q(\boldsymbol{\lambda},\boldsymbol{\theta})}\left[\log\frac{p(Y,\boldsymbol{\lambda},\boldsymbol{\theta}|X)}{q(\boldsymbol{\lambda},\boldsymbol{\theta})}\right],$ (5)
-where $q(\boldsymbol{\lambda},\boldsymbol{\theta})$ is parametrized by learnable variational parameters.
+where $\sigma^{\theta}_{\ell},\mu^{\lambda}_{\ell},\sigma^{\lambda}_{\ell}$ are hyper-parameters.
+For simplicity, we assumed a Gaussian prior for $\bm{\lambda}$, but please note that, when strict positivity is required, one should formally turn to Folded Normal priors.
+The neural network is parametrized by realizations $\bm{\lambda}\sim p(\bm{\lambda}),\bm{\theta}\sim\bm{\theta}$ – it relies on a finite number of neurons as detailed later and in Section 3.1 – and it outputs either class probabilities (classification) or the mean of a Gaussian distribution (regression) to parametrize $p(y_{i}|\bm{\lambda},\bm{\theta},x_{i})$ depending on the task.
+Maximizing Equation (1), however, requires computing the evidence $\int p(Y,\bm{\lambda},\bm{\theta}|X)d\bm{\lambda}d\bm{\theta}$, which is intractable.
+Therefore, we turn to mean-field variational inference (Jordan et al., 1999; Blei et al., 2017) to maximize an expected lower bound (ELBO) instead.
+This requires to define a variational distribution over the latent variables $q(\bm{\lambda},\bm{\theta})$ and re-phrase the objective as:
+$\log p(Y|X)\geq\mathbb{E}_{q(\bm{\lambda},\bm{\theta})}\left[\log\frac{p(Y,\bm{\lambda},\bm{\theta}|X)}{q(\bm{\lambda},\bm{\theta})}\right],$ (5)
+where $q(\bm{\lambda},\bm{\theta})$ is parametrized by learnable variational parameters.
 Before continuing, we define the truncated width $D_{\ell}$, that is the finite number of neurons at layer $\ell$, as the quantile function evaluated at $k$, with $k$ a hyper-parameter, of a distribution [^1] General functions are allowed if a threshold can be computed.
-$f_{\ell}$ with infinite support over $\mathbb{N}^{+}$ and parametrized by $\lambda_{\ell}$; Appendix A and B provide formal requirements about $f_{\ell}$.
+$f_{\ell}$ with infinite support over $\mathbb{N}^{+}$ and parametrized by $\lambda_{\ell}$; Appendix A provides desirable properties of $f_{\ell}$ in a similar vein to Nazaret & Blei (2022).
 In other words, we find the integer such that the cumulative mass function of $f_{\ell}$ takes value $k$, and that integer $D_{\ell}$ is the truncated width of layer $\ell$.
 We implement $f_{\ell}$ as a discretized exponential distribution adhering to Def.
-A.2, following the discretization strategy of Roy [49]: For every natural $x$, the discretized distribution relies on the exponential’s cumulative distribution function:
-$f_{\ell}(x;\lambda_{\ell})=(1-e^{\lambda_{\ell}(x+1)})-(1-e^{\lambda_{\ell}(x)}).$ (6)
+A.2, following the discretization strategy of Roy (2003): For every $x\in\mathbb{N}^{+}$, the discretized distribution relies on the exponential’s cumulative distribution function:
+$f_{\ell}(x;\lambda_{\ell})=(1-e^{-\lambda_{\ell}(x+1)})-(1-e^{-\lambda_{\ell}x}).$ (6)
 We choose the exponential because it is a monotonically decreasing function and allows us to impose an ordering of importance among neurons, as detailed in Section 3.1.
-Then, we can factorize the variational distribution $q(\boldsymbol{\lambda},\boldsymbol{\theta})$ into:
-$q(\boldsymbol{\lambda},\boldsymbol{\theta})=q(\boldsymbol{\lambda})q(\boldsymbol{\theta}|\boldsymbol{\lambda})$ $q(\boldsymbol{\lambda})=\prod_{\ell=1}^{L}q(\lambda_{\ell})=\prod_{\ell=1}^{L}\mathcal{N}(\lambda_{\ell};\nu_{\ell},1)$ (7)
-$q(\boldsymbol{\theta}|\boldsymbol{\lambda})=\prod_{\ell=1}^{L}\prod_{n=1}^{D_{\ell}}q(\theta_{\ell n})\prod_{D_{\ell}+1}^{\infty}p(\theta_{\ell n})$ $q(\theta_{\ell n})=\mathcal{N}(\theta_{\ell n};\text{diag}(\rho_{\ell n}),\mathbf{I}).$ (8)
-Here, $\nu_{\ell},\rho_{\ell n}$ are learnable variational parameters and, as before, we define $\boldsymbol{\rho}_{\ell}=\left\{\rho_{\ell n}\right\}_{n=1}^{D_{\ell}}$, $\boldsymbol{\rho}=\left\{\boldsymbol{\rho}_{\ell}\right\}_{\ell=1}^{L}$ and $\boldsymbol{\nu}=\left\{\nu_{\ell}\right\}_{\ell=1}^{L}$.
+Then, we can factorize the variational distribution $q(\bm{\lambda},\bm{\theta})$ into:
+$q(\bm{\lambda},\bm{\theta})=q(\bm{\lambda})q(\bm{\theta}|\bm{\lambda})$ $q(\bm{\lambda})=\prod_{\ell=1}^{L}q(\lambda_{\ell})=\prod_{\ell=1}^{L}\mathcal{N}(\lambda_{\ell};\nu_{\ell},1)$ (7)
+$q(\bm{\theta}|\bm{\lambda})=\prod_{\ell=1}^{L}\prod_{n=1}^{D_{\ell}}q(\theta_{\ell n})\prod_{n^{\prime}=D_{\ell}+1}^{\infty}p(\theta_{\ell n^{\prime}})$ $q(\theta_{\ell n})=\mathcal{N}(\theta_{\ell n};\rho_{\ell n},\mathbf{I}).$ (8)
+Here, $\nu_{\ell},\rho_{\ell n}$ are learnable variational parameters and, as before, we define $\bm{\rho}_{\ell}=\left\{\rho_{\ell n}\right\}_{n=1}^{D_{\ell}}$, $\bm{\rho}=\left\{\bm{\rho}_{\ell}\right\}_{\ell=1}^{L}$ and $\bm{\nu}=\left\{\nu_{\ell}\right\}_{\ell=1}^{L}$.
 Note that the set of variational parameters is finite as it depends on $D_{\ell}$.
-By expanding Equation 5 using the above definitions and approximating the expectations at the first order, i.e., $\mathbb{E}_{q(\boldsymbol{\lambda})}[f(\boldsymbol{\lambda})]$=$f(\boldsymbol{\nu})$ and $\mathbb{E}_{q(\boldsymbol{\theta}|\boldsymbol{\lambda})}[f(\boldsymbol{\theta})]=f(\boldsymbol{\rho})$ as in Nazaret and Blei [42], we obtain the final form of the ELBO [^2] The higher $k$ the better the ELBO approximation.
-(the full derivation is in Appendix C):
-$\sum_{\ell}^{L}\log\frac{p(\nu_{\ell};\mu_{\ell}^{\lambda},\sigma_{\ell}^{\lambda})}{q(\nu_{\ell};\nu_{\ell})}+\sum_{\ell}^{L}\sum_{n=1}^{D_{\ell}}\log\frac{p(\rho_{\ell n};\sigma_{\ell}^{\theta})}{q(\rho_{\ell n};\rho_{\ell n})}+\sum_{i=1}^{N}\log p(y_{i}|\boldsymbol{\lambda}\text{=}\boldsymbol{\nu},\boldsymbol{\theta}\text{=}\boldsymbol{\rho},x_{i}),$ (9)
+By expanding Equation 5 using the above definitions and approximating the expectations at the first order, i.e., $\mathbb{E}_{q(\bm{\lambda})}[f(\bm{\lambda})]$$f(\bm{\nu})$ and $\mathbb{E}_{q(\bm{\theta}|\bm{\lambda})}[f(\bm{\theta})]=f(\bm{\rho})$ as in Nazaret & Blei (2022), we obtain the final form of the ELBO (the full derivation is in Appendix C):
+$\sum_{\ell}^{L}\log\frac{p(\nu_{\ell};\mu_{\ell}^{\lambda},\sigma_{\ell}^{\lambda})}{q(\nu_{\ell};\nu_{\ell})}+\sum_{\ell}^{L}\sum_{n=1}^{D_{\ell}}\log\frac{p(\rho_{\ell n};\sigma_{\ell}^{\theta})}{q(\rho_{\ell n};\rho_{\ell n})}+\sum_{i=1}^{N}\log p(y_{i}|\bm{\lambda}\text{=}\bm{\nu},\bm{\theta}\text{=}\bm{\rho},x_{i}),$ (9)
 where distributions’ parameters are made explicit to distinguish them.
 The first two terms in the ELBO regularize the width of the layers and the magnitude of the parameters when priors are informative, whereas the third term accounts for the predictive performance.
-In practice, the finite variational parameters $\boldsymbol{\nu},\boldsymbol{\rho}$ are those used to parametrize the neural network rather than sampling $\boldsymbol{\lambda},\boldsymbol{\theta}$, which enables easy optimization via backpropagation.
+In practice, the finite variational parameters $\bm{\nu},\bm{\rho}$ are those used to parametrize the neural network rather than sampling $\bm{\lambda},\bm{\theta}$, which enables easy optimization via backpropagation.
 Maximizing Equation (9) will update each variational parameter $\nu_{\ell}$, which in turn will change the value of $D_{\ell}$ during training.
@@
 The latter can be initially set to same value across layers since they can freely adapt later, or it can be sampled from the prior $p(\lambda_{\ell})$.
-Therefore, we have two more hyper-parameters compared to the fixed-width network, but we make some considerations: i) it is always possible to use an uninformative prior over $\lambda_{\ell}$, removing the extra hyper-parameters letting the model freely adapt the width of each layer (as is typical of frequentist approaches); ii) the choice of higher level of hyper-parameters is known to be less stringent than that of hyper-parameters themselves [21, 5], so we do not need to explore many values of $\mu_{\ell}^{\lambda}$ and $\sigma_{\ell}^{\lambda}$; iii) our experiments suggest that AWNN can converge to similar widths regardless of the starting point $\nu_{\ell}$, so that we may just need to perform model selection over one/two sensible initial values; iv) the more data, the less the priors will matter.
+Therefore, we have two more hyper-parameters compared to the fixed-width network, but we make some considerations: i) it is always possible to use an uninformative prior over $\lambda_{\ell}$, removing the extra hyper-parameters letting the model freely adapt the width of each layer (as is typical of frequentist approaches); ii) the choice of higher level of hyper-parameters is known to be less stringent than that of hyper-parameters themselves (Goel & Degroot, 1981; Bernardo & Smith, 2009), so we do not need to explore many values of $\mu_{\ell}^{\lambda}$ and $\sigma_{\ell}^{\lambda}$; iii) our experiments suggest that AWN can converge to similar widths regardless of the starting point $\nu_{\ell}$, so that we may just need to perform model selection over one/two sensible initial values; iv) the more data, the less the priors will matter.
 3.1 Imposing a Soft Ordering on Neurons’ Importance
-Now that the learning objective has been formalized, the missing ingredient is the definition of the neural network $p(y_{i}|\boldsymbol{\lambda}\text{=}\boldsymbol{\nu},\boldsymbol{\theta}\text{=}\boldsymbol{\rho},x_{i})$ of Equation 4 as a modified MLP.
-Compared to a standard MLP, we use the variational parameters $\boldsymbol{\nu}$ that affect the truncation width at each hidden layer, whereas $\boldsymbol{\rho}$ are the weights.
+Now that the learning objective has been formalized, the missing ingredient is the definition of the neural network $p(y_{i}|\bm{\lambda}\text{=}\bm{\nu},\bm{\theta}\text{=}\bm{\rho},x_{i})$ of Equation 4 as a modified MLP.
+Compared to a standard MLP, we use the variational parameters $\bm{\nu}$ that affect the truncation width at each hidden layer, whereas $\bm{\rho}$ are the weights.
 We choose a monotonically decreasing function $f_{\ell}$, thus when a new neuron is added its relative importance is low and will not drastically impact the network output and hidden activations.
@@
 The first is to regularize the magnitude of the weights thanks to the prior $p(\theta_{\ell+1n})$, so that it may be difficult to compensate for the least important neurons that have a high $1/f_{\ell}(j)$.
-The second and less obvious strategy is to prevent the units’ activations of the current layer to compensate for high values by bounding their range, e.g., using a ReLU6 or tanh activation [51].
+The second and less obvious strategy is to prevent the units’ activations of the current layer to compensate for high values by bounding their range, e.g., using a ReLU6 or tanh activation (Sandler et al., 2018).
 We apply both strategies to our experiments, although we noted that they do not seem strictly necessary in practice.
-3.2 Rescaled Weight Initialization for Deep AWNN
-Rescaling the activations of hidden units using Equation 10 causes activations of deeper layers to quickly decay to zero for an AWNN MLP with ReLU nonlinearity initialized using the well known Kaiming scheme [25].
-This affects convergence since gradients get close to zero and it becomes slow to train deep AWNN MLPs.
+3.2 Rescaled Weight Initialization for Deep AWN
+Rescaling the activations of hidden units using Equation 10 causes activations of deeper layers to quickly decay to zero for an AWN MLP with ReLU nonlinearity initialized using the well known Kaiming scheme (He et al., 2015).
+This affects convergence since gradients get close to zero and it becomes slow to train deep AWN MLPs.
 We therefore derive a rescaled Kaiming weight that guarantees that the variance of activation across layers is constant at initialization.
@@
 Consider an MLP with activations as in Equation 10 and ReLU nonlinearity.
-At initialization, given $\alpha_{j}^{\ell}=\sigma\left(\sum_{k=1}^{D_{\ell-1}}w^{\ell}_{jk}h^{\ell-1}_{k}\right),\mathrm{Var}[w_{j*}^{\ell}]=\frac{2}{\sum_{j=1}^{D_{\ell-1}}f^{2}_{\ell}(j)}\Rightarrow\mathrm{Var}[\alpha_{j}^{\ell}]\approx\mathrm{Var}[\alpha_{j}^{\ell-1}]$.
+At initialization, given $\alpha_{j}^{\ell}=\sigma\left(\sum_{k=1}^{D_{\ell-1}}w^{\ell}_{jk}h^{\ell-1}_{k}\right),\mathrm{Var}[w_{jk}^{\ell}]=\frac{2}{\sum_{j=1}^{D_{\ell-1}}f^{2}_{\ell}(j)}\Rightarrow\mathrm{Var}[\alpha_{j}^{\ell}]\approx\mathrm{Var}[\alpha_{j}^{\ell-1}]$.
 Proof.
@@
 This change drastically impacts overall convergence on the SpiralHard synthetic dataset (described in Section 4), where it appears it would be otherwise hard to converge using a standard Kaiming initialization.
-Algorithm 1 AWNN Training Procedure
-1:  Input: Dataset $\mathcal{D}$, initialized AWNN model $\mathcal{M}$ (Section 3.2)
-2:  Output: Trained AWNN Model $\mathcal{M}$
+Algorithm 1 AWN Training Procedure
+1:Input: Dataset $\mathcal{D}$, initialized AWN model $\mathcal{M}$ (Section 3.2)
+2:Output: Trained AWN Model $\mathcal{M}$
 3:for each training epoch do
 4:  for batch in $\mathcal{D}$ do
-5:       $\mathcal{M}$)
+5:   update_width($\mathcal{M}$)
 6:   $\hat{y}\leftarrow\mathcal{M}(\text{batch})$
-7:       $\text{loss}\leftarrow\text{ELBO}(\mathcal{M},\text{batch},\hat{y})$ // Eq. 9
+[…]
 8:   $\mathcal{M}\leftarrow\text{backpropagation}(\mathcal{M},\text{loss})$
-9:  function $\mathcal{M}$):
-[…]
+9:function update_width($\mathcal{M}$):
+10:  for layer $\ell$ in $\mathcal{M}\text{.hidden\_layers}$ do
 11:   $D_{\ell}\leftarrow\text{quantile function of }f_{\ell}(\cdot;\nu_{\ell})\text{\ evaluated at }k$
-12:       Use $D_{\ell}$ to update $\boldsymbol{\rho}_{\ell},\boldsymbol{\rho}_{\ell+1}$ // add/remove neurons
+12:   Use $D_{\ell}$ to update $\bm{\rho}_{\ell},\bm{\rho}_{\ell+1}$ // add/remove neurons
 Algorithm 1 summarizes the main changes to the training procedure, namely the new initialization and the update of the model’s truncated width at each training step.
@@
 They are used at the very end of CNNs, in each Transformer layer, and they process messages coming from neighbors in DGNs.
-Our experiments focus on MLPs to showcase AWNN’s broad applicability, but there are many other scenarios where one can apply AWNN’s principles.
+Our experiments focus on MLPs to showcase AWN’s broad applicability, but there are many other scenarios where one can apply AWN’s principles.
 For instance, one could impose a soft ordering of importance on CNNs’ filters at each layer, therefore learning the number of filters during training.
-From a more theoretical perspective, we believe one could draw connections between our technique and the Information Bottleneck principle [56], which seeks maximally representative (i.e., performance) and compact representations (e.g., width).
-Finally, the analysis of different functions $f$ to soft-order neurons, for instance a power-law distribution or a sigmoid-like function, may have a different impact on convergence and performance and could be empirically investigated.
-AWNN needs to change the network at training time with an associated overhead.
-Our current implementation has only a 3-4$\times$ overhead on the smallest datasets, the worst case scenario for our approach, where the inference cost is negligible compared to the network update.
-For more realistic tasks, we believe our method is advantageous compared to the cost of a grid-search or alternating training/updating approaches.
-In addition, further optimizations are possible: i) by updating the width every $M$ mini-batches rather than at every step; ii) by allocating a larger network and dynamically grow it only if needed; iii) by creating ad-hoc support of AWNN on deep learning libraries.
+However, doing so would require a distinct formalization, which is why this lies beyond the scope of the present work.
+From a more theoretical perspective, we believe one could draw connections between our technique and the Information Bottleneck principle (Tishby et al., 2000), which seeks maximally representative (i.e., performance) and compact representations (e.g., width).
+In addition, we could try to revisit the theoretical results of Caron et al.
+(2025) in the NTK regime to better understand convergence properties of adaptive-width neural networks.
+since both are based on asymmetric rescaling of activations.
 4 Experiments and Setup
-The purpose of the empirical analysis is not to claim AWNN is generally better than the fixed-width baseline.
-Rather, we demonstrate how AWNN overcomes the problem of fixing the number of neurons by learning it end-to-end, thus reducing the amount of hyper-parameter configurations to test.
+The purpose of the empirical analysis is not to claim AWN is generally better than the fixed-width baseline.
+Rather, we demonstrate how AWN overcomes the problem of fixing the number of neurons by learning it end-to-end, thus reducing the amount of hyper-parameter configurations to test.
 As such, due to the nature of this work and similarly to Mitchell et al.
-[40], we use the remaining space to thoroughly study the behavior of AWNN, so that it becomes clear how to use it in practice.
-We first quantitatively verify that AWNN does not harm the performance compared to baseline models and compare the chosen width by means of grid-search model selection with the learned width of AWNN.
-Second, we check that AWNN chooses a larger width for harder tasks, which can be seen as increasing the hypotheses space until the neural network finds a good path to convergence.
-Third, we verify that convergence speed is not significantly altered by AWNN, so that the main limitation lies in the extra overhead for adapting the network at each training step.
-As a sanity check, we study conditions under which AWNN’s learned width does not seem to depend on starting hyper-parameters, so that their choice does not matter much.
-Finally, we analyze other practical advantages of training a neural network under the AWNN framework: the ability to compress information during training or post training, and the resulting trade-offs.
+(2023), we use the remaining space to thoroughly study the behavior of AWN, so that it becomes clear how to use it in practice.
+We first quantitatively verify that AWN does not harm the performance compared to baseline models and compare the chosen width by means of grid-search model selection with the learned width of AWN.
+Second, we check that AWN chooses a larger width for harder tasks, which can be seen as increasing the hypotheses space until the neural network finds a good path to convergence.
+Third, we verify that convergence speed is not significantly altered by AWN, so that the main limitation lies in the extra overhead for adapting the network at each training step.
+As a sanity check, we study conditions under which AWN’s learned width does not seem to depend on starting hyper-parameters, so that their choice does not matter much.
+In addition, we analyze other practical advantages of training a neural network under the AWN framework: the ability to compress information during training or post training, and the resulting trade-offs.
+Finally, we analyze the impact of different families of functions $f_{\ell}(j;\nu_{\ell})$ compared to the exponential mainly used in this work.
 Further analyses are in the Appendix.
-We compare baselines that undergo proper hyper-parameter tuning (called “Fixed”) against its AWNN version, where we replace any fixed MLP with an adaptive one.
+We compare baselines that undergo proper hyper-parameter tuning (called “Fixed”) against its AWN version, where we replace any fixed MLP with an adaptive one.
 First, we train an MLP on 3 synthetic tabular tasks of increasing binary classification difficulty, namely a double moon, a spiral, and a double spiral that we call SpiralHard.
 A stratified hold-out split of 70% training/10% validation/20% test for risk assessment is chosen at random for these datasets.
-Similarly, we consider a ResNet-20 [26] trained on 3 image classification tasks, namely MNIST [32], CIFAR10, and CIFAR100 [31], where data splits and preprocessing are taken from the original paper and AWNN is applied to the downstream classifier.
-In the sequential domain, we implemented a basic adaptive Recurrent Neural Network (RNN) and evaluated on the PMNIST dataset [67] with the same data split as MNIST.
-In the graph domain, we train a Graph Isomorphism Network [65] on the NCI1 and REDDIT-B classification tasks, where topological information matters, using the same split and evaluation setup of Errica et al.
-[16].
-Here, the first 1 hidden layer MLP as well as the one used in each graph convolutional layer are replaced by adaptive AWNN versions.
+Then, we consider 3 larger real-world tabular tasks with higher feature dimensionality, namely pol, MiniBooNE, and credit card clients (Beyazit et al., 2023).
+Similarly, we consider a ResNet-20 (He et al., 2016) trained on 3 image classification tasks, namely MNIST (LeCun, 1998), CIFAR10, and CIFAR100 (Krizhevsky, 2009), where data splits and preprocessing are taken from the original paper and AWN is applied to the downstream classifier.
+In the sequential domain, we implemented a basic adaptive Recurrent Neural Network (RNN) and evaluated on the PMNIST dataset (Zenke et al., 2017) with the same data split as MNIST.
+In the graph domain, we train a Graph Isomorphism Network (Xu et al., 2019) on the NCI1 and REDDIT-B classification tasks, where topological information matters, using the same split and evaluation setup of Errica et al.
+(2020).
+Here, the first 1 hidden layer MLP as well as the one used in each graph convolutional layer are replaced by adaptive AWN versions.
 On all these tasks, the metric of interest is the accuracy.
-Finally, for the textual domain we train a Transformer architecture [58] on the Multi30k English-German translation task [13], using a pretrained GPT-2 Tokenizer, and we evaluate the cross-entropy loss over the translated words.
+Finally, for the textual domain we train a Transformer architecture (Vaswani et al., 2017) on the Multi30k English-German translation task (Elliott et al., 2016), using a pretrained GPT-2 Tokenizer, and we evaluate the cross-entropy loss over the translated words.
 On tabular, image, and text-based tasks, an internal validation set (10%) for model selection is extracted from the union of outer training and validation sets, and the best configuration chosen according to the internal validation set is retrained 10 times on the outer train/validation/test splits, averaging test performances after an early stopping strategy on the validation set.
-Due to space reasons, we report datasets statistics and the hyper-parameter tried for the fixed and AWNN versions in Appendix F and G, respectively [^3] Code to reproduce results is in the supplementary material..
+Due to space reasons, we report datasets statistics and the hyper-parameter tried for the fixed and AWN versions in Appendix F and G, respectively [^2] Code to reproduce results is available at https://github.com/nec-research/Adaptive-Width-Neural-Networks..
 We ran the experiments on a server with 64 cores, 1.5TB of RAM, and 4 NVIDIA A40 with 48GB of memory.
 5 Results
-Table 1: Performances and total width of MLP layers for the fixed and AWNN versions of the various models used.
+Table 1: Performances and total width of MLP layers for the fixed and AWN versions of the various models used.
 The exact width chosen by model selection on the graph datasets is unknown since we report published results.
 “Linear" means the chosen downstream classifier is a linear model.
-Fixed AWNN Width (Fixed) Width (AWNN)
+Fixed AWN Width (Fixed) Width (AWN)
 Mean (Std) Mean (Std) Mean (Std)
@@
 SpiralHard 98.0 (2.0) 100.0 (0.0) 32 227.4 (32.4)
+pol 99.3 (0.2) 99.2 (0.1) 48 84 (11.0)
+MiniBooNE 92.9 (0.3) 93.2 (0.1) 32 53 (11.1)
+credit card 81.6 (0.1) 81.8 (0.1) 16 51 (12.0)
 PMNIST 91.1 (0.4) 95.7 (0.2) 24 806.3 (44.5)
@@
 We begin by discussing the quantitative results of our experiments: Table 1 reports means and standard deviations across the 10 final training runs.
-In terms of performance, we observe that AWNN is more stable or accurate than a fixed MLP on DoubleMoon, Spiral, SpiralHard, and PMNIST; all other things being equal, it seems that using more neurons and their soft ordering are the main contributing factors to these improvements.
-On the image datasets, performances of AWNN are comparable to those of the fixed baseline but for CIFAR100, due to an unlucky run that did not converge.
-There, AWNN learns a smaller total width compared to grid search.
-Results on graph datasets are interesting in two respects: First, the performance on REDDIT-B is significantly improved by AWNN both in terms of average performance and stability of results; second, and akin to PMNIST, the total learned width is significantly higher than those tried in Xu et al.
-[65], Errica et al.
-[16], meaning a biased choice of possible widths had a profound influence on risk estimation for DGN models (i.e., GIN).
+In terms of performance, we observe that AWN is more stable or accurate than a fixed MLP on tabular and sequential datasets; all other things being equal, it seems that using more neurons and their soft ordering are the main contributing factors to these improvements.
+On the image datasets, performances of AWN are comparable to those of the fixed baseline but for CIFAR100, due to an unlucky run that did not converge.
+There, AWN learns a smaller total width compared to grid search.
+Results on graph datasets are interesting in two respects: First, the performance on REDDIT-B is significantly improved by AWN both in terms of average performance and stability of results; second, and akin to PMNIST, the total learned width is significantly higher than those tried in Xu et al.
+(2019); Errica et al.
+(2020), meaning a biased choice of possible widths had a profound influence on risk estimation for DGN models (i.e., GIN).
 This result makes it evident that it is important to let the network decide how many neurons are necessary to solve the task.
 Appendix H shows what happens when we retrain Fixed baselines using the total width as the width of each layer.
-Finally, the results on the Multi30k show that the AWNN Transformer learns to use 200x parameters less than the fixed Transformer for the feed-forward networks, achieving a statistically comparable test loss.
-This result is appealing when read through the lenses of modern deep learning, as the power required by some neural networks such as Large Language Models [9] is so high that cannot be afforded by most institutions, and it demands future investigations.
+Finally, the results on the Multi30k show that the AWN Transformer learns to use 200x parameters less than the fixed Transformer for the feed-forward networks, achieving a statistically comparable test loss.
+This result is appealing when read through the lenses of modern deep learning, as the power required by some neural networks such as Large Language Models (Brown et al., 2020) is so high that cannot be afforded by most institutions, and it demands future investigations.
 Adaptation to Task Difficulty and Convergence
-Intuitively, one would expect that AWNN learned larger widths for more difficult tasks.
+Intuitively, one would expect that AWN learned larger widths for more difficult tasks.
 This is indeed what happens on the tabular datasets (and image datasets, see Appendix I) where some tasks are clearly harder than others.
@@
 It is also interesting to observe that the total width for a multi-layer MLP on SpiralHard is significantly lower than that achieved by a single-layer MLP, which is consistent with the circuit complexity theory arguments put forward in Bengio et al.
-[3], Mhaskar et al.
-[37].
-It also appears that convergence is not affected by the introduction of AWNN, as investigated in Figure 2 (right), which was not obvious considering the parametrization constraints encouraged by the rescaling of neurons’ activations.
+(2006); Mhaskar et al.
+(2017).
+It also appears that convergence is not affected by the introduction of AWN, as investigated in Figure 2 (right), which was not obvious considering the parametrization constraints encouraged by the rescaling of neurons’ activations.
 Figure 2: (Left) The learned width adapts to the increasing difficulty of the task, from the DoubleMoon to SpiralHard.
-(Right) AWNN reaches perfect test accuracy with a comparable amount of epochs on DoubleMoon and Spiral, while it converges faster on SpiralHard.
+(Right) AWN reaches perfect test accuracy with a comparable amount of epochs on DoubleMoon and Spiral, while it converges faster on SpiralHard.
 Training Stability Analysis
-To support our argument that AWNN can reduce the time spent performing hyper-parameter selection, we check whether AWNN learns a consistent amount of neurons across different training runs and hyper-parameter choices.
+To support our argument that AWN can reduce the time spent performing hyper-parameter selection, we check whether AWN learns a consistent amount of neurons across different training runs and hyper-parameter choices.
 Figure 3 reports the impact of the batch size and starting width averaged across the different configurations tried during model selection.
 Smaller batch sizes cause more instability, but in the long run we observe convergence to a similar width.
-Convergence with respect to different rates holds, instead, for the bounded ReLU6 activation; Appendix J shows that unbounded activations may cause the network to converge more slowly to the same width, which is in accord with the considerations about counterbalancing the rescaling effect of Section 3.1.
+Convergence with respect to different rates holds, instead, for the bounded ReLU6 activation; Appendix K shows that unbounded activations may cause the network to converge more slowly to the same width, which is in accord with the considerations about counterbalancing the rescaling effect of Section 3.1.
 Therefore, whenever possible, we recommend using bounded activations.
@@
 We demonstrate the effect of an informative prior by performing a width-annealing experiment on the SpiralHard dataset.
-We set an uninformative $p(\boldsymbol{\theta})$ and ReLU6 nonlinearity.
+We set an uninformative $p(\bm{\theta})$ and ReLU6 nonlinearity.
 At epoch 1000, we introduce $p(\lambda_{\ell})=\mathcal{N}(\lambda_{\ell};0.05,1)$, and gradually anneal the standard deviation up to $0.1$ at epoch $2500$.
@@
 This technique might be particularly useful to compress large models with billions of parameters.
-Figure 4: It is possible to regularize the width at training time by increasing the magnitude of the loss term $\log\frac{p(\boldsymbol{\nu})}{q(\boldsymbol{\nu})}$.
+Figure 4: It is possible to regularize the width at training time by increasing the magnitude of the loss term $\log\frac{p(\bm{\nu})}{q(\bm{\nu})}$.
 The total width is reduced by more than 50% (left) while preserving accuracy (right).
-The inset plot refers to the loss term that AWNN tries to maximize.
+The inset plot refers to the loss term that AWN tries to maximize.
 Post-hoc Truncation Achieves a Trade-off between Performance and Compute Resources
@@
 (Right) The distribution of neurons’ activations for all Spiral test samples follows an exponential-like curve.
+Impact of Different Functions $f_{\ell}(j;\nu_{\ell})$
+We conclude our main analyses with an ablation study on the use of different decreasing functions $f_{\ell}(j;\nu_{\ell})$, considering the SpiralHard dataset, to understand how different functions have different properties.
+We tested: i) a Power Law distribution $f_{\ell}(j;\nu_{\ell},\alpha)=Z(j+j_{sat})^{-\nu_{\ell}}$, with $Z$ being a normalizing term, initial $\nu_{\ell}\in[2.,3.]$ and low degree saturation $j_{sat}$ chosen between $[0.,2.]$; ii) a sigmoid-like function $f_{\ell}(j;\nu_{\ell},b)=1-\sigma(j-\nu_{\ell})$, where $\sigma$ is the sigmoidal activation, $\nu_{\ell}\in[128,256]$ is the transition point.
+The results are reported in Table 2.
+When averaged over all possible hyper-parameter configurations, there is no significant deviation in performance between distribution, and there is always at least one configuration attaining the best performance.
+However, the Power Law distribution introduces more neurons than the exponential due to its long-tail.
+Instead, the Sigmoidal function tends to allocates less neurons, but the drawback is the loss of the ability to truncate the network after training, since the importance of neurons is very close to 1 before the transition point and the transition itself, for the chosen fixed parameters, is relatively sharp.
+The choice of the function family $f_{\ell}(j;\nu_{\ell})$ remains an interesting direction of further exploration, given that each family has different properties that might be more or less suitable depending on the use cases.
+Table 2: Analysis of the impact of different families of importance functions on SpiralHard validation performance and learned width, averaged across different hyper-parameter configurations.
+$f_{\ell}(j;\nu_{\ell})$ Mean Accuracy Max Accuracy Mean Total Width
+Exponential 80.27 (19.9) 100.00 954.4 (1083.9)
+Power Law 81.82 (16.6) 100.00 2952.4 (3371.6)
+Sigmoidal 76.85 (18.3) 100.00 426.8 (268.4)
 6 Conclusions
@@
 A by-product of neurons’ ordering is the ability to easily compress the network during or after training, which is relevant in the context of foundational models trained on large data, which are believed to require billions of parameters.
-Finally, we have tested AWNN on different models and data domains to prove its broad scope of applicability: a Transformer architecture achieved a similar loss with 200x less parameters.
-Appendix A Background Notions
-Since we were mostly inspired by the work of Nazaret and Blei [42], we complement the main text with an introduction to truncated distributions.
+Finally, we have tested AWN on different models and data domains to prove its broad scope of applicability: a Transformer architecture achieved a similar loss with 200x less parameters.
+7 Reproducibility Statement
+The code provided in the supplementary material relies on libraries for automatic experimentation, which enforce reproducibility of the results.
+These libraries require that the user specifies every detail, from data splitting strategies to hyper-parameter selection and final evaluations, and subsequently the experiment is automatized and results are provided.
+Appendix A Truncated Distributions’ Notions
+Since we were mostly inspired by the work of Nazaret & Blei (2022), we complement the main text with an introduction to truncated distributions.
 In particular, by truncating a distribution at its quantile function evaluated at $k$ (see Appendix B for a visual explanation), the support of the distribution becomes finite and countable, hence we can compute expectations in a finite number of steps.
-A truncated distribution should ideally satisfy some requirements defined in Nazaret and Blei [42] and illustrated below.
-Definition A.1.
-A variational family $Q={q(\boldsymbol{\omega})}$ over $\mathbb{N}^{+}$ is unbounded with connected and bounded members if
+A truncated distribution should ideally satisfy some requirements defined in Nazaret & Blei (2022) and illustrated below.
+Definition A.1 (Nazaret & Blei (2022)).
+A variational family $Q={q(x;\bm{\omega})}$ over $\mathbb{N}^{+}$ is unbounded with connected and bounded members if
   1. 1.
@@
   3. 3.
-    Each parameter in the set $\boldsymbol{\omega}$ is a continuous variable.
+    Each parameter in the set $\bm{\omega}$ is a continuous variable.
 The first condition allows us to compute the expectation over any $q\in Q$ in finite time, condition 2 ensures there is a parametrization that assigns enough probability mass to each point in the support of $q$, and condition 3 is necessary for learning via backpropagation.
-An important distinction with Nazaret and Blei [42] is that we do not need nor want condition 2 to be satisfied.
+An important distinction with Nazaret & Blei (2022) is that we do not need nor want condition 2 to be satisfied.
 As a matter of fact, we want to enforce a strict ordering of neurons where the first is always the most important one.
-Condition 2 is useful, as done in Nazaret and Blei [42], to assign enough ‘importance” to a specific layer of an adaptive-depth architecture, while previous layers are still functional to the final result.
+Condition 2 is useful, as done in Nazaret & Blei (2022), to assign enough ‘importance” to a specific layer of an adaptive-depth architecture, while previous layers are still functional to the final result.
 In the context of adaptive-width networks, however, this mechanism may leave a lot of neurons completely unutilized, letting the network grow indefinitely.
-That is why we require that each distribution $q$ is monotonically decreasing and all possible widths should have non-zero importance.
+That is why we require that each distribution $f_{\ell}$ is monotonically decreasing and all possible widths should have non-zero importance.
 We change the definition for the distributions we are interested in as follows:
 Definition A.2.
-A variational family $Q={q(\boldsymbol{\omega})}$ over $\mathbb{N}^{+}$ is unbounded with connected, bounded, and decreasing members if
+A family of distributions $Q={f(x;\bm{\lambda})}$ over $\mathbb{N}^{+}$ is unbounded with connected, bounded, and decreasing members if
   1. 1.
-    $\forall q\in Q$, $\operatorname{support}(q)$ is bounded
+    $\forall f\in Q$, $\operatorname{support}(f)$ is bounded
   2. 2.
-    $\forall q\in Q$, $q$ is monotonically decreasing and $\forall x\in\mathbb{N}^{+},q(x)>0$
+    $\forall f\in Q$, $f$ is monotonically decreasing and $\forall x\in\mathbb{N}^{+},f(x)>0$
   3. 3.
-    Each parameter in the set $\boldsymbol{\omega}$ is a continuous variable.
+    Each parameter in the set $\bm{\lambda}$ is a continuous variable.
 Appendix B How to Compute $D_{\ell}$ in Practice
 To aid the understanding of how one computes $D_{\ell}$, we provide a visualization of the exponential distribution, its discretized version (Equation 6) and the quantile function evaluated at $k=0.9$ in Figure 6 below.
+Note that typically, newly introduced neurons are multiplied by a very small importance, and their random weights are sampled from a standard Gaussian.
+This makes their initial contribution very small, so it does not drastically alter the loss/gradients/convergence.
+From a backpropagation viewpoint, the reason why new neurons are introduced is that previous neurons need to become relatively more important.
+So, when the chosen importance distributions grow too slowly, as in the case of a Power law family of distributions, this can slow down convergence because it takes a long time to make the last neurons more important, and in addition a large tail or unimportant neurons is added if the chosen quantile is too high.
+The exponential distribution, instead, worked pretty well without having to tune it extensively.
 Figure 6: On the left, we provide a visualization of the exponential density with $\lambda=0.5$, its cumulative density, and the quantile function evaluated at $k=0.9$, which returns a value close to $5$.
@@
 The full ELBO derivation is as follows:
-$\mathbb{E}_{q(\boldsymbol{\lambda},\boldsymbol{\theta})}\left[\log\frac{p(Y,\boldsymbol{\lambda},\boldsymbol{\theta}|X)}{q(\boldsymbol{\lambda},\boldsymbol{\theta})}\right]=$ (11)
-$\mathbb{E}_{q(\boldsymbol{\lambda})}\mathbb{E}_{q(\boldsymbol{\theta}|\boldsymbol{\lambda})}[\log p(Y,\boldsymbol{\lambda},\boldsymbol{\theta}|X)-\log q(\boldsymbol{\lambda},\boldsymbol{\theta})]=$ (12)
-$\mathbb{E}_{q(\boldsymbol{\lambda})}\mathbb{E}_{q(\boldsymbol{\boldsymbol{\theta}|\lambda})}[\log(p(Y|\boldsymbol{\lambda},\boldsymbol{\theta},X)p(\boldsymbol{\theta})p(\boldsymbol{\lambda}))-\log(q(\boldsymbol{\lambda})q(\boldsymbol{\theta}|\boldsymbol{\lambda}))]$ (13)
+$\mathbb{E}_{q(\bm{\lambda},\bm{\theta})}\left[\log\frac{p(Y,\bm{\lambda},\bm{\theta}|X)}{q(\bm{\lambda},\bm{\theta})}\right]=$ (11)
+$\mathbb{E}_{q(\bm{\lambda})}\mathbb{E}_{q(\bm{\theta}|\bm{\lambda})}[\log p(Y,\bm{\lambda},\bm{\theta}|X)-\log q(\bm{\lambda},\bm{\theta})]=$ (12)
+$\mathbb{E}_{q(\bm{\lambda})}\mathbb{E}_{q(\bm{\theta|\lambda})}[\log(p(Y|\bm{\lambda},\bm{\theta},X)p(\bm{\theta})p(\bm{\lambda}))-\log(q(\bm{\lambda})q(\bm{\theta}|\bm{\lambda}))]$ (13)
 where we factorized probabilities according to the independence assumptions of the graphical model. Then
-$\mathbb{E}_{q(\boldsymbol{\lambda})}\mathbb{E}_{q(\boldsymbol{\boldsymbol{\theta}|\lambda})}[\log p(Y|\boldsymbol{\lambda},\boldsymbol{\theta},X)+\log p(\boldsymbol{\lambda})+\log p(\boldsymbol{\theta})-\log q(\boldsymbol{\lambda})-\log q(\boldsymbol{\theta}|\boldsymbol{\lambda})]=$ (14)
-$\mathbb{E}_{q(\boldsymbol{\lambda})}\mathbb{E}_{q(\boldsymbol{\boldsymbol{\theta}|\lambda})}[\log\frac{p(\boldsymbol{\lambda})}{q(\boldsymbol{\lambda})}+\log\frac{p(\boldsymbol{\theta})}{q(\boldsymbol{\theta}|\boldsymbol{\lambda})}+\log p(Y|\boldsymbol{\lambda},\boldsymbol{\theta},X)]\approx$ (15)
-$\log\frac{p(\boldsymbol{\nu})}{q(\boldsymbol{\nu})}+\log\frac{p(\boldsymbol{\rho})}{q(\boldsymbol{\rho}|\boldsymbol{\nu})}+\log p(Y|\boldsymbol{\nu},\boldsymbol{\rho},X)]$ (16)
+$\mathbb{E}_{q(\bm{\lambda})}\mathbb{E}_{q(\bm{\theta|\lambda})}[\log p(Y|\bm{\lambda},\bm{\theta},X)+\log p(\bm{\lambda})+\log p(\bm{\theta})-\log q(\bm{\lambda})-\log q(\bm{\theta}|\bm{\lambda})]=$ (14)
+$\mathbb{E}_{q(\bm{\lambda})}\mathbb{E}_{q(\bm{\theta|\lambda})}[\log\frac{p(\bm{\lambda})}{q(\bm{\lambda})}+\log\frac{p(\bm{\theta})}{q(\bm{\theta}|\bm{\lambda})}+\log p(Y|\bm{\lambda},\bm{\theta},X)]\approx$ (15)
+$\log\frac{p(\bm{\nu})}{q(\bm{\nu})}+\log\frac{p(\bm{\rho})}{q(\bm{\rho}|\bm{\nu})}+\log p(Y|\bm{\nu},\bm{\rho},X)]$ (16)
 where we use the first-order approximation in the last step.
 Note that when computing $\frac{p(\theta)}{q(\theta|\lambda)}$, the products from $D_{\ell+1}$ to infinity cancel out.
-Equation 13 follows by simply applying definitions, the iid assumption $p(Y|\boldsymbol{\nu},\boldsymbol{\rho},X)=\prod_{i}^{N}p(y_{i}|\boldsymbol{\nu},\boldsymbol{\rho},x_{i})$, and by transforming products into summations using the logarithm.
-An important note about the first-order (Taylor) approximation $E_{q(\boldsymbol{\lambda};\boldsymbol{\nu})q(\boldsymbol{\theta}|\boldsymbol{\lambda};\boldsymbol{\rho})}[f(\boldsymbol{\lambda},\boldsymbol{\theta})]\approx E_{q(\boldsymbol{\lambda};\boldsymbol{\nu})}[f(\boldsymbol{\lambda},\boldsymbol{\rho})]\approx f(\boldsymbol{\nu},\boldsymbol{\rho})$ is that the function $f(\boldsymbol{\lambda},\boldsymbol{\theta})=\log\frac{p(\boldsymbol{\lambda})}{q(\boldsymbol{\lambda})}+\log\frac{p(\boldsymbol{\theta})}{q(\boldsymbol{\theta}|\boldsymbol{\lambda})}+\log p(Y|\boldsymbol{\lambda},\boldsymbol{\theta},X)$ is not differentiable with respect to $D_{\ell}$ when there is a jump in the quantile evaluated at $k$, so it would seem we cannot apply it.
-This also happens in Nazaret and Blei [42], Equation 11, when the function is a neural network containing non-differentiable activations like ReLU, but we could not find a discussion about it.
+Equation 13 follows by simply applying definitions, the iid assumption $p(Y|\bm{\nu},\bm{\rho},X)=\prod_{i}^{N}p(y_{i}|\bm{\nu},\bm{\rho},x_{i})$, and by transforming products into summations using the logarithm.
+An important note about the first-order (Taylor) approximation $E_{q(\bm{\lambda};\bm{\nu})q(\bm{\theta}|\bm{\lambda};\bm{\rho})}[f(\bm{\lambda},\bm{\theta})]\approx E_{q(\bm{\lambda};\bm{\nu})}[f(\bm{\lambda},\bm{\rho})]\approx f(\bm{\nu},\bm{\rho})$ is that the function $f(\bm{\lambda},\bm{\theta})=\log\frac{p(\bm{\lambda})}{q(\bm{\lambda})}+\log\frac{p(\bm{\theta})}{q(\bm{\theta}|\bm{\lambda})}+\log p(Y|\bm{\lambda},\bm{\theta},X)$ is not differentiable with respect to $D_{\ell}$ when there is a jump in the quantile evaluated at $k$, so it would seem we cannot apply it.
+This also happens in Nazaret & Blei (2022), Equation 11, when the function is a neural network containing non-differentiable activations like ReLU, but we could not find a discussion about it.
 Importantly, the function is not differentiable at a countable set of points (one for each possible width), which has a measure null.
@@
 Here, we want to show that the change in the ELBO after a change in the number of neurons is bounded.
-Therefore, we provide a theoretical result (Proposition D.1) that shows that the ELBO satisfies the Lipschitz continuity.
-Proposition D.1 (AWNN ELBO Lipschitz continuity).
-The ELBO loss of Equation 9, with respect to the change in the depth $D_{\ell}$ for the layer $\ell$, is Lipschitz continuous.
+Therefore, we provide a theoretical result (proposition˜D.1) that shows that the ELBO satisfies the Lipschitz continuity.
+Proposition D.1 (AWN ELBO Lipschitz continuity).
+The ELBO loss of eq.˜9, with respect to the change in the depth $D_{\ell}$ for the layer $\ell$, is Lipschitz continuous.
 Proof.
-We focus on the term involving $D_{\ell}$ of the ELBO, we write Equation 9 as
-$\log\frac{p(\boldsymbol{\nu})}{q(\boldsymbol{\nu})}+\log\frac{p(\boldsymbol{\rho})}{q(\boldsymbol{\rho}|\boldsymbol{\nu})}+\log p(Y|\boldsymbol{\nu},\boldsymbol{\rho},X)$
-where only the second and last terms depend on $\mathbf{D}=\{D_{\ell}\}_{\ell=1}^{L}$ Let’s first define
-$\log\frac{p(\boldsymbol{\rho})}{q(\boldsymbol{\rho}|\boldsymbol{\nu})}=\sum_{{\ell}=1}^{L}\sum_{n=1}^{D_{\ell}}\log\frac{p(\rho^{\ell}_{n})}{q(\rho^{\ell}_{n}|\boldsymbol{\nu})}=\sum_{{\ell}=1}^{L}f_{1}(D_{\ell})$
+We focus on the term involving $D_{\ell}$ of the ELBO, we write eq.˜9 as
+$\log\frac{p(\bm{\nu})}{q(\bm{\nu})}+\log\frac{p(\bm{\rho})}{q(\bm{\rho}|\bm{\nu})}+\log p(Y|\bm{\nu},\bm{\rho},X)$
+where only the second and last terms depend on $\mathbf{D}=\{D_{\ell}\}_{\ell=1}^{L}$ . Let’s first define
+$\log\frac{p(\bm{\rho})}{q(\bm{\rho}|\bm{\nu})}=\sum_{{\ell}=1}^{L}\sum_{n=1}^{D_{\ell}}\log\frac{p(\rho^{\ell}_{n})}{q(\rho^{\ell}_{n}|\bm{\nu})}=\sum_{{\ell}=1}^{L}f_{1}(D_{\ell})$
 We have that $f_{1}(D_{\ell})$ is Lipschitz continuous, indeed, when $D_{\ell}$ changes to $D_{l^{\prime}}$, we have
-$|f_{1}(D_{\ell^{\prime}})-f_{1}(D_{\ell})|$ $=|\sum_{n=D_{\ell}}^{D_{\ell^{\prime}}}\log\frac{p(\boldsymbol{\rho}_{n})}{q(\boldsymbol{\rho}_{n}|\boldsymbol{\nu})}|$ (17)
-$\leq~\sum_{n=D_{\ell}}^{D_{\ell^{\prime}}}|\log\frac{p(\boldsymbol{\rho}_{n})}{q(\boldsymbol{\rho}_{n}|\boldsymbol{\nu})}|$ (18)
-$\leq\max_{n}|\log\frac{p(\boldsymbol{\rho}_{n})}{q(\boldsymbol{\rho}_{n}|\boldsymbol{\nu})}||D_{\ell^{\prime}}-D_{\ell}|$ (19)
+$|f_{1}(D_{\ell^{\prime}})-f_{1}(D_{\ell})|$ $=|\sum_{n=D_{\ell}}^{D_{\ell^{\prime}}}\log\frac{p(\bm{\rho}_{n})}{q(\bm{\rho}_{n}|\bm{\nu})}|$ (17)
+$\leq \sum_{n=D_{\ell}}^{D_{\ell^{\prime}}}|\log\frac{p(\bm{\rho}_{n})}{q(\bm{\rho}_{n}|\bm{\nu})}|$ (18)
+$\leq\max_{n}|\log\frac{p(\bm{\rho}_{n})}{q(\bm{\rho}_{n}|\bm{\nu})}||D_{\ell^{\prime}}-D_{\ell}|$ (19)
 Therefore
 $|f_{1}(D_{\ell^{\prime}})-f_{1}(D_{\ell})|\leq M|D_{\ell^{\prime}}-D_{\ell}|$
-with $M=\max_{n}|\log\frac{p(\boldsymbol{\rho}_{n})}{q(\boldsymbol{\rho}_{n}|\boldsymbol{\nu})}|$. We now look at the last term,
-$f_{2}(\mathbf{D})=\log p(Y|\boldsymbol{\nu},\boldsymbol{\rho},X)$
+with $M=\max_{n}|\log\frac{p(\bm{\rho}_{n})}{q(\bm{\rho}_{n}|\bm{\nu})}|$. We now look at the last term,
+$f_{2}(\mathbf{D})=\log p(Y|\bm{\nu},\bm{\rho},X)$
 This function is a neural network followed by the squared norm.
-Therefore, $f_{2}(\mathbf{D})$ is continuous almost everywhere [59], therefore Lipschitz on both parameters and the input.
+Therefore, $f_{2}(\mathbf{D})$ is continuous almost everywhere (Virmaux & Scaman, 2018), therefore Lipschitz on both parameters and the input.
 When we change $D_{\ell}$ to $D_{\ell^{\prime}}$, we are adding network parameters, in particular, if we look at the two-layer network, we have
@@
 $f^{\prime}(t)=\left(f_{1}^{\prime}(t),\dots,f_{D}^{\prime}(t)\right)\in\mathbb{R}^{D\times 1}$
-For reasons that will become clear later, we may want to compute the gradient of the loss function with respect to the intermediate activations $\boldsymbol{\alpha}^{\ell}$ at a given layer, that is $\nabla\mathcal{L}(\boldsymbol{\alpha}^{\ell})=\left(\frac{\partial\mathcal{L}(\alpha_{1}^{\ell})}{\partial\alpha_{1}^{\ell}},\dots,\frac{\partial\mathcal{L}(\alpha_{N^{\ell}}^{\ell})}{\partial\alpha_{N^{\ell}}^{\ell}}\right)$.
+For reasons that will become clear later, we may want to compute the gradient of the loss function with respect to the intermediate activations $\bm{\alpha}^{\ell}$ at a given layer, that is $\nabla\mathcal{L}(\bm{\alpha}^{\ell})=\left(\frac{\partial\mathcal{L}(\alpha_{1}^{\ell})}{\partial\alpha_{1}^{\ell}},\dots,\frac{\partial\mathcal{L}(\alpha_{N^{\ell}}^{\ell})}{\partial\alpha_{N^{\ell}}^{\ell}}\right)$.
 We focus on the $i$-th partial derivative $\frac{\partial\mathcal{L}(\alpha_{i}^{\ell})}{\partial\alpha_{i}^{\ell}}$, where the only variable is $\alpha_{i}^{\ell}$.
-Then, we view the computation of the loss function starting from $\alpha_{i}^{\ell}$ as a composition of a function $\boldsymbol{\alpha}^{\ell+1}:\mathbb{R}\rightarrow\mathbb{R}^{N^{\ell+1}}=\left(\alpha_{1}^{\ell+1}(\alpha_{i}^{\ell})\dots,\alpha_{{N^{\ell+1}}}^{\ell+1}(\alpha_{i}^{\ell})\right)$ and another function (abusing the notation) $\mathcal{L}:\mathbb{R}^{N^{\ell+1}}\rightarrow\mathbb{R}$ that computes the loss value starting from $\boldsymbol{\alpha}^{\ell+1}$.
+Then, we view the computation of the loss function starting from $\alpha_{i}^{\ell}$ as a composition of a function $\bm{\alpha}^{\ell+1}:\mathbb{R}\rightarrow\mathbb{R}^{N^{\ell+1}}=\left(\alpha_{1}^{\ell+1}(\alpha_{i}^{\ell})\dots,\alpha_{{N^{\ell+1}}}^{\ell+1}(\alpha_{i}^{\ell})\right)$ and another function (abusing the notation) $\mathcal{L}:\mathbb{R}^{N^{\ell+1}}\rightarrow\mathbb{R}$ that computes the loss value starting from $\bm{\alpha}^{\ell+1}$.
 By the chain rule:
@@
 Proof.
-Let us start from the first case of $\sigma^{\prime}(0)\approx 1$. Using the Taylor expansions for the moments of functions of random variables as in Glorot and Bengio [20]
+Let us start from the first case of $\sigma^{\prime}(0)\approx 1$. Using the Taylor expansions for the moments of functions of random variables as in Glorot & Bengio (2010)
 $\mathrm{Var}[\alpha^{\ell}_{i}]=\mathrm{Var}\left[\sigma\left(\sum_{j=1}^{D_{\ell-1}}w^{\ell}_{ij}\alpha^{\ell-1}_{j}p^{\ell-1}_{j}\right)\right]$ (25)
@@
 We can also abstract from the indexes, since the weight variables are i.i.d.
-and from that it follows that $\mathrm{Var}[\alpha^{\ell}_{i}]=\mathrm{Var}[\alpha^{\ell}_{j}]\ \forall i,j,\ \ i,j\in\{1,\dots,D_{\ell}\}$, obtaining [^4] It is worth noting that, in standard MLPs, $p_{j}^{\ell}=1$ so we recover the derivation of Glorot and Bengio [20], since $\sum_{j=1}^{D_{\ell-1}}\left(p^{\ell-1}_{j}\right)^{2}$ would be equal to $D_{\ell-1}$.
-In Glorot and Bengio [20] our $D_{\ell-1}$ is denoted as “$n_{i^{\prime}}$” (see Equation 5).
+and from that it follows that $\mathrm{Var}[\alpha^{\ell}_{i}]=\mathrm{Var}[\alpha^{\ell}_{j}]\ \forall i,j,\ \ i,j\in\{1,\dots,D_{\ell}\}$, obtaining [^3] It is worth noting that, in standard MLPs, $p_{j}^{\ell}=1$ so we recover the derivation of Glorot & Bengio (2010), since $\sum_{j=1}^{D_{\ell-1}}\left(p^{\ell-1}_{j}\right)^{2}$ would be equal to $D_{\ell-1}$.
+In Glorot & Bengio (2010) our $D_{\ell-1}$ is denoted as $n_{i^{\prime}}$” (see Equation 5).
 $\mathrm{Var}[\alpha^{\ell}]\approx\mathrm{Var}[w^{\ell}]\mathrm{Var}[\alpha^{\ell-1}]\sum_{j=1}^{D_{\ell-1}}\left(p^{\ell-1}_{j}\right)^{2},$ (31)
@@
 Condition on the gradients
-From a backpropagation perspective, a similar desideratum would be to ensure that $\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{i}^{\ell})}{\partial\alpha_{i}^{\ell}}\right]=\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{i}^{\ell+1})}{\partial\alpha_{i}^{\ell+1}}\right]$ [^5] Note that what we impose is different from Glorot and Bengio [20], where the specific position $i$ is irrelevant.
+From a backpropagation perspective, a similar desideratum would be to ensure that $\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{i}^{\ell})}{\partial\alpha_{i}^{\ell}}\right]=\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{i}^{\ell+1})}{\partial\alpha_{i}^{\ell+1}}\right]$ [^4] Note that what we impose is different from Glorot & Bengio (2010), where the specific position $i$ is irrelevant.
 Here, we are asking that the variance of the gradients for neurons in the same position $i$, but at different layers, stays constant.
 Alternatively, we could impose an equivalence for all $i\neq i^{\prime}$..
-Using Equation 22, and considering as in Glorot and Bengio [20] that at initialization we are in a linear regime where $\sigma^{\prime}(x)\approx 1$,
+Using Equation 22, and considering as in Glorot & Bengio (2010) that at initialization we are in a linear regime where $\sigma^{\prime}(x)\approx 1$,
 $\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{i}^{\ell})}{\partial\alpha_{i}^{\ell}}\right]=\mathrm{Var}\left[\sum_{j=1}^{D_{\ell+1}}\frac{\partial\mathcal{L}(\alpha_{j}^{\ell+1})}{\partial\alpha_{j}^{\ell+1}}\frac{\partial\alpha_{j}^{\ell+1}(\alpha_{i}^{\ell})}{\partial\alpha_{i}^{\ell}}\right]=\mathrm{Var}\left[\sum_{j=1}^{D_{\ell+1}}\frac{\partial\mathcal{L}(\alpha_{j}^{\ell+1})}{\partial\alpha_{j}^{\ell+1}}\frac{\partial\alpha_{j}^{\ell+1}(\alpha_{i}^{\ell})}{\partial\alpha_{i}^{\ell}}\right]$ (33)
@@
 Because $p(x)$ is symmetric, $\mathbb{E}[x]=0$. Then
-[…]
+$\frac{1}{2}\int_{-\infty}^{{+\infty}}x^{2}p(x)dx=\frac{1}{2}\int_{-\infty}^{{+\infty}}(x-\mathbb{E}[x])^{2}p(x)dx=\frac{1}{2}\mathrm{Var}[x].$ (44)
 Using Lemma E.1, we can rewrite Equation 26 as
@@
 From a backpropagation perspective, since $\sigma^{\prime}(x)\not\approx 1$, we use similar arguments as in He et al.
-[25]: we assume that the pre-activations $\alpha$ have zero mean, then the derivative of the ReLU can have values zero or one with equal probability.
+(2015): we assume that the pre-activations $\alpha$ have zero mean, then the derivative of the ReLU can have values zero or one with equal probability.
 One can show, by expanding the definition of variance, that $\mathrm{Var}[\sigma^{\prime}\left(\sum_{k=1}^{D_{\ell}}w^{\ell+1}_{jk}\alpha^{\ell}_{k}p^{\ell}_{k}\right)w^{\ell+1}_{ji}]=\frac{1}{2}\mathrm{Var}[w^{\ell+1}_{ji}]$.
@@
 Comparison of convergence for different initializations
-Below, we provide a comparison of convergence between an AWNN ReLU MLP initialized with the standard Kaiming scheme and one initialized according to our theoretical results (“Kaiming+").
+Below, we provide a comparison of convergence between an AWN ReLU MLP initialized with the standard Kaiming scheme and one initialized according to our theoretical results (“Kaiming+").
 Figure 7: (Left) Effect of the rescaled initialization scheme (“Kaiming+”) on a ReLU-based MLP, where neurons’ activations are computed using Equation 10.
@@
 The synthetic datasets DoubleMoon, Spiral, and SpiralHard are shown in Figure 8.
-These binary classification datasets have been specifically created to analyze the behavior of AWNN in a controlled scenario where we are sure that the difficulty of the task increases.
+These binary classification datasets have been specifically created to analyze the behavior of AWN in a controlled scenario where we are sure that the difficulty of the task increases.
 We provide the code to generate these datasets in the supplementary material.
-Figure 8: The DoubleMoon, Spiral, and SpiralHard synthetic datasets are used to test AWNN’s inner workings, and are ordered difficulty. Orange and blue colors denote different classes.
-Table 2 reports information on their base statistics and the data splits created to carry out model selection (inner split) and risk assessment (outer split), respectively.
+Figure 8: The DoubleMoon, Spiral, and SpiralHard synthetic datasets are used to test AWN’s inner workings, and are ordered difficulty. Orange and blue colors denote different classes.
+Table 3 reports information on all datasets’ statistics and the data splits created to carry out model selection (inner split) and risk assessment (outer split), respectively.
 Table 3: Dataset statistics and number of samples in each split are shown.
 # Samples # Features # Classes Outer Split (TR/VL/TE) Inner Split (TR/VL)
-DoubleMoon 2500 2 2 1750/250/500 1800/200
-Spiral 2500 2 2 1750/250/500 1800/200
-SpiralHard 5000 2 2 3500/500/1000 3600/400
-MNIST 70000 28x28 10 53000/7000/10000 54000/6000
-CIFAR10 60000 32x32x3 10 44000/6000/10000 45000/5000
-CIFAR100 60000 32x32x3 100 44000/6000/10000 45000/5000
-NCI1 4110 37 2 3288/411/411 3329/370
-REDDIT-B 2000 1 2 1800/100/100 1710/190
-Multi30k 31000 128x50257 Multilabel 29000/1000/1000 29000/1000
+DoubleMoon 5000 2 2 3600/400/1000 3600/400
+Spiral 5000 2 2 3600/400/1000 3600/400
+SpiralHard 10000 2 2 7200/800/2000 7200/800
+pol 15000 48 2 10800/1200/3000 10800/1200
+MiniBooNE 130064 50 2 93645/10406/26013 93645/10406
+credit card clients 30000 23 2 21600/2400/6000 21600/2400
+MNIST 70000 28x28 10 48610/9723/11667 48610/9723
+CIFAR10 60000 32x32x3 10 41665/8334/10001 41665/8334
+CIFAR100 60000 32x32x3 100 41665/8334/10001 41665/8334
+NCI1 4110 37 2 3328/370/412 3328/370
+REDDIT-B 2000 1 2 1620/180/200 1620/180
+Multi30k 31000 128x50257 Multilabel 29046/968/1000 29046/968
 Note that for Multi30k, the maximum sequence length is 128 and the number of possible tokens is 50527.
@@
 For a detailed description of the graph datasets, the reader can refer to Errica et al.
-[16].
-Appendix G Hyper-parameters for fixed and AWNN models
-For each data domain, we discuss the set of hyper-parameters tried during model selection for the fixed baseline and for AWNN.
+(2020).
+Appendix G Hyper-parameters for fixed and AWN models
+For each data domain, we discuss the set of hyper-parameters tried during model selection for the fixed baseline and for AWN.
 We try our best to keep architectural choices identical with the exception of the width of the MLPs used by each model.
-In all experiments, we either run patience-based early stopping [46] or simply select the epoch with best validation score.
+In all experiments, we either run patience-based early stopping (Prechelt, 1998) or simply select the epoch with best validation score.
 G.1 Synthetic Tabular Datasets and Permuted MNIST
-We report the hyper-parameter configurations tried for the fixed MLP/RNN and AWNN in Table 3.
-In particular, the starting width of AWNN under the quantile function evaluated at $k=0.9$ is approximately $256$ hidden neurons.
+We report the hyper-parameter configurations tried for the fixed MLP/RNN and AWN in Table 4.
+In particular, the starting width of AWN under the quantile function evaluated at $k=0.9$ is approximately $256$ hidden neurons.
 We also noticed that a learning rate value of $0.1$ makes learning the width unstable, so we did not use it in our experiments.
-The total number of configurations for the fixed models is 180, for AWNN is 18.
-We also did not impose any prior on an expected size to see if AWNN recovers a similar width compared to the best fixed model (please refer to Table 1) starting from the largest configuration among the fixed models.
-Table 3: Hyper-parameter configurations for standard MLP/RNN and AWNN versions on tabular datasets.
-Hyper-Parameter DoubleMoon Spiral SpiralHard
-Batch Size 32 $[32,128]$ (MLP/RNN), 128 (AWNN) $[32,128]$ (MLP/RNN), 128 (AWNN)
+The total number of configurations for the fixed models is 180, for AWN is 18.
+We also did not impose any prior on an expected size to see if AWN recovers a similar width compared to the best fixed model (please refer to Table 1) starting from the largest configuration among the fixed models.
+Table 4: Hyper-parameter configurations for standard MLP/RNN and AWN versions on tabular datasets.
+Hyper-Parameter DoubleMoon Spiral SpiralHard/pol/MiniBooNE/credit c.
+Batch Size 32 $[32,128]$ (MLP/RNN), 128 (AWN) $[32,128]$ (MLP/RNN), 128 (AWN)
 Epochs 500 1000 5000
@@
 Non-linearity [ReLU, LeakyReLU, ReLU6]
-Optimizer Adam, learning rate $\in[0.1,0.01]$ (MLP/RNN), $0.01$ (AWNN)
-AWNN Specific
+Optimizer Adam, learning rate $\in[0.1,0.01]$ (MLP/RNN), $0.01$ (AWN)
+AWN Specific
 Quantile Fun. Threshold $k$ $0.9$
@@
 After previous results on tabular datasets showed that a LeakyReLU performs very well, we fixed it in the rest of the experiments.
-Table 4: Hyper-parameter configurations for standard MLP and AWNN versions on image classification datasets.
+Table 5: Hyper-parameter configurations for standard MLP and AWN versions on image classification datasets.
 Hyper-Parameter MNIST / CIFAR10 / CIFAR100
@@
 Scheduler Multiply learning rate by $0.1$ at epochs $50,100,150$
-AWNN Specific
+AWN Specific
 Quantile Fun. Threshold $k$ $0.9$
@@
 Prior over $\lambda$ Uninformative
-In this case, we test 5 different configurations of the fixed baseline, whereas we do not have to perform any model selection for AWNN.
+In this case, we test 5 different configurations of the fixed baseline, whereas we do not have to perform any model selection for AWN.
 The rate of the exponential distribution has been chosen to have a starting width of approximately $128$ neurons, which is the median value among the ones tried for the fixed model.
@@
 Due to the cost of training a Transformer architecture from scratch on a medium size dataset like Multi30k, we use most hyper-parameters’ configurations of the base Transformer in Vaswani et al.
-[58], with the difference that AWNN is applied to the MLPs in each encoder and decoder layer.
+(2017), with the difference that AWN is applied to the MLPs in each encoder and decoder layer.
 We train an architecture with 6 encoder and 6 decoder layers for 500 epochs, with a patience of 250 epochs applied to the validation loss.
@@
 The embedding size is fixed to 512 and the batch size is 128.
-The AWNN version starts with an exponential distribution rate of 0.004, corresponding to approximately 512 neurons, and a $\sigma^{\alpha}_{\ell}$=1.
+The AWN version starts with an exponential distribution rate of 0.004, corresponding to approximately 512 neurons, and a $\sigma^{\alpha}_{\ell}$=1.
 G.4 Graph Classification Tasks
-We train the AWNN version of Graph Isomorphism Network, following the setup of Errica et al.
-[16] and reusing the published results.
+We train the AWN version of Graph Isomorphism Network, following the setup of Errica et al.
+(2020) and reusing the published results.
 We test 12 configurations as shown in the table below.
 Please note that an extra MLP with 1 hidden layer is always placed before the sequence of graph convolutional layers.
-Table 5: Hyper-parameter configurations for AWNN versions on graph classification datasets.
+Table 6: Hyper-parameter configurations for AWN versions on graph classification datasets.
 Hyper-Parameter NCI1 / REDDIT-B
@@
 Optimizer Adam, learning rate $0.01$
-AWNN Specific
+AWN Specific
 Patience 500 epochs
@@
 Appendix H Re-training a fixed network with the learned width
-While on DoubleMoon AWNN learns a similar number of neurons chosen by the model selection procedure and the convergence to the perfect solution over 10 runs seem much more stable, this is not the case on Spiral, SpiralHard and REDDIT-B, where AWNN learns a higher total width over layers.
-To investigate whether the width learned by AWNN on these datasets is indeed the reason behind the better performances (either a more stable training or better accuracy), we run again the experiments on the baseline networks, this time by fixing the width of each layer to the average width over hidden layers learned by AWNN and reported in Table 1.
+While on DoubleMoon AWN learns a similar number of neurons chosen by the model selection procedure and the convergence to the perfect solution over 10 runs seem much more stable, this is not the case on Spiral, SpiralHard and REDDIT-B, where AWN learns a higher total width over layers.
+To investigate whether the width learned by AWN on these datasets is indeed the reason behind the better performances (either a more stable training or better accuracy), we run again the experiments on the baseline networks, this time by fixing the width of each layer to the average width over hidden layers learned by AWN and reported in Table 1.
 We call this experiment “Fixed+”.
@@
 Fixed 99.5(0.5) 98.0(2.0) 87.0(4.4)
-AWNN 99.8(0.1) 100.0(0.0) 90.2(1.3)
+AWN 99.8(0.1) 100.0(0.0) 90.2(1.3)
 Fixed+ 100.0(0.1) 83.2(17.0) 90.7(1.4)
 The results suggest that by using more neurons than those selected by grid search – possibly because the validation results were identical between different configurations – it is actually possible to improve both accuracy and stability of results on Spiral and REDDIT-B, while we had a convergence issue on SpiralHard.
-Notice that the number of neurons learned by AWNN on REDDIT-B was well outside of the range investigated by Xu et al.
-[65] and Errica et al.
-[16], demonstrating that the practitioner’s bias can play an important role on the final performance.
+Notice that the number of neurons learned by AWN on REDDIT-B was well outside of the range investigated by Xu et al.
+(2019) and Errica et al.
+(2020), demonstrating that the practitioner’s bias can play an important role on the final performance.
+Because we are not aware of the best width achieved by the grid search approach of Errica et al.
+(2020) on REDDIT-B, we ran an additional experiment on fixed models, testing smaller and larger widths compared to the one selected by AWN (average of $\sim$793).
+What we discover is the following.
+Table 7: We report REDDIT-B validation performances, averaged across the 10 outer folds, for a fixed-width DGN.
+The other hyper-parameters match those selected by AWN during model selection, as per this specific experiment.
+# Total Width Mean Std
+50 89.28 3.50
+250 92.89 1.45
+500 92.67 2.03
+793 92.83 3.55
+1000 93.06 1.60
+1250 92.78 2.33
+1500 91.78 2.41
+Though these results are statistically comparable, the average performance obtained using AWN’s learned width seems to be close to a local optimum for the fixed-width model.
 Appendix I Further qualitative results on image classification datasets
-Because it is well known that MNIST, CIFAR10, and CIFAR100 are datasets of increasing difficulty, we attach a figure depicting the learned number of neurons and the convergence of AWNN on these datasets, following the practices highlighted in He et al.
-[26] and described in Table 4 to bring the ResNet20 to convergence.
+Because it is well known that MNIST, CIFAR10, and CIFAR100 are datasets of increasing difficulty, we attach a figure depicting the learned number of neurons and the convergence of AWN on these datasets, following the practices highlighted in He et al.
+(2016) and described in Table 5 to bring the ResNet20 to convergence.
 Akin to the tabular datasets, we see a clear pattern where the learned number of neurons, averaged over the 10 final training runs, increases together with the task difficulty.
-The AWNN version of the ResNet20 converges in the same amount of epochs as the fixed one.
+The AWN version of the ResNet20 converges in the same amount of epochs as the fixed one.
 Figure 9: We present an analysis similar to Figure 2 but for image classification datasets of increasing difficulty. One can observe the jumps caused by the learning rate’s scheduling strategy.
-Appendix J Non-linearity ablation study
+Appendix J Comparison with Neural Architecture Search (NAS)
+Although complementary to our approach, we thought instructive to run a some basic NAS methods on the image classification datasets.
+Most NAS methods fix an upper bound on the maximum width, whereas AWN’s methodology does not.
+The truncation ability of AWN is not intended as a substitute for pruning but rather as a confirmation of our claims.
+In fact, NAS and pruning can actually be combined with AWN and do not necessarily be seen as “competing” strategies.
+The results are shown in Table 8.
+That said, we compared the performance of AWN against simple NAS methods on the vision datasets.
+We tested grid search, random search (Following Wu et al.
+(2020)), local search (White et al., 2021), and Bayesian Optimization (Barber, 2012), fixing the same budget of 5 for all methods and width values between 0 (Linear classifier) and 512 neurons.
+Results are comparable to those of AWN.
+Table 8: Comparison between different NAS approaches and AWN. We remind the reader that NAS is an orthogonal and complementary research direction to AWN.
+Grid Search Random Search Local Search Bayes. Optim. AWN
+MNIST 99.6(0.1) 99.6(0.0) 99.5 99.4 99.7(0.0)
+CIFAR10 91.4(0.2) 90.1(0.5) 90.6 91.2 91.4(0.2)
+CIFAR100 66.5(0.4) 64.9(1.1) 64.9 65.9 63.1(4.0)
+Appendix K Non-linearity ablation study
 Below, we report an ablation study on the impact of non-linear activation functions.
@@
 The decreasing trend for LeakyReLU and ReLU activations may suggest that these configurations are also converging to a similar value than ReLU6, but they are taking a much longer time.
-Appendix K Learned width for each layer on Graph Datasets
+Appendix L Learned width for each layer on Graph Datasets
 Figure 11: Learned neurons per layer (from 1 to 5), averaged over 10 final runs for each of the 10 best configurations selected in the outer folds.
@@
 One can observe similar trends in most cases.
-Appendix L Comparison with Neural Architecture Search (NAS)
-Although complementary to our approach, we thought instructive to run a some basic NAS methods on the image classification datasets.
-Most NAS methods fix an upper bound on the maximum width, whereas AWNN’s methodology does not.
-The truncation ability of AWNN is not intended as a substitute for pruning but rather as a confirmation of our claims.
-In fact, NAS and pruning can actually be combined with AWNN and do not necessarily be seen as “competing” strategies.
-The results are shown in Table 6.
-That said, we compared the performance of AWNN against simple NAS methods on the vision datasets.
-We tested grid search, random search (Following Wu et al.
-[64]), local search [60], and Bayesian Optimization [1], fixing the same budget of 5 for all methods and width values between 0 (Linear classifier) and 512 neurons.
-Results are comparable to those of AWNN.
-Table 6: Comparison between different NAS approaches and AWNN. We remind the reader that NAS is an orthogonal and complementary research direction to AWNN.
-Grid Search Random Search Local Search Bayes. Optim. AWNN
-MNIST 99.6(0.1) 99.6(0.0) 99.5 99.4 99.7(0.0)
-CIFAR10 91.4(0.2) 90.1(0.5) 90.6 91.2 91.4(0.2)
-CIFAR100 66.5(0.4) 64.9(1.1) 64.9 65.9 63.1(4.0)
+Appendix M Quantile Ablation Study
+Figure 12: We report learned width and validation accuracy trends averaged over 1-layer AWN configurations on SpiralHard.
+We see how a higher quantile $k$ (i.e., a better ELBO approximation) grants more stable performances.
+All architectures keep adapting to the task.
+Appendix N Regularization Ablation Study
+In Section 3, we argued there are two ways we can maximize the importance rescaling effect, namely by using bounded activation functions and by regularizing the weights.
+In this section, we show with an ablation that these strategies are generally not required, in the sense that AWN is still able to solve the tasks, although it might introduce a more neurons than needed.
+The table below compares our original results on tabular data (Table 1) with those where we did not consider ReLU6 and where we imposed an almost null regularizer on the weights ($\sigma_{\ell}^{\theta}=100$).
+Table 9: We compare performances and learned width of AWN, where we removed regularization and bounded activations from hyper-parameters, with the results in the main paper.
+AWN AWN - NoReg Width (AWN) Width (AWN - NoReg)
+Mean (Std) Mean (Std) Mean (Std) Mean (Std)
+DoubleMoon 100.0 (0.0) 100.0 (0.2) 8 (2.8) 17 (16.6)
+Spiral 99.8 (0.1) 100.0 (0.0) 66 (8.7) 108 (26.8)
+SpiralHard 100.0 (0.0) 100.0 (0.0) 227 (32.4) 539 (809.8)
+pol 99.2 (0.1) 99.2 (0.1) 84 (11.0) 2335 (552.8)
+MiniBooNE 93.2 (0.1) 93.8 (0.2) 53 (11.1) 4907 (1141)
+credit card 81.8 (0.1) 81.8 (0.1) 51 (12.0) 53 (13)
+We see that in all cases, not using regularization nor bounded activations leads to comparable performances but definitely higher widths.
+For some of the configurations tried, the width tends to grow a lot, which is why we set a maximum width of 5000 in the interest of time, although performances do not seem to decrease in spite of that.
+Further investigation into the results revealed that the weight regularization is what helps the most in controlling a stable behavior of the width, while the bounded activation did not have a clear impact on that.
+We therefore recommend some degree of regularization to make sure that the effect of the importance rescaling is properly taken into account by the optimization process.
+Appendix O Wall-Clock Time Comparison
+To further highlight the advantages brought by AWN, we report the average runtime (in seconds) required to complete a single model selection configuration for AWN and for the fixed model.
+In cases where model selection is not performed (e.g., AWN on vision tasks), we instead report the average runtime of a final run.
+Results for NCI1 and REDDIT are excluded from this comparison, since they were taken directly from the literature.
+Based on these runtimes, we compute two measures: i) the ratio of average single-run times (Single-Run Ratio), and ii) the ratio of total time required by AWN and the fixed models to complete the entire model selection process (Model Selection Ratio).
+Table 10: We report average wall-clock times for AWN and the fixed model, measured either over model selection configurations or final runs when no selection is performed.
+We also provide two ratios: i) single-run wall-clock times and ii) total model selection time (fixed models always test at least five widths).
+Dataset Avg Fixed (s) Avg AWN (s) Single-Run Ratio ($\downarrow$) Model Selection Ratio ($\downarrow$)
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+The results confirm that AWN’s overhead of adapting the network vanishes for larger datasets with higher feature dimensionality.
+Interestingly, on some of these datasets the average single-run wall time is even shorter than the fixed models.
+By being able to reduce hyper-parameter tuning of the width while preserving most performances, the total model selection time ratios are consistently in favor of AWN, and in principle scale favorably with the number of widths tested.
+Note that memory consumption scales with the width of the layers, but our methodology does not introduce any noticeable extra memory requirements.
+Appendix P Impact of Removing Most Important Neurons
+We complement the truncation analysis of Figure 5 by showing what happens when we remove individual neurons or a group of them, starting from the most important ones.
+This analysis provides further insights about how neurons behave in AWN.
+Figure 13: We show the impact of removing the most important neurons, either in batch (top) or individually (bottom) on Spiral performances, using the same experiment as Figure 5.
+We observe that performance quickly drops, as expected, and we reach random performance after removing the first $\sim$40 neurons out of 83.
+The second analysis prunes individual neurons to understand how much their absence influences performance – which can have a different effect compared to removal of a set of neurons – to see if there exist two neighboring neurons that contribute similarly to the performance drop or do not contribute at all.
+Removing most of the first 25 neurons out of 83 contributes majorly to performance degradation, but interestingly the removal of some neurons does not affect performances.
+Instead, removing any subsequent neuron has little impact (but for one) on performances, as one would expect.
+This suggests that we may be able to compress the neural network even more: we leave this investigation to future work.